### v1

# Enhancing the Harrow-Hassidim-Lloyd (HHL) algorithm in systems with large condition numbers

Abstract
Although the Harrow-Hassidim-Lloyd (HHL) algorithm offers an exponential speedup in system size for treating linear equations of the form $A\vec{x}=\vec{b}$ on quantum computers when compared to their traditional counterparts, it faces a challenge related to the condition number ($\mathcal{\kappa}$) scaling of the $A$ matrix. In this work, we address the issue by introducing the post-selection-improved HHL (Psi-HHL) approach that involves a simple yet effective modification of the HHL algorithm to extract a feature of |$\ket{x}$, and which leads to achieving optimal behaviour in $\mathcal{\kappa}$ (linear scaling) for large condition number situations. This has the important practical implication of having to use substantially fewer shots relative to the traditional HHL algorithm. We carry out two sets of simulations, where we go up to 26-qubit calculations, to demonstrate the ability of Psi-HHL to handle situations involving large $\mathcal{\kappa}$ matrices via: (a) a set of toy matrices, for which we go up to size $64\times 64$ and $\mathcal{\kappa}$ values of up to $\approx$ 1 million, and (b) a deep-dive into quantum chemistry, where we consider matrices up to size $256\times 256$ that reach $\mathcal{\kappa}$ of about 466. The molecular systems that we consider are Li2, RbH, and CsH. Although the feature of |$\ket{x}$ considered in our examples is an overlap between the input and output states of the HHL algorithm, our approach is general and can be applied in principle to any transition matrix element involving |$\ket{x}$.
preprint: APS/123-QED
I Introduction
[…]
Despite possessing the aforementioned advantages, the HHL algorithm relies upon the probability of successfully post-selecting $1$ on the HHL ancilla qubit, that is, $P(1)$ [1]. It is known that there is a substantial decrease in $P(1)$ with increase in the condition number of the problem matrix $A$, which we shall denote by $\mathcal{\kappa}$. In other words, the fraction of usable shots in the HHL algorithm becomes very low when large $\mathcal{\kappa}$ input matrices are involved. Unfortunately, it is not uncommon to encounter the issue while employing HHL for some interesting problems (for example, see Ref. [3] for application of the algorithm to hydrological systems), including quantum chemistry, which we shall discuss in detail in our Results section. Several works in literature have addressed the issue of handling large $\mathcal{\kappa}$ matrices either directly (where $\mathcal{\kappa}$ of the $A$ matrix itself is altered via approaches such as preconditioning) or indirectly [4, 5, 6, 7, 8, 3, 9]. For a thorough survey, we direct the reader to Section 1.2 of Ref. [10].
In this work, we propose a new method to address the issue of handling $A$ matrices with large condition numbers, by introducing the post-selection-improved HHL (Psi-HHL) algorithm, which relies upon the intuition of extracting the correct signal by subtracting a wrong signal from a mixed signal (that contains both the correct and wrong signals). A signal, in the context of our work, refers to a scalar that is a feature of |$\ket{x}$. For the purposes of this work, we pick the overlap between the input and output states of HHL, which in the case of molecules corresponds to their correlation energies (up to some normalization constants), as our signal, although it is to be noted that Psi-HHL itself is applicable to features as general as transition matrix elements. The wrong signal is obtained by successfully post-selecting outcome bit ‘0’ instead of ‘1’ (we call this approach HHL1). The mixed signal is generated in an HHL quantum circuit by introducing an appropriately chosen $2\times 2$ unitary just before measurement of the HHL ancilla (we name this procedure HHL2). Figure 1 presents schematic diagrams outlining the workings of the HHL and the Psi-HHL algorithms.
This simple modification involving execution of two HHL routines enables one to carry out calculations for treating systems of linear equations involving large-$\mathcal{\kappa}$ matrices without having to contend with a large number of shots. In fact, we argue that the condition number scales linearly with system size with our approach for large $\mathcal{\kappa}$ situations, while still staying in the realm of a typical HHL computation. This compares favourably with several other notable works in literature, including but not limited to the original HHL work: $\mathcal{O}(\mathcal{\kappa}^{3})$ and $\mathcal{O}(\mathcal{\kappa}^{2})$ without and with amplitude amplification respectively [1], $\mathcal{O}(\mathcal{\kappa}\mathrm{log}^{3}\mathcal{\kappa})$ with variable time amplitude amplification [5], and works by Childs et al: $\mathcal{O}(\mathcal{\kappa}\mathrm{polylog}(\mathcal{\kappa}))$ [11], Wossnig et al: $\mathcal{O}(\mathcal{\kappa}^{2})$, and Costa et al’s discrete quantum adiabatic approach based quantum linear solver algorithm [12]: $\mathcal{O}(\mathcal{\kappa})$. In particular, the last approach also achieves linear scaling in condition number, but we do so without invoking any additional algorithmic machinery beyond the scope of the traditional HHL algorithm. In order to provide a very qualitative perspective, we note that for condition number of about 1 million, $\mathcal{\kappa}^{3}:\mathcal{\kappa}^{2}:\mathcal{\kappa}\mathrm{log}^{3}\mathcal{\kappa}:\mathcal{\kappa}$ is $\sim 10^{18}:10^{12}:7.9\times 10^{9}:10^{6}$. For completeness, we add that the best known classical approach, conjugate gradient, scales as $\mathcal{O}(\sqrt{\mathcal{\kappa}})$ [13], but its scaling with respect to system size is, of course, exponentially more expensive than that for HHL.
The paper is organized as follows: Section II presents the theory and methodology involved in our work by introducing the Psi-HHL approach, including a discussion on the complexity theoretic aspects (in $\mathcal{\kappa}$) of Psi-HHL. Section III presents our results for Psi-HHL applied to (a) suitably chosen sets of $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$ toy matrices, as well as for (b) molecular calculations ($64\times 64$ and $256\times 256$ $A$ matrix sizes) in the linearized coupled cluster (LCC) HHL and Psi-HHL frameworks. We conclude in Section IV.
Figure 1: (a) Schematic of the HHL algorithm, where 1 is post-selected and for which $P(1)$ would be low for large $\mathcal{\kappa}$. |$\ket{b}$ is the input state register with $n_{b}$ qubits, while $n_{r}$ refers to the number of clock register qubits. The Hong-Ou-Mandel (HOM) module is a controlled-SWAP test without an additional ancilla qubit, and we use it to extract |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ | (interpreted as the correlation energy, $E_{\mathrm{c}}^{\mathrm{HHL}}$, when HHL is applied to molecular calculations), which is the feature of |$\ket{x}$ that we are interested in. Sub-figure (b) provides a block-diagram description of the HHL and the HOM modules. $QPE$ refers to the quantum phase estimation while $CR(\theta)$ refers to the controlled-rotation module that inverts the eigenvalues of $A$. The dashed cambered rectangle around the CNOT gate in the HOM circuit module, and which is accompanied by $(i)$, indicates that the gate is controlled on the ith qubit of the |$\ket{b}$ register. Sub-figure (c) consists of HHL1 in the top panel, where $P(0)$ is post-selected, and HHL2 in the bottom panel, where the traditional HHL module is appended with an additional mixing unitary, $R_{Y}(2\alpha)$, and where $P^{\prime}(1)$ is post-selected. Psi-HHL relies on subtracting the final outcomes of HHL1 and HHL2, and as the expression at the bottom of panel (c) indicates, we recover the correlation energy via the procedure to within a simple factor. In our notation, $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the correlation energies obtained from HHL and Psi-HHL respectively. Further, $a_{i},b_{i},c_{i},d_{i}\in\mathbb{R}$ are constants.
II Theory and Methodology
II.1 The HHL algorithm
We briefly sketch the relevant details of the HHL algorithm (we refer the readers to Refs. [14, 15, 16, 17] for detailed descriptions of the algorithm) in Figs. 1(a) and 1(b), where an amplitude-encoded input |$\arrowvert b\rangle=\sum_{i}b_{i}\arrowvert\nu_{i}\rangle$, such that $\sum_{i}|b_{i}|^{2}=1$ where {|$\{\arrowvert\nu_{i}\rangle\}$ refers to the eigenbasis of $A$, is fed into the HHL algorithm along with $n_{r}$ number of clock register qubits and one HHL ancillary qubit. The HHL algorithm itself involves QPE, followed by a non-trivial controlled-rotation module, and then a QPE module. At the end of the circuit, we have the following state vector,
[…]
In the above expression, $\{\tilde{\lambda}_{i}\}$ refers to the estimates of the actual eigenvalues of $A$ ($1\geq\tilde{\lambda}_{i}\geq 1/\mathcal{\kappa}$) obtained via the QPE module of HHL. Typically, $C$ is chosen to be the minimum eigenvalue of $A$, but it is worth noting that incurring such a large classical overhead can always be circumvented via approaches such as AdaptHHL [2]. When one post-selects the outcome ‘1’ on the HHL ancilla qubit, one can think of it as choosing the ‘correct’ signal.
The solution vector $|x\rangle\propto A^{-1}|b\rangle$, and the input vector $|b\rangle$ can be written as,
$|x\rangle=|x\rangle_{\rm un}/\||x\rangle_{\rm un}\|\ \mathrm{and}\ \ket{b}=\ket{b}_{\rm un}/\|\ket{b}_{\rm un}\|$ (2)
respectively, where subscript ‘$\mathrm{un}$’ refers to unnormalized states— $|b\rangle_{\rm un}$ is the unnormalized input state after amplitude encoding and $|x\rangle_{\rm un}=\sum_{i}\frac{b_{i}C}{\tilde{\lambda}_{i}}|\nu_{i}\rangle$, with $\sum_{i}|b_{i}|^{2}=1$. The probability of successfully post-selecting outcome ‘$1$’ is given as $P(1)=\||x\rangle_{\rm un}\|^{2}$. Finally, when one wants to predict a feature of the solution vector, say, the correlation energy of a molecular system (proportional to the overlap between |$\ket{x}$ and |$\ket{b}$), it can be done by appending an additional circuit (the Hong-Ou-Mandel module [18]), abbreviated as HOM in Fig. 1(b) that inputs the output, |$\arrowvert x\rangle$, of the HHL algorithm and |$\arrowvert b\rangle$, to output the overlap between the two quantities. The correlation energy is then calculated as
$E_{c}^{\rm HHL}$ $=$ |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ |
$=$ |$-\|\ket{b}_{\rm{un}}\|^{2}\sum_{i}|b_{i}|^{2}\frac{C}{\tilde{\lambda}_{i}}.$
For the sake of simplicity, we set this quantity hereafter as the feature of |$\ket{x}$ that is of interest to us, for the rest of this work, although Psi-HHL can be applied to more general features, such as transition matrix elements involving |$\ket{x}$. Therefore, even for non-quantum chemistry examples that we shall later consider, we will use the term ‘correlation energy’ to refer to the quantity from Eq. II.1. The number of trials needed to successfully post-select outcome `$1$’ is $\mathcal{O}(\mathcal{\kappa}^{2})$. Thus, in the event of large $\mathcal{\kappa}$, the success probability falls quickly, making it practically difficult to carry computations using large $\mathcal{\kappa}$ matrices. In this work, we demonstrate a new approach— ‘Psi-HHL’, towards handling large $\mathcal{\kappa}$ matrices by leveraging the probability of successfully post-selecting the ‘wrong’ outcome `$0$’, $P(0)$, in HHL.
II.2 The Psi-HHL approach
In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Fig. 1(c). The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the traditional HHL given in Fig. 1(a) and Fig. 1(b). The HHL2 entails traditional HHL module appended with a mixing $2\times 2$ unitary, given as $R_{Y}(2\alpha)$ and post-selecting 1 from the mixed signal. The role of this unitary is to mix the coefficients of $|0\rangle$ and $|1\rangle$ in the traditional HHL (see Eq. 1), thereby outputting a ‘mixed’ signal. The core idea behind Psi-HHL is to subtract the correlation energies associated with HHL1 and HHL2 thus leaving us with the correlation energy corresponding to the ‘correct’ signal. The gain in taking this route is the efficiency that we achieve in the number of shots expended. We summarize this in the following proposition:
Proposition.
Psi-HHL yields the same correlation energy as traditional HHL through two implementations of HHL, but is more efficient in the number of shots than HHL, especially when dealing with $A$ matrices with large condition numbers.
We expound on the first part of the proposition in the following paragraphs, and the second in sub-section II.3.
II.2.1 HHL1: Computing the wrong signal
The first HHL module, HHL1, serves to post-select the ‘wrong’ signal of the HHL ancilla qubit, that is, the state attached to |$\ket{0}$ in Eq.1. We re-write the state at the output of all the registers $|\psi_{\rm out}\rangle$ except the clock qubits register, as follows,
[…]
$=$ $(|0\rangle\otimes|x_{\mathrm{w}}\rangle+|1\rangle\otimes|x_{\mathrm{r}}\rangle).$
Note that here $|x_{\mathrm{w}}\rangle$ and $|x_{\mathrm{r}}\rangle$ are the unnormalised vectors in the superposition. The subscripts $\rm w$ and $\rm r$ refer to ‘wrong’ and ‘correct’ respectively, where for the latter, we use $\rm r$ since $\rm c$ is used already to denote ‘correlation’. The normalised solution vector $|x_{0}\rangle$ which appears at the output register after successfully post-selecting outcome ‘00’ is given by,
|$\ket{x_{\rm 0}}$ $=$ $|x_{\mathrm{w}}\rangle/\||x_{\mathrm{w}}\rangle\|$ (5)
$=$ $\sum_{i}b_{i}\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}\ket{\nu_{i}}/\||x_{\mathrm{w}}\rangle\|,$
where,
$\||x_{\rm w}\rangle\|=\sqrt{P(0)}=\left(\sum_{i}\left|b_{i}\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}\right|^{2}\right)^{1/2}.$ (6)
Here, $P(0)$ is the probability of successfully post-selecting outcome ‘0’. The associated correlation energy is then given as,
$E_{\rm c}^{\rm w}=-\||b\rangle_{\rm un}\|^{2}\ \||x_{\rm w}\rangle\|\ |\langle b|x_{0}\rangle|.$ (7)
Here, $|b\rangle_{\rm un}$ is the unnormalized input vector after amplitude encoding, and
$|\langle b|x_{0}\rangle|=\sum_{i}|b_{i}|^{2}\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}.$ (8)
II.2.2 HHL2: Computing the mixed signal
The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancilla before the measurement as illustrated in the lower panel of Fig. 1(c). This seemingly trivial addition to the algorithm introduces a new degree of freedom via the angle $\alpha$, enabling one to boost the probability of successful matrix inversion in the case of large $\mathcal{\kappa}$. We recall that $R_{Y}(2\alpha)$ is a rotation about the $Y$ axis on the Bloch sphere as given below
$R_{Y}(2\alpha)=\begin{pmatrix}\cos({\alpha})&-\sin({\alpha})\\ \sin({\alpha})&\cos({\alpha})\end{pmatrix}.$ (9)
The state vector at the output of HHL2 module $|\psi^{\prime}_{\rm out}\rangle$, ignoring the clock qubits register, after applying the unitary in Eq. 9 is given below,
$|\psi^{\prime}_{\rm out}\rangle$ $=$ $\sum_{i}b_{i}\left(\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}\cos({\alpha})-\frac{C}{\tilde{\lambda}_{i}}\sin({\alpha})\right)\ket{0}\otimes\ket{\nu_{i}}$ (10)
$+$ $b_{i}\left(\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}\sin({\alpha})+\frac{C}{\tilde{\lambda}_{i}}\cos({\alpha})\right)\ket{1}\otimes\ket{\nu_{i}}$
$=$ $(|0\rangle\otimes|x_{\rm m}\rangle+|1\rangle\otimes|x^{\prime}_{\rm m}\rangle).$
Note that in Eq. 10, |$\ket{x_{\mathrm{m}}}$ and |$\ket{x^{\prime}_{\mathrm{m}}}$ are unnormalized vectors. The subscript ‘m’ denotes ‘mixed’, and while the unprimed vector accompanies |$\ket{0}$, the primed one accompanies |$\ket{1}$. The solution vector at the output after successfully post-selecting outcome ‘$1$’ is given below
|$\ket{x^{\prime}}$ $=$ |$\frac{\ket{x^{\prime}_{\rm m}}}{\|\ket{x^{\prime}_{\rm m}}\|}$
$=$ $\sum_{i}b_{i}\left(\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}\sin({\alpha})+\frac{C}{\tilde{\lambda}_{i}}\cos({\alpha})\right)\ket{\nu_{i}}/\|\ket{x^{\prime}_{\rm m}}\|,$
where
|$\|\ket{x^{\prime}_{\rm m}}\|$ $=$ $\sqrt{P^{\prime}(1)}$ (11)
$=$ $\left(\sum_{i}\left|b_{i}\left(\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}\sin({\alpha})+\frac{C}{\tilde{\lambda}_{i}}\cos({\alpha})\right)\right|^{2}\right)^{1/2}.$
Here, $P^{\prime}(1)$ is the probability of successfully obtaining outcome ‘1’ in HHL2. The correlation energy is then obtained as
$E_{\mathrm{c}}^{\rm m}=-\||b\rangle_{\rm un}\|^{2}\ \||x_{\rm m}^{\prime}\rangle\|\ |\langle b|x^{\prime}\rangle|,$ (12)
where
$|\langle b|x^{\prime}\rangle|=\sum_{i}|b_{i}|^{2}\left(\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}\sin(\alpha)+\frac{C}{\tilde{\lambda}_{i}}\cos(\alpha)\right).$ (13)
We choose $\alpha$ such that $\sin({\alpha})$ and $\cos({\alpha})$ are positive and $P^{\prime}(1)\gg P^{\prime}(0)$. $P^{\prime}(0)$ refers to the probability of post-selecting outcome ‘0’ in HHL2.
II.2.3 Psi-HHL: Subtracting signals
In order to prove this statement, we subtract Eq. 12 and Eq. 7, thus obtaining
$E_{c}^{\mathrm{Psi-HHL}}\mathrm{cot(\alpha)}$ $=$ $\frac{E_{\mathrm{c}}^{\mathrm{m}}}{\sin(\alpha)}-E_{\mathrm{c}}^{\mathrm{w}}$ (14)
$=$ |$-\|\ket{b}_{\rm un}\|^{2}\sum_{i}|b_{i}|^{2}\frac{C}{\tilde{\lambda}_{i}}\rm cot(\alpha).$
By comparing the right hand sides of Eq. II.1 and Eq. 14, we see that the expressions for $E_{\mathrm{c}}^{\rm HHL}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ differ only by a factor of $\rm cot(\alpha)$ as shown below.
$E_{\mathrm{c}}^{\mathrm{Psi-HHL}}=\frac{\frac{E_{\mathrm{c}}^{\mathrm{m}}}{\rm sin(\alpha)}-E_{\mathrm{c}}^{\mathrm{w}}}{\rm cot(\alpha)}=E_{\mathrm{c}}^{\rm HHL}.$ (15)
Although theoretically, we obtain the same correlation energy from Psi-HHL, we observe that this technique is superior to traditional HHL in the way it handles large $\mathcal{\kappa}$ matrices, due to the optimal probability of success that one achieves while post-selecting, as we explain in the subsequent section. Therefore, $E_{c}^{\mathrm{Psi-HHL}}$ can predict the true correlation energy, $E_{c}$, to a reasonable level of precision with fewer shots even for values of $\mathcal{\kappa}$ where HHL procedure fails to do so, which we illustrate in the Results section with examples. We find through our examples that for up to fairly large values of $\mathcal{\kappa}$ that are well beyond those that HHL can handle, Psi-HHL predicts the correlation energy with fewer shots. It is also important to note that throughout the derivation discussed above, we can replace |$\ket{b}$ with $O\ket{y}$, where $O$ is a suitably block-encoded observable [19], and |$\ket{y}$ an arbitrary state that can be an output of another algorithm, and our results from Psi-HHL still hold. We also add that in problems where it is not immediately obvious a priori that the condition number is large, one cannot determine whether to employ HHL or Psi-HHL. However, this is not a problem as Psi-HHL intrinsically involves the regular HHL procedure, and hence if $P(1)$ is significant, one can choose not to opt for the Psi-HHL procedure.
II.3 Complexity aspects
We now demonstrate the complexity (in $\mathcal{\kappa}$) of the Psi-HHL algorithm. Since for the purposes of deriving condition number behaviour at the measurement-post-selection part of HHL, the distinction between the estimated eigenvalues of $A$, that is, $\tilde{\lambda}_{i}$, and the actual eigenvalues, $\lambda_{i}$, does not matter, we denote all estimated eigenvalues in this section as simply $\lambda_{i}$ for the sake of convenience.
In order to elaborate on the second part of our Proposition, we invoke the probability with which one post selects the desired outcome successfully in terms of $\kappa$.
a) Success probability of the post-selection process in HHL1: The probability of successfully post-selecting the outcome ‘0’, $P(0)$, in HHL1 as given in Eq. 6 can be expanded as [20]
$P(0)$ $=$ $\sum_{i}\left|b_{i}\sqrt{1-\frac{C^{2}}{{\lambda}_{i}^{2}}}\right|^{2}$ (16)
$\approx$ $\|\sqrt{(\mathbb{I}-C^{2}A^{-2})}|b\rangle\|^{2},$
where $\mathbb{I}$ is the identity operator written in the eigenbasis of operator $A$, constant $C=O(1/\mathcal{\kappa})$, where the condition number $\mathcal{\kappa}$ is taken as $\lambda_{\rm max}/\lambda_{\rm min}$ and $|b\rangle$ is the input to the HHL circuit. The set of values $\bigg{\{}\sqrt{1-\frac{C^{2}}{\lambda_{i}^{2}}}\bigg{\}}$ are the eigenvalues of the operator $\sqrt{\mathbb{I}-C^{2}A^{-2}}$, with the largest eigenvalue being $\sqrt{1-\frac{C^{2}}{\lambda_{\rm max}^{2}}}$, and therefore
$P(0)$$\leq$ $\|\sqrt{(\mathbb{I}-C^{2}A^{-2})}\|_{\rm op}^{2}\||b\rangle\|^{2}$
$P(0)$$\leq$ $1-\frac{C^{2}}{\lambda_{\rm max}^{2}}$ (17)
$\leq$ $1-\frac{1}{\mathcal{\kappa}^{2}}.$
The first line uses the property of sub-multiplicativity of matrix norms, that is, $\|AB\|\leq\|A\|\|B\|$. In Eq. II.3, we recall that we set $C=\lambda_{\rm min}$, .$\|.\|_{\rm op}$ is the operator norm, and $\||b\rangle\|=1$. Note that the smallest eigenvalue of this operator is $\sqrt{1-\frac{C^{2}}{\lambda_{\rm min}^{2}}}=0$, and corresponds to $P(0)=0$, which leads to expending a large number of shots in HHL1. One could circumvent this issue by appropriately choosing a scaled $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, leading to $P(0)\geq 1-\gamma^{2}$. This idea has been illustrated via an example as shown in Fig. 2, where we choose the scaling $\gamma=0.5$.
Refer to caption Refer to caption
(a) (b)
Figure 2: Figure showing probabilities of post-selecting various outcomes in Psi-HHL. Sub-figures (a) and (b) show the behaviour of $P(0)$ and $P^{\prime}(1)$ respectively with $\mathcal{\kappa}$, for $A$ being a $4\times 4$ diagonal matrix and with six different choices for the vector $\vec{b}$: $P(0)_{\mathrm{vmin}}$ corresponds to the case where $\vec{b}$ is the eigenvector associated with the minimum eigenvalue of $A$, $P(0)_{\mathrm{vmax}}$ is the case where $\vec{b}$ is the eigenvector associated with the maximum eigenvalue of $A$, $P(0)_{\lambda_{2}}$ and $P(0)_{\lambda_{3}}$ correspond to the situations where $\vec{b}$ is the eigenvector associated with two other eigenvalues of $A$, $P(0)_{\mathrm{equal}}$ is the case where $\vec{b}$ is in equal superposition of all eigenstates of $A$ and finally $P(0)_{\mathrm{unequal}}$ considers the scenario where $\vec{b}$ is in an arbitrary superposition of eigenstates of $A$. In the left-most figure of sub-figure (a), we choose the parameter $C$ as $\lambda_{\mathrm{min}}$, that is, the minimum eigenvalue of $A$, whereas in right-most figure of sub-figure (a), $C=0.5\lambda_{\mathrm{min}}$, that is, $\gamma$ is set to 0.5. Figures (b) represent $P^{\prime}(1)$ where the legend’s labels have a similar interpretation to those of $P(0)$. The HHL and Psi-HHL calculations were done using 1 million shots, and $\alpha$ was set to $60^{\circ}$ for Psi-HHL.
b) Success probability of the post-selection process in HHL2: A similar expression for the probability of successful post-selection in HHL2, $P^{\prime}(1)$, can be obtained by further expanding Eq. 11 as follows,
$P^{\prime}(1)$$\approx$ $\rm\|[\sqrt{(\mathbb{I}-C^{2}A^{-2}})\mathrm{sin}(\alpha)$ (18)
$+$ $(CA^{-1})\mathrm{cos}(\alpha)]|b\rangle\|^{2}.$
The operator $B\coloneq\sqrt{\mathbb{I}-C^{2}A^{-2}}\sin(\alpha)+(CA^{-1})\cos(\alpha)$ in its eigenbasis $\{|\nu_{i}\rangle\}$ could be given in the diagonal form as
$\mathrm{diag}(\cdots,\sqrt{1-C^{2}/\lambda^{2}_{\mathrm{max}}}\sin(\alpha)+(C/\lambda_{\mathrm{max}})\cos(\alpha),\cdots,$
$\sqrt{1-C^{2}/\lambda^{2}_{\mathrm{min}}}\sin(\alpha)+(C/\lambda_{\mathrm{min}})\cos(\alpha),\cdots),$
where the entries that occur in the parentheses after ‘diag’ are the diagonal elements of the matrix. Note that the largest and smallest eigenvalue of the matrix given in Eq. II.3 depend on the choice of $\alpha$. When $\alpha=\pi/2$, the component of that eigenvalue attached to $\sin(\alpha)$ remains, and we do not extract the desired solution. However, when $\alpha$ is too small, say 00, the component attached to $\cos(\alpha)$ remains and we get the same complexity as the traditional HHL approach. Therefore, $\alpha$ is chosen to satisfy the following properties:
  • The success probability, $P^{\prime}(1)$ is large enough.
  • The solution component that matters to us does not vanish in the superpositions that occur along the diagonal entries in Eq. II.3.
We demonstrate the case where $\alpha$ is close to $\pi/2$ and $C=\lambda_{\min}$, in which case the operator in Eq. II.3 could be written in the form where the largest eigenvalue is the first entry and the smallest eigenvalue is the last entry, as given below,
$\mathrm{diag}(\cdots,\sqrt{1-1/\mathcal{\kappa}^{2}}\sin(\alpha)+(1/\mathcal{\kappa})\cos(\alpha),\cdots,\cos(\alpha)).$
Using once again the sub-multiplicativity of matrix norms in Eq. 18, we can write
$P^{\prime}(1)$$\leq$ $\rm\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
$+$ $CA^{-1}\cos(\alpha)]\|^{2}_{\rm op}\||b\rangle\|^{2},$
where .$\|.\|_{\rm op}$ is the operator norm yielding the largest eigenvalue of the operator, and $\||b\rangle\|=1$. Then $P^{\prime}(1)$ is bounded the following way,
$\cos^{2}(\alpha)\leq P^{\prime}(1)$$\leq$ $\rm\bigg{|}\bigg{[}\bigg{(}\sqrt{1-\frac{1}{\mathcal{\kappa}^{2}}}\bigg{)}sin(\alpha)+\frac{1}{\mathcal{\kappa}}cos(\alpha)\bigg{]}\bigg{|}^{2}$ (22)
$=$ $\rm\bigg{(}1-\frac{1}{\mathcal{\kappa}^{2}}\bigg{)}\sin^{2}(\alpha)+\frac{1}{\mathcal{\kappa}^{2}}cos^{2}(\alpha)$
$+$ $2\left(\sqrt{1-\frac{1}{\mathcal{\kappa}^{2}}}\right)\frac{1}{\mathcal{\kappa}}\sin(\alpha)cos(\alpha)$
$\approx$ $\rm sin^{2}(\alpha)+\frac{1}{\mathcal{\kappa}}\ sin(2\alpha)+\frac{1}{\mathcal{\kappa}^{2}}\cos(2\alpha)$
-- $\frac{1}{\mathcal{\kappa}^{3}}\sin(\alpha)\cos(\alpha).$
In this bound, the left hand side is the square of the smallest eigenvalue of the matrix given in Eq. II.3, while the right hand side is the square of the largest eigenvalue of the matrix in Eq. II.3.
For the choice of $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, one can push the lower bound for the choice of given $\alpha$. This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Fig. 1(a). We show this in Fig. 2, where for a given $\alpha=60^{\circ}$, we obtain
$\rm\bigg{|}\bigg{[}\bigg{(}\sqrt{1-\frac{\gamma^{2}}{\mathcal{\kappa}^{2}}}\bigg{)}sin(\alpha)+\frac{\gamma}{\mathcal{\kappa}}cos(\alpha)\bigg{]}\bigg{|}^{2}$$\leq$ $P^{\prime}(1)$ (23)
$\leq$ $\bigg{|}\bigg{[}(\sqrt{1-{\gamma^{2}}})\mathrm{sin}(\alpha)$
$+$ ${\gamma}\mathrm{cos}(\alpha)\bigg{]}\bigg{|}^{2}.$
While the $A$ and $B$ matrices share the same basis {|$\{\ket{\nu_{i}}\}$, an eigenstate of $B$ matrix that achieves the lower bound for $P^{\prime}(1)$ may not often correspond to the largest/smallest eigenvalue of $A$ matrix, for a given choice of angle $\alpha$.
Note that for an arbitrary input $|b\rangle$ which is a nontrivial superposition of all the eigenvectors $\{|\nu_{i}\rangle\}$ of $A$, we have the following condition,
${P^{\prime}(1)}$ $=$ $\left(\sum_{i}\left|b_{i}\left(\sqrt{1-\frac{C^{2}}{{\lambda}_{i}^{2}}}\sin({\alpha})+\frac{C}{{\lambda}_{i}}\cos({\alpha})\right)\right|^{2}\right)$ (24)
$=$ $\left(\sum_{i}\mid b_{i}\mid^{2}{\lambda}_{i}^{\prime 2}\right).$
Here, $\{\lambda_{i}^{\prime}\}$ are the eigenvalues of the operator $\rm\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)+CA^{-1}\cos(\alpha)$. Eq. 24 is a convex combination of the square of eigenvalues, $\lambda_{i}^{\prime 2}$, therefore, an arbitrary input $|b\rangle$ will yield a $P^{\prime}(1)$ satisfying the condition given in Eq. 22. We augment our findings with the simulations in the subsequent Sec. III.
III Results and discussions
In this section, we consider two sets of examples to compare the performance of Psi-HHL with that of HHL:
  • Toy $A$ matrices of sizes $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$: For practical reasons associated with computational resources, we do not choose matrices whose $\mathcal{\kappa}$ grows with the $A$ matrix size. Instead, we generate a sequence of $A$ matrices with systematically increasing condition numbers for every system size. For example, for $4\times 4$ matrix size, we begin with a sparse and diagonally dominant matrix (the latter is not a requirement but simplifies matters for our illustrations) such that its smallest diagonal entry is $2^{-(n_{r}-1)}$ and the largest 1. Recall that $n_{r}$ is the number of QPE clock register qubits. We generate the sequence by varying $n_{r}$ from 2 to 20. In such a sequence, $\mathcal{\kappa}$ grows exponentially with $n_{r}$, thus enabling us to reach the failure point of HHL rather rapidly and with existing computational resources. We also note that changing the position of the minimum diagonal element of $A$ along the diagonal does not affect our analysis.
  • Quantum chemistry: Since quantum chemistry is one of the main applications of interest with quantum computers, it is worth exploring this domain for our analysis [2]. However, since we no longer are considering toy examples, we begin by qualitatively assessing the applicability of HHL to the LCC problem by:
    1. 1.
      Checking how the $A$ matrices grow with system size (and along the way, defining the system size), thereby providing a handle on the classical hardness of the problem, and
    2. 2.
      Checking how $\mathcal{\kappa}$ itself grows as a function of what we set to be the system size. Only after this qualitative analysis do we move to our main results for this problem domain.
  • We carry out the following analyses:
    1. 1.
      For the toy matrices, we compare HHL and Psi-HHL, and demonstrate the ability of the latter in predicting correlation energies across a range of condition number values. As we shall explain later in this section, we find that since $\mathcal{\kappa}$ grows slowly for the quantum chemistry case, it is not computationally feasible to carry out the analysis. We quantify the quality of our results for the correlation energies through percentage fraction difference, abbreviated as PFD, given by $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{HHL}}}{E_{\mathrm{c}}}\times 100$ for HHL and $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{Psi-HHL}}}{E_{\mathrm{c}}}\times 100$ for Psi-HHL. We denote the correlation energy obtained on a traditional computer by $E_{\mathrm{c}}$, whereas $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the values for the quantity obtained from HHL and Psi-HHL respectively.
    2. 2.
      For both the toy matrices as well as the quantum chemistry matrices, we study the PFDs of HHL and Psi-HHL approaches versus the number of shots.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Refer to caption Refer to caption
(d) (e)
Figure 3: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for toy matrices. Sub-figures whose top panels present the average probabilities of HHL and Psi-HHL with varying values of $\mathcal{\kappa}$ for different $A$ matrix sizes: $4\times 4$ (sub-figure (a)), $8\times 8$ (sub-figure (b)), $16\times 16$ (sub-figure (c)), $32\times 32$ (sub-figure (d)), and $64\times 64$ (sub-figure (e)). The bottom panel of each sub-figure represents the average Percentage Fraction Differences (PFD) of the overlap, $E_{c}$, calculated with HHL and Psi-HHL with respect to the classical value versus $\mathcal{\kappa}$. Each data point results from 10 repeated runs of $10^{6}$ shots each. Also, all the computations have been done with $A$ not diagonal and $\vec{b}$ in unequal superposition, as described in the main text. We also note here that the approximate equality is due to the fit coefficients in the figure being rounded off to two decimal places.
III.1 Toy matrices: $4\times 4$ through $64\times 64$
We begin by discussing the details of our $A$ matrices and $\vec{b}$ vectors. We pick our $4\times 4$ matrix to be non-diagonal, and is given by
$A=\begin{pmatrix}2^{-(n_{r}-1)}&0.00&0.00&0.00\\ 0.00&0.75&0.10\times 10^{-3}&0.00\\ 0.00&0.10\times 10^{-3}&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix},$
and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is defined as the number of non-zero entries in the row with the most number of non-zero elements and is set to 2. Our choices for the other matrices, $8\times 8$ through $64\times 64$ can be found in Figure A1 of the Appendix. We have set the following values for sparsity: 3 for $8\times 8$, 4 for $16\times 16$, 17 for $32\times 32$, and 7 for $64\times 64$ matrices. Throughout, we have ensured that the eigenvalues of the $A$ matrix lies between $1/\mathcal{\kappa}$ and 1. We provide more examples for the $4\times 4$ matrix size in Section A1 of the Appendix.
Figures 3(a) through 3(e) provides our findings for $4\times 4$ through the $64\times 64$ examples considered. Each data point in the sub-figures is the average result obtained over 10 repetitions, with each repetition involving 1 million shots. We pick $\alpha=60^{\circ}$ for the Psi-HHL calculations. The corresponding data is given in Table A1 of the Appendix. Our results obtained using other choices of angles, $70^{\circ}$ and $80^{\circ}$, are provided in Tables A2 and A3 respectively. We now return to Figure 3, where each sub-figure provides two sets of data: one which compares the probabilities, $P(0)$, $P(1)$, $P^{\prime}(0)$, and $P^{\prime}(1)$, obtained by executing HHL and Psi-HHL, and the other providing the PFDs for the two methods. We choose condition numbers all the way till about 1 million. Figures A2 and A3 of the Appendix shows these results for $\alpha$ of $60^{\circ}$, $70^{\circ}$, and $80^{\circ}$, with $10^{5}$ and $10^{6}$ shots respectively. The figures show that our data fits for HHL align with the expected behaviour of $P(1)\sim\mathcal{O}(\mathcal{\kappa}^{-2})$. We also find that $P^{\prime}(1)$ scales as $\sim\mathcal{O}(1)$ for large condition numbers (with the fit parameter $a=0.75$ matching with the Y-axis intercept for large condition number values). For completeness, we verify that the fit probabilities add to one (see Table A4 of the Appendix).
Refer to caption Refer to caption
(a) (b)
Figure 4: Figure showing a comparison of the spread in PFDs as well as the average PFD (over 10 repetitions), obtained using HHL (sub-figure (a)) and Psi-HHL (sub-figure (b)), against a range of three different numbers of shots and for the case where $A$ is of size $4\times 4$, and for which $\mathcal{\kappa}\approx 2^{12}=4096$.
We now comment on the PFDs. We immediately see that HHL fails at some value of $\mathcal{\kappa}$ for the chosen number of shots (indicated by ‘nan’ values in Tables A1, A2, and A3 of the Appendix). This is due to the fact that as $P(1)$ becomes very small, the fraction of the total number of shots that yield 1 upon measurement tend to very small values. Therefore, the number of shots available for the HOM module become too small to give any reasonable result beyond a certain value of $\mathcal{\kappa}$. The bottom panel of each sub-figure of Figure 3 shows that the error bars, in which the upper and lower points of the bars reflect the largest and the smallest PFD values obtained using HHL, increases as the condition number increases, till it reaches a critical value of $\mathcal{\kappa}$, beyond which HHL fails. In contrast, we see that for Psi-HHL, even at $\mathcal{\kappa}\ \approx$ 1 million, the length of the error bar is about 17 percent.
We now verify that HHL does not fail upon increasing the number of shots, but the number requires to be about 1 billion in order to get a comparable range of PFDs as Psi-HHL, while Psi-HHL only incurs 10 million shots, as Figure 4 shows. In particular, HHL gives a maximum PFD of 1.51 percent and a minimum of $-2.98$ percent with a billion shots, whereas Psi-HHL gives a maximum and minimum PFD of 2.33 and $-2.97$ percent respectively with only 10 million shots.
We finally comment on our results from Figure 3 on the larger size matrices ($8\times 8$ through $64\times 64$), where for Psi-HHL, we consider $\alpha=60^{\circ}$. Figures A4 through A7 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A4 also giving the information for $10^{5}$ shots. We observe that broadly, the trends are similar to what we find for the smaller $4\times 4$ cases. Figure A8 of the Appendix summarizes the situations where HHL fails for all of the toy matrices considered, as a heat map. We find that for 1 million shots, HHL starts returning ‘nan’ values between $\mathcal{\kappa}$ of $2^{8}$ and $2^{10}$. In view of the computational cost involved, we did not go all the way up to a condition number of $2^{20}$ for matrices above $4\times 4$ size, and we stopped much earlier for the $64\times 64$ case.
III.2 Quantum chemistry matrices
Quantum chemistry is regarded as one of the killer applications of quantum computing [21]. Therefore, it is interesting to consider the performance of Psi-HHL in this scenario. In particular, we consider quantum chemistry problems via the LCC approach [2], for this purpose. Quantum chemistry on quantum computers can involve different approaches to selecting systems of interest depending on the application domain, including but not limited to performing calculations on numerous candidate molecules of progressively increasing size, picking one molecule and systematically improving the precision in one’s calculations via suitable active space choices (as well as better quality single particle bases, incorporating more physical effects in the molecular non-relativistic Hamiltonian, etc), etc. For the purposes of this work, since we only want to compare performance of the HHL and Psi-HHL approaches, we opt for the substantially simpler second approach.
Refer to caption Refer to caption Refer to caption
(a) (b) (c)
Figure 5: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for three representative molecules. PFDs of Li2 (sub-figure (a)), RbH (sub-figure (b)), and CsH (sub-figure (c)) molecules versus number of shots (denoted in the figure as ‘No. of shots’). The centre of each unfilled circle represents the average PFD over a range of 5 repetitions. Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$, while for RbH, we evaluate a 25-qubit circuit involving an $A$ matrix size of $256\times 256$. For CsH, we carry out 26-qubit calculations where the $A$ matrix size is $256\times 256$. The zero PFD value is marked with a horizontal line for ease of visualization.
III.2.1 The $A$ matrix dimension
Since we are now going beyond toy examples, we begin with some preliminary qualitative remarks. We begin with the growth of the LCC $A_{mol}$ matrix. We make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$, where we recall that $n=2^{N}$. We go from $A_{mol}$ to $A$ via padding, which we explain in Section A2 of the Appendix.
  • The number of rows (or columns), $m$, of the $A_{mol}$ matrix, which we shall hereafter think of as the system size, is the number of particle-hole excitations that one considers for a calculation, which in turn is decided by choice of single particle basis and active space size. We denote the number of occupied orbitals as $n_{o}$ and virtuals by $n_{v}$. It is common to employ the LCC approach by only considering single and double excitations (termed LCCSD), in which case $A_{mol}$ matrix grows as $m\sim n_{o}^{2}n_{v}^{2}$ for fixed values of $n_{o}$ and $n_{v}$. As we include more excitations systematically, such as triples, quadruples, etc, while keeping the number of orbitals fixed, we observe that the $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$, which is exponential. $\mathcal{E}$ refers to the excitation rank.
  • For a given excitation rank, $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially, for example, see Figure A9 in the Appendix. For the purposes of this work, we restrict ourselves to LCCSD approximation ($\mathcal{E}=2$), and study the size of $A_{mol}$ with increase in number of unoccupied orbitals, $n_{v}$, in a molecule, while keeping $n_{o}$ fixed at a reasonable number.
For our analyses, we do not attempt to consider a molecule in a given active space size to study the variation of $m$ with choice of single particle basis, and defer the study for a future work. We choose three examples: the light Li2, the moderately heavy RbH, and the heavy CsH neutral diatomic molecules, all in their equilibrium bond lengths (Li2: 5.06 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr)  [22, 23]. The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule. We pick the Sapporo double zeta basis sets for RbH and CsH (Rb: Sapporo-DZP, Cs: Sapporo-DKH3-DZP-2012, and H: Sapporo-DZP) [24], while for Li2, we choose the STO-6G basis [25]. We employ the $C_{2v}$ point group symmetry for RbH and CsH computations, while for the case of Li2, we pick the $D_{2h}$ symmetry. We fix $n_{o}$ to 3, 13 , and 18 for Li2, RbH, and CsH respectively. The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector are discussed in Section A2.A of the Appendix.
III.2.2 Condition number of the $A$ matrix
Although HHL offers an exponential advantage in terms of system size, it is worth checking if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the systems that we considered. We point out towards the end of Section A2.A of the Appendix that it is sufficient to analyze $\mathcal{\kappa}$ versus $m$; this curve upper bounds $\mathcal{\kappa}$ versus $n$. We increase the active space, and thus $m$, in a systematic manner by carrying out an orbital character analysis, whose details we explain in Section A2.B of the Appendix. Furthermore, for this analysis, we only consider RbH and CsH, and not Li2 as the system is too small (in the STO-6G basis) to carry out condition number versus system size analysis. For RbH, we consider $n_{o}=13$ for our calculations, and vary the number of virtuals from 1 to 21, which corresponds to $A_{mol}$ matrix size of $40\times 40$ to $9065\times 9065$ respectively. For CsH, $n_{o}=18$ and virtuals are varied from 1 to 16, while the $A_{mol}$ matrix size goes from $67\times 67$ to $11058\times 11058$.
Our results shown in Figure A10 of the Appendix indicate that for RbH as well as for CsH, as we go from $m$ of 40 to 9065 and 67 to 11059 respectively, $\mathcal{\kappa}$ grows polylogarithmically (fourth and sixth order respectively). This behaviour is also consistent with what one may expect from chemistry intuition: we begin by referring the reader to Table A10 of the Appendix, where we list the largest and smallest eigenvalues of the $A$ matrix for each value of $m$, for both RbH and CsH. The data shows that the increase in $\mathcal{\kappa}$ as we add virtual orbitals is driven by the decrease of the lowest eigenvalue of the $A$ matrix. We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock configuration. The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the HF configuration). Thus, we expect that the decrease in the lowest eigenvalue of $A$ saturates as we keep adding virtual orbitals, and eventually, the lowest eigenvalue of the $A$ matrix will converge to a certain value. Thus, the increase of $\mathcal{\kappa}$ saturates too, thus leading one to expect that the condition number would vary as a polylogarithmic function of system size. We add at this juncture that while considering only two molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may a reasonable direction to consider. However, further analysis would be required to assess what one would expect for $\mathcal{\kappa}(m)$ in general for an HHL-LCC computation. We also note that we carried out a similar analysis for sparsity, and found that it too grows polylogarithmically for RbH and CsH (see Figure A11 of the Appendix).
III.2.3 HHL and Psi-HHL results
We now discuss our results for the performance of the HHL and Psi-HHL algorithms for the molecules considered in this work.
We begin by conducting a study similar to the one shown in Figure 4 for the toy matrices. It is not practical to carry out an analysis similar to that in Figure 3, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon. For the purposes of our analysis in Figure 4, we vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 5 repetitions. For this analysis, we consider Li2, RbH and CsH. The details are:
  1. 1.
    Li2: 21-qubit circuit with an $A$ matrix size of $64\times 64$ (built out of single and double excitations arising from 3 occupied and 8 virtual orbitals) in the STO-6G basis, with $\mathcal{\kappa}=64.07$.
  2. 2.
    RbH: 25-qubit circuit with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 13 occupied and 3 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=153.17$.
  3. 3.
    CsH: 26-qubit calculations with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 18 occupied and 2 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=466.59$. Although RbH and CsH have the same $A$ matrix size, the latter involves a 26-qubit calculation because we set an additional ancilla for CsH due to its larger atomic number.
We note that for RbH, the same calculation would have incurred more qubits for VQE-UCCSD (32-qubit circuit) and QPE-CASCI (with $n_{r}=8$, 40-qubit circuit) computations, while we carry out a 25-qubit circuit evaluation in the case of HHL-LCC. Similarly, for CsH, VQE-UCCSD: 40 qubits, QPE-CASCI: 48 qubits, and HHL-LCC: 26 qubits.
Our results are presented in Figure 5, with the accompanying data given in Table A11 of the Appendix. For Li2, the condition number is very small, and hence we only study the performance of HHL and Psi-HHL between 500 and 10000 shots. The data indicates that for Li2, the Psi-HHL approach gives better average PFDs, especially in the regime of fewer number of shots (500 and 1000 shots). In the same regime, the spread in PFDs is also lesser for Psi-HHL. For the cases of RbH and CsH, as we increase the number of shots, the error bars for both HHL and Psi-HHL are seen to become much smaller.
IV Conclusion
We propose the Psi-HHL algorithm to efficiently handle problems involving $A$ matrices with large condition numbers. The approach involves post-selecting 0 in HHL (the ‘wrong signal’), followed by post-selecting 1 from an HHL-like circuit that contains an additional mixing $2\times 2$ unitary (the ‘mixed signal’), and taking the difference between the wrong and the mixed signals to extract the correct signal.
We present arguments to show that the complexity of Psi-HHL is optimal in condition number for large $\mathcal{\kappa}$ regimes; the complexity of the algorithm scales as $\mathcal{\kappa}$. This is an improvement over earlier approaches in literature, such as HHL with amplitude amplification (scales quadratically, and incurs several additional gates due to its iterative nature) and HHL with variable time amplitude amplification (scales as $\mathcal{\kappa}\mathrm{log}^{3}\mathcal{\kappa}$). It is also important to note that while the optimal $\mathcal{\kappa}$ behaviour can be recovered using a discrete quantum adiabatic approach based quantum linear solver algorithm [12], we achieve the same result using Psi-HHL, which is a simple modification to the HHL algorithm.
We apply our method to different toy matrices, where we systematically vary the condition number all the way to about 1 million while keeping the number of shots fixed to 1 million, and show that while HHL fails beyond a certain value, Psi-HHL continues to predict good results. We also demonstrate the performance of our proposed Psi-HHL approach for quantum chemistry examples, where we once again find that Psi-HHL yields better quality results in much fewer number of shots. En route, we also carry out a qualitative analysis to check suitability of treating LCC equations using the HHL algorithm.
We anticipate that our work will pave the way for development of non-trivial Psi-HHL variants that would allow it to be integrated as an algorithmic primitive in larger algorithmic frameworks that use traditional HHL as a subroutine.
Acknowledgements.
We acknowledge National Supercomputing Mission (NSM) for providing computing resources of ‘PARAM Siddhi-AI’, under National PARAM Supercomputing Facility (NPSF), C-DAC, Pune and supported by the Ministry of Electronics and Information Technology (MeitY) and Department of Science and Technology (DST), Government of India. VSP thanks Prof. Aram Harrow for a very fruitful discussion on HHL as well as HHL-LCC. VSP and AJ thank Dr. Shibdas Roy for nice discussions on complexity. VSP and PBT thank Prof. Debashis Mukherjee for useful discussions on the quantum chemical sections of the work. KS acknowledges support from Quantum Leap Fragship Program (Grant No. JPMXS0120319794) from Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan; Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from Japan Science and Technology Agency (JST), Japan; and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from Japan Society for the Promotion of Sciences (JSPS), Japan.
Appendix
Appendix A1 Details on calculations involving toy matrices from size $4\times 4$ through $64\times 64$
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure A1: Our choices for (a) $8\times 8$, (b) $16\times 16$, (c) $32\times 32$, and (d) $64\times 64$ toy matrices for comparing the performance of HHL and the Psi-HHL algorithms. The cell with a star symbol denotes the matrix element with the variable, $n_{r}$ (the number of clock register qubits from the HHL circuit), which sets $\mathcal{\kappa}$.
In our Results section in the main text, we discussed the case of $A$ not diagonal and the amplitude encoded state corresponding to $\vec{b}$ in an unequal superposition. We also provide our $4\times 4$ through $64\times 64$ matrices that we considered in the main text in Figure A1.
We now provide additional data on two more choices for our $4\times 4$ $A$ matrices:
  1. 1.
    $A$ is diagonal and the amplitude encoded state corresponding to $\vec{b}$ is in an equal superposition. In particular, $A=\begin{pmatrix}2^{-(n_{r}-1)}&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, $\vec{b}=\begin{pmatrix}1.00\\ 1.00\\ 1.00\\ 1.00\end{pmatrix}$. The sparsity, $s$, of $A$ is set to 1.
  2. 2.
    $A$ is diagonal and the amplitude encoded state associated with $\vec{b}$ is in an unequal superposition. In particular, $A$ is the same as in case 1 given above, and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$.
We also present these results for $\alpha$ set to 60, 70, and 80 degrees (the main text only discussed the first case). Furthermore, while the main text considers 1 million shots, we also perform the same analyses for $10^{5}$ shots. All of these results are presented in Figures A2 and A3 for the cases considering $10^{5}$ and $10^{6}$ shots respectively. The accompanying data is provided in Tables A1, A2, and A3, for 60, 70, and 80 degrees examples respectively.
Refer to caption Refer to caption Refer to caption
(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
Refer to caption Refer to caption Refer to caption
(d) $\alpha=60^{\circ}$ (e) $\alpha=70^{\circ}$ (f) $\alpha=80^{\circ}$
Refer to caption Refer to caption Refer to caption
(g) $\alpha=60^{\circ}$ (h) $\alpha=70^{\circ}$ (i) $\alpha=80^{\circ}$
Figure A2: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{5}$ shots throughout. Sub-figures (a) through (c) consider the case where $A$ is diagonal and $\vec{b}$ is in equal superposition. Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Subfigures (g) through (i) consider the case where $A$ is not diagonal and $\vec{b}$ is in unequal superposition.
Refer to caption Refer to caption Refer to caption
(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
Refer to caption Refer to caption Refer to caption
(d) $\alpha=60^{\circ}$ (e) $\alpha=70^{\circ}$ (f) $\alpha=80^{\circ}$
Refer to caption Refer to caption Refer to caption
(g) $\alpha=60^{\circ}$ (h) $\alpha=70^{\circ}$ (i) $\alpha=80^{\circ}$
Figure A3: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{6}$ throughout. Sub-figures (a) through (c) consider the case where $A$ is diagonal, and $\vec{b}$ is in equal superposition. Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal, and $\vec{b}$ is in unequal superposition.
Table A1: Table presenting the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$) for $4\times 4$ matrices, and in particular, for three cases: $A$ diagonal and $\vec{b}$ in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ equal’), $A$ diagonal and $\vec{b}$ not in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ unequal’), and $A$ not diagonal and $\vec{b}$ in an unequal superposition (denoted in the table as ‘$A$ not diag, $\vec{b}$ unequal’). The string ‘nan’ refers to situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result.
$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
[…]
$2^{2}$ 8 -0.44(0.65, 0.35) -0.22(0.21, 0.79) -1.01(0.92, 0.08) -0.10(0.06, 0.94) -0.87(0.92, 0.08) -0.04(0.08, 0.92)
$2^{3}$ 9 -0.55(0.72, 0.28) -0.07(0.27, 0.73) -5.82(0.97, 0.03) -0.32(0.15, 0.85) -5.72(0.97, 0.03) 0.08(0.15, 0.85)
$2^{4}$ 10 -0.81(0.74, 0.26) 0.32(0.32, 0.68) 14.39(0.99, 0.01) -0.69(0.2, 0.80) 17.20(0.99, 0.01) 0.26(0.2, 0.80)
$2^{5}$ 11 -0.17(0.75, 0.25) -0.02(0.35, 0.65) 3.96(0.99, 0.01) 2.85(0.23, 0.77) -0.12(0.99, 0.01) 0.32(0.22, 0.78)
$2^{6}$ 12 0.30(0.75, 0.25) 0.05(0.37, 0.63) -6.73(0.99, 0.01) 1.66(0.24, 0.76) -6.73(0.99, 0.01) -1.56(0.24, 0.76)
$2^{7}$ 13 nan(0.75, 0.25) -0.45(0.37, 0.63) nan(0.99, 0.01) 1.91(0.25, 0.75) nan(0.99, 0.01) -2.62(0.25, 0.75)
$2^{8}$ 14 nan(0.75, 0.25) -0.01(0.37, 0.63) nan(0.99, 0.01) -2.00(0.25, 0.75) nan(0.99, 0.01) -3.87(0.25, 0.75)
$2^{9}$ 15 nan(0.75, 0.25) -0.22(0.37, 0.63) nan(0.99, 0.01) -7.50(0.25, 0.75) nan(0.99, 0.01) 8.02(0.25, 0.75)
$2^{10}$ 16 nan(0.75, 0.25) -0.26(0.37, 0.63) nan(0.99, 0.01) -2.33(0.25, 0.75) nan(0.99, 0.01) -7.77(0.25, 0.75)
$2^{11}$ 17 nan(0.75, 0.25) 0.08(0.37, 0.63) nan(0.99, 0.01) 2.65(0.25, 0.75) nan(0.99, 0.01) -0.55(0.25, 0.75)
$2^{12}$ 18 nan(0.75, 0.25) 0.03(0.37, 0.63) nan(0.99, 0.01) 7.93(0.25, 0.75) nan(0.99, 0.01) -2.41(0.25, 0.75)
$2^{13}$ 19 nan(0.75, 0.25) 0.30(0.38, 0.62) nan(0.99, 0.01) 4.67(0.25, 0.75) nan(0.99, 0.01) 3.12(0.25, 0.75)
$2^{14}$ 20 nan(0.75, 0.25) 0.41(0.38, 0.62) nan(0.99, 0.01) 9.11(0.25, 0.75) nan(0.99, 0.01) -5.60(0.25, 0.75)
$2^{15}$ 21 nan(0.75, 0.25) 0.13(0.38, 0.62) nan(0.99, 0.01) 9.42(0.26,0.74) nan(0.99, 0.01) -1.11(0.25, 0.75)
$2^{16}$ 22 nan(0.75, 0.25) 0.54(0.38, 0.62) nan(0.99, 0.01) -0.25(0.25, 0.75) nan(0.99, 0.01) 2.43(0.25, 0.75)
$2^{17}$ 23 nan(0.75, 0.25) -0.03(0.38, 0.62) nan(0.99, 0.01) -0.40(0.26, 0.74) nan(0.99, 0.01) -7.87(0.25, 0.75)
$2^{18}$ 24 nan(0.75, 0.25) 0.26(0.37, 0.63) nan(0.99, 0.01) -8.56(0.26, 0.74) nan(0.99, 0.01) 9.69(0.26, 0.74)
$2^{19}$ 25 nan(0.75, 0.25) 0.19(0.37, 0.63) nan(0.99, 0.01) 6.91(0.26, 0.74) nan(0.99, 0.01) 1.06(0.25, 0.75)
$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -0.58(0.37, 0.63) nan(0.99, 0.01) 0.30(0.25, 0.75) nan(0.99, 0.01) 6.15(0.25, 0.75)
$2^{2}$ 8 -0.57(0.65, 0.35) 0.00(0.21, 0.79) -0.87(0.92, 0.08) 0.03(0.08, 0.92) -0.88(0.92, 0.08) 0.04(0.08, 0.92)
$2^{3}$ 9 -0.61(0.72, 0.28) -0.14(0.27, 0.73) -5.82(0.97, 0.03) -0.04(0.15, 0.85) -5.82(0.97, 0.03) 0.00(0.15, 0.85)
$2^{4}$ 10 -0.63(0.74, 0.26) -0.04(0.31, 0.69) 14.66(0.99, 0.01) 0.11(0.20, 0.8) 14.95(0.99, 0.01) -0.15(0.20, 0.80)
$2^{5}$ 11 0.27(0.75, 0.25) -0.06(0.34, 0.66) 1.84(0.99, 0.01) 0.16(0.22, 0.78) 0.61(0.99, 0.01) 0.44(0.22, 0.78)
$2^{6}$ 12 0.04(0.75, 0.25) 0.00(0.36, 0.64) -0.72(0.99, 0.01) 0.01(0.24, 0.76) 1.68(0.99, 0.01) -0.70(0.24, 0.76)
$2^{7}$ 13 -0.07(0.75, 0.25) 0.20(0.37, 0.63) 1.35(0.99, 0.01) -1.56(0.25, 0.75) -4.87(0.99, 0.01) -0.11(0.25, 0.75)
$2^{8}$ 14 -0.13(0.75, 0.25) -0.02(0.37, 0.63) 0.09(0.99, 0.01) -0.04(0.25, 0.75) -4.04(0.99, 0.01) -1.15(0.25, 0.75)
$2^{9}$ 15 0.05(0.75, 0.25) 0.06(0.37, 0.63) -2.82(0.99, 0.01) 2.82(0.25, 0.75) nan(0.99, 0.01) 2.26(0.25, 0.75)
$2^{10}$ 16 0.04(0.75, 0.25) -0.04(0.37, 0.63) nan(0.99, 0.01) -0.12(0.25, 0.75) nan(0.99, 0.01) -0.59(0.25, 0.75)
$2^{11}$ 17 0.19(0.75, 0.25) -0.05(0.37, 0.63) nan(0.99, 0.01) -1.49(0.25, 0.75) nan(0.99, 0.01) 1.41(0.25, 0.75)
$2^{12}$ 18 0.08(0.75, 0.25) -0.20(0.37, 0.63) nan(0.99, 0.01) 0.11(0.25, 0.75) nan(0.99, 0.01) 1.26(0.25, 0.75)
$2^{13}$ 19 0.12(0.75, 0.25) 0.01(0.37, 0.63) nan(0.99, 0.01) 0.13(0.25, 0.75) nan(0.99, 0.01) -3.37(0.25, 0.75)
$2^{14}$ 20 -0.03(0.75, 0.25) -0.00(0.37, 0.63) nan(0.99, 0.01) -4.52(0.25, 0.75) nan(0.99, 0.01) 1.47(0.25, 0.75)
$2^{15}$ 21 0.05(0.75, 0.25) -0.05(0.37, 0.63) nan(0.99, 0.01) -0.70(0.25, 0.75) nan(0.99, 0.01) -0.79(0.25, 0.75)
$2^{16}$ 22 0.10(0.75, 0.25) 0.12(0.38, 0.63) nan(0.99, 0.01) 1.57(0.25, 0.75) nan(0.99, 0.01) 0.84(0.25, 0.75)
$2^{17}$ 23 -0.03(0.75, 0.25) -0.15(0.38, 0.63) nan(0.99, 0.01) -0.40(0.25, 0.75) nan(0.99, 0.01) 0.05(0.25, 0.75)
$2^{18}$ 24 0.11(0.25, 0.75) 0.00(0.37, 0.63) nan(0.99, 0.01) -1.22(0.25, 0.75) nan(0.99, 0.01) -3.58(0.25, 0.75)
$2^{19}$ 25 0.08(0.75, 0.25) -0.03(0.37, 0.63) nan(0.99, 0.01) 0.78(0.25, 0.75) nan(0.99, 0.01) 1.41(0.25, 0.75)
$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.19(0.37, 0.63) nan(0.99, 0.01) -1.52(0.25, 0.75) nan(0.99, 0.01) 0.01(0.25, 0.75)
Table A2: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=70^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Table A1.
$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 8 -0.04(0.65, 0.35) -0.24(0.23, 0.77) -1.01(0.92, 0.08) -0.18(0.02, 0.98) -0.87(0.92, 0.08) 0.58(0.02, 0.98)
$2^{3}$ 9 -0.55(0.72, 0.28) -0.28(0.25, 0.75) -5.82(0.97, 0.03) -0.50(0.06, 0.94) -5.72(0.97, 0.03) -0.41(0.06, 0.94)
$2^{4}$ 10 -0.81(0.74, 0.26) 0.51(0.27, 0.73) 14.39(0.99, 0.01) 0.33(0.12, 0.88) 17.20(0.99, 0.01) 0.44(0.09, 0.91)
$2^{5}$ 11 -0.17(0.75, 0.25) 0.05(0.29, 0.71) 3.96(0.99, 0.01) -0.17(0.1, 0.90) -0.12(0.99, 0.01) 1.17(0.10, 0.90)
$2^{6}$ 12 0.30(0.75, 0.25) 0.54(0.30, 0.70) nan(0.99, 0.01) -1.16( 0.12, 0.88) -6.73(0.99, 0.01) 0.99(0.11, 0.89)
$2^{7}$ 13 nan(0.75, 0.25) 0.74(0.30,0.70) nan(0.99, 0.01) 0.30(0.12, 0.88) nan(0.99, 0.01) 2.20(0.02, 0.88)
$2^{8}$ 14 nan(0.75, 0.25) 0.40(0.30, 0.69) nan(0.99, 0.01) -7.03(0.12, 0.88) nan(0.99, 0.01) 3.54(0.02, 0.88)
$2^{9}$ 15 nan(0.75, 0.25) -0.31(0.31, 0.69) nan(0.99, 0.01) 2.88(0.12, 0.88) nan(0.99, 0.01) 1.96(0.02, 0.88)
$2^{10}$ 16 nan(0.75, 0.25) 0.02(0.31, 0.69) nan(0.99, 0.01) 5.28(0.12, 0.88) nan(0.99, 0.01) 4.93(0.02, 0.88)
$2^{11}$ 17 nan(0.75, 0.25) -0.30(0.31, 0.69) nan(0.99, 0.01) 2.52(0.12, 0.88) nan(0.99, 0.01) -2.36(0.02, 0.88)
$2^{12}$ 18 nan(0.75, 0.25) 0.48(0.31, 0.69) nan(0.99, 0.01) -3.50(0.12, 0.88) nan(0.99, 0.01) 0.81(0.02, 0.88)
$2^{13}$ 19 nan(0.75, 0.25) 0.38(0.31, 0.69) nan(0.99, 0.01) -9.08(0.12, 0.88) nan(0.99, 0.01) -2.22(0.02, 0.88)
$2^{14}$ 20 nan(0.75, 0.25) 0.15(0.31, 0.69) nan(0.99, 0.01) -0.46(0.12, 0.88) nan(0.99, 0.01) 0.33(0.02, 0.88)
$2^{15}$ 21 nan(0.75, 0.25) 1.26(0.31, 0.69) nan(0.99, 0.01) 3.66(0.12, 0.88) nan(0.99, 0.01) 6.98(0.02, 0.88)
$2^{16}$ 22 nan(0.75, 0.25) -0.34(0.31, 0.69) nan(0.99, 0.01) 5.62(0.12, 0.88) nan(0.99, 0.01) -17.07(0.02, 0.88)
$2^{17}$ 23 nan(0.75, 0.25) 1.31(0.31, 0.69) nan(0.99, 0.01) 5.56(0.12, 0.88) nan(0.99, 0.01) 2.02(0.02, 0.88)
$2^{18}$ 24 nan(0.75, 0.25) 0.00(0.69, 0.31) nan(0.99, 0.01) 0.00(0.12, 0.88) nan(0.99, 0.01) 18.88(0.13, 0.87)
$2^{19}$ 25 nan(0.75, 0.25) 0.33(0.31, 0.69) nan(0.99, 0.01) 0.33(0.12, 0.88) nan(0.99, 0.01) 6.42( 0.12, 0.88)
$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -0.27(0.31, 0.69) nan(0.99, 0.01) -6.81(0.12, 0.88) nan(0.99, 0.01) 6.75(0.12, 0.88)
$2^{2}$ 8 -0.57(0.64, 0.36) -0.01(0.23, 0.77) -0.87(0.92, 0.08) -0.02(0.02, 0.98) -0.88(0.92, 0.08) -0.01(0.02, 0.98)
$2^{3}$ 9 -0.61(0.72, 0.28) -0.16(0.22, 0.76) -5.82(0.97, 0.03) 0.04(0.06,0.94) -5.82(0.97, 0.03) 0.00(0.06, 0.94)
$2^{4}$ 10 -0.63(0.74, 0.26) -0.05(0.27, 0.73) 14.66(0.99, 0.01) -0.26(0.09, 0.91) 14.95(0.99, 0.01) -0.27(0.09, 0.91)
$2^{5}$ 11 0.27(0.75, 0.25) -0.08(0.29, 0.71) 1.84(0.99, 0.01) -0.26(0.10, 0.90) 0.61(0.99, 0.01) 0.50(0.11, 0.89)
$2^{6}$ 12 0.04(0.75, 0.25) 0.02(0.3, 0.70) -0.72(0.99, 0.01) 1.32(0.11, 0.89) 1.68(0.99, 0.01) -0.27(0.12, 0.88)
$2^{7}$ 13 -0.07(0.75, 0.25) -0.07(0.31, 0.69) 1.35(0.99, 0.01) -0.24(0.12, 0.88) -4.87(0.99, 0.01) 0.40(0.12, 0.88)
$2^{8}$ 14 -0.13(0.75, 0.25) 0.12(0.31, 0.69) 0.09(0.99, 0.01) 0.43(0.12, 0.88) -4.04(0.99, 0.01) 1.91(0.12, 0.88)
$2^{9}$ 15 0.05(0.75, 0.25) 0.19(0.31, 0.69) -2.82(0.99, 0.01) -1.81(0.12, 0.88) nan(0.99, 0.01) 1.12(0.12, 0.88)
$2^{10}$ 16 -0.04(0.75, 0.25) -0.26(0.31, 0.69) nan(0.99, 0.01) 1.19(0.12, 0.88) nan(0.99, 0.01) -2.59(0.12, 0.88)
$2^{11}$ 17 0.18(0.75, 0.25) 0.04(0.31, 0.69) nan(0.99, 0.01) -1.68(0.12, 0.88) nan(0.99, 0.01) -0.85(0.12, 0.88)
$2^{12}$ 18 0.07(0.75, 0.25) -0.44(0.31, 0.69) nan(0.99, 0.01) 2.36(0.12, 0.88) nan(0.99, 0.01) -5.96(0.12, 0.88)
$2^{13}$ 19 0.12(0.75, 0.25) 0.15(0.31, 0.69) nan(0.99, 0.01) 5.19(0.12, 0.88) nan(0.99, 0.01) 1.52(0.12, 0.88)
$2^{14}$ 20 -0.04(0.75, 0.25) 0.25(0.31, 0.69) nan(0.99, 0.01) 2.50(0.12, 0.88) nan(0.99, 0.01) -1.89(0.12, 0.88)
$2^{15}$ 21 0.06(0.75, 0.25) -0.33(0.31, 0.69) nan(0.99, 0.01) 0.41(0.12, 0.88) nan(0.99, 0.01) -3.65(0.12, 0.88)
$2^{16}$ 22 0.10(0.75, 0.25) -0.01(0.31, 0.69) nan(0.99, 0.01) -3.09(0.12, 0.88) nan(0.99, 0.01) -1.88(0.12, 0.88)
$2^{17}$ 23 -0.03(0.75, 0.25) -0.07(0.31, 0.69) nan(0.99, 0.01) 2.54(0.12, 0.88) nan(0.99, 0.01) -0.92(0.12, 0.88)
$2^{18}$ 24 0.11(0.75, 0.25) 0.07(0.31, 0.69) nan(0.99, 0.01) -0.47(0.88, 0.12) nan(0.99, 0.01) -0.46(0.88, 0.12)
$2^{19}$ 25 -0.08(0.75, 0.25) 0.09(0.31, 0.69) nan(0.99, 0.01) 1.81(0.12, 0.88) nan(0.99, 0.01) -0.89(0.12, 0.88)
$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.02(0.31, 0.69) nan(0.99, 0.01) -2.00(0.12, 0.88) nan(0.99, 0.01) 3.65(0.12, 0.88)
Table A3: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=80^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Tables A1 and A2.
$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 8 -0.44(0.65, 0.35) 0.43(0.21, 0.79) -1.01(0.92, 0.08) 0.06(0.01, 0.99) -0.87(0.92, 0.08) 0.31(0.02, 0.98)
$2^{3}$ 9 -0.55(0.72, 0.28) 0.29(0.27, 0.73) -5.82(0.97, 0.03) -0.59(0.02, 0.98) -5.72(0.97, 0.03) -0.92(0.02, 0.98)
$2^{4}$ 10 -0.81(0.74, 0.26) 2.47(0.32, 0.68) 14.39(0.99, 0.01) 2.59(0.02, 0.98) 17.20(0.99, 0.01) 1.55(0.02, 0.98)
$2^{5}$ 11 -0.17(0.75, 0.25) 0.67(0.35, 0.65) 3.96(0.99, 0.01) 1.63(0.03, 0.97) -0.12(0.99, 0.01) 1.89(0.03, 0.97)
$2^{6}$ 12 0.30(0.75, 0.25) -0.17(0.36, 0.64) -6.73(0.99, 0.01) 5.34(0.04, 0.96) -6.73(0.99, 0.01) -4.59(0.03, 0.97)
$2^{7}$ 13 nan(0.75, 0.25) -0.30(0.37, 0.63) nan(0.99, 0.01) 0.55(0.04, 0.96) nan(0.99, 0.01) 0.37(0.04, 0.96)
$2^{8}$ 14 nan(0.75, 0.25) -2.17(0.37, 0.63) nan(0.99, 0.01) 6.89(0.04, 0.96) nan(0.99, 0.01) 0.46(0.04, 0.96)
$2^{9}$ 15 nan(0.75, 0.25) 0.04(0.37, 0.63) nan(0.99, 0.01) -2.64(0.04, 0.96) nan(0.99, 0.01) 4.65(0.04, 0.96)
$2^{10}$ 16 nan(0.75, 0.25) 1.01(0.37, 0.63) nan(0.99, 0.01) 3.21(0.04, 0.96) nan(0.99, 0.01) -4.06(0.04, 0.96)
$2^{11}$ 17 nan(0.75, 0.25) -0.20(0.37, 0.63) nan(0.99, 0.01) -5.46(0.04, 0.96) nan(0.99, 0.01) -0.93(0.04, 0.96)
$2^{12}$ 18 nan(0.75, 0.25) -0.55(0.37, 0.63) nan(0.99, 0.01) -4.80(0.04, 0.96) nan(0.99, 0.01) 9.42(0.04, 0.96)
$2^{13}$ 19 nan(0.75, 0.25) -0.17(0.37, 0.63) nan(0.99, 0.01) 2.18(0.04, 0.96) nan(0.99, 0.01) 7.83(0.04, 0.96)
$2^{14}$ 20 nan(0.75, 0.25) 0.56(0.37, 0.63) nan(0.99, 0.01) 5.32(0.04, 0.96) nan(0.99, 0.01) -0.86(0.04, 0.96)
$2^{15}$ 21 nan(0.75, 0.25) -0.11(0.37, 0.63) nan(0.99, 0.01) 5.89(0.04, 0.96) nan(0.99, 0.01) 3.92(0.04, 0.96)
$2^{16}$ 22 nan(0.75, 0.25) 0.24(0.37, 0.63) nan(0.99, 0.01) -12.26(0.04, 0.96) nan(0.99, 0.01) 9.60(0.04, 0.96)
$2^{17}$ 23 nan(0.75, 0.25) -0.94(0.37, 0.63) nan(0.99, 0.01) -2.94(0.04, 0.96) nan(0.99, 0.01) -5.61(0.04, 0.96)
$2^{18}$ 24 nan(0.75, 0.25) -1.64(0.26, 0.74) nan(0.99, 0.01) -0.60(0.04, 0.96) nan(0.99, 0.01) 5.92(0.04, 0.96)
$2^{19}$ 25 nan(0.75, 0.25) 1.47(0.27, 0.73) nan(0.99, 0.01) -0.60(0.04, 0.96) nan(0.99, 0.01) -4.92(0.04, 0.96)
$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -1.70(0.26, 0.74) nan(0.99, 0.01) -17.20(0.04, 0.96) nan(0.99, 0.01) 5.03(0.04, 0.96)
$2^{2}$ 8 -0.57(0.64, 0.36) -0.13(0.28, 0.72) -0.87(0.92, 0.08) 0.00(0.02, 0.98) -0.88(0.92, 0.08) 0.03(0.02, 0.98)
$2^{3}$ 9 -0.61(0.72, 0.28) 0.14(0.24, 0.76) -5.82(0.97, 0.03) -0.19(0.01, 0.99) -5.82(0.97, 0.03) -0.06(0.01, 0.99)
$2^{4}$ 10 -0.63(0.74, 0.26) -0.03(0.25, 0.75) 14.66(0.99, 0.01) -0.22(0.02, 0.98) 14.95(0.99, 0.01) -0.02(0.02, 0.98)
$2^{5}$ 11 0.27(0.75, 0.25) 0.64(0.26, 0.74) 1.84(0.99, 0.01) 0.03(0.03, 0.97) 0.61(0.99, 0.01) 1.11(0.03, 0.97)
$2^{6}$ 12 0.04(0.75, 0.25) -0.05(0.26, 0.74) -0.72(0.99, 0.01) 1.21(0.04, 0.96) 1.68(0.99, 0.01) 0.68(0.03, 0.97)
$2^{7}$ 13 -0.07(0.75, 0.25) 0.17(0.26, 0.74) 1.35(0.99, 0.01) -1.22(0.04, 0.96) -4.87(0.99, 0.01) 0.24(0.03, 0.97)
$2^{8}$ 14 -0.13(0.75, 0.25) -0.17(0.26, 0.74) 0.09(0.99, 0.01) -3.19(0.04, 0.96) -4.04(0.99, 0.01) 0.32(0.04, 0.96)
$2^{9}$ 15 0.05(0.75, 0.25) -0.50(0.26, 0.74) 2.82(0.99, 0.01) -2.13(0.04, 0.96) nan(0.99, 0.01) -2.48(0.04, 0.96)
$2^{10}$ 16 -0.03(0.75, 0.25) -0.30(0.26, 0.74) nan(0.99, 0.01) -0.59(0.04, 0.96) nan(0.99, 0.01) -0.24(0.04, 0.96)
$2^{11}$ 17 0.18(0.75, 0.25) 0.47(0.26, 0.74) nan(0.99, 0.01) -2.93(0.04, 0.96) nan(0.99, 0.01) 0.05(0.04, 0.96)
$2^{12}$ 18 0.07(0.75, 0.25) 0.40(0.26, 0.74) nan(0.99, 0.01) 0.62(0.04, 0.96) nan(0.99, 0.01) -2.73(0.04, 0.96)
$2^{13}$ 19 0.11(0.75, 0.25) 0.14(0.26, 0.74) nan(0.99, 0.01) 0.12(0.04, 0.96) nan(0.99, 0.01) 6.03(0.04, 0.96)
$2^{14}$ 20 -0.03(0.75, 0.25) 0.62(0.26, 0.74) nan(0.99, 0.01) 2.52(0.04, 0.96) nan(0.99, 0.01) 0.28(0.04, 0.96)
$2^{15}$ 21 0.05(0.75, 0.25) -0.34(0.26, 0.74) nan(0.99, 0.01) 1.03(0.04, 0.96) nan(0.99, 0.01) -6.30(0.04, 0.96)
$2^{16}$ 22 0.10(0.75, 0.25) -0.59(0.26, 0.74) nan(0.99, 0.01) -3.21(0.04, 0.96) nan(0.99, 0.01) -2.40(0.04, 0.96)
$2^{17}$ 23 -0.03(0.75, 0.25) 0.49(0.26, 0.74) nan(0.99, 0.01) 0.45(0.04, 0.96) nan(0.99, 0.01) 2.72(0.04, 0.96)
$2^{18}$ 24 0.11(0.75, 0.25) 0.14(0.27, 0.73) nan(0.99, 0.01) -2.46(0.04, 0.96) nan(0.99, 0.01) -1.52(0.04, 0.96)
$2^{19}$ 25 0.08(0.75, 0.25) 0.01(0.27, 0.73) nan(0.99, 0.01) -1.95(0.04, 0.96) nan(0.99, 0.01) -0.72(0.04, 0.96)
$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.33(0.27, 0.73) nan(0.99, 0.01) -2.07(0.04, 0.96) nan(0.99, 0.01) -2.46(0.04, 0.96)
Table A4: Summary of the fitted probabilities of HHL and Psi-HHL for the 4x4 case $A$ not diagonal and $\vec{b}$ in an unequal superposition of states, $\alpha=60^{\circ},70^{\circ},80^{\circ}$ and the number of shots is fixed to $10^{6}$.
$\alpha$ $\approx\mathcal{\kappa}$ $P_{\text{Fit}}(0)$ $P_{\text{Fit}}(1)$ $P_{\text{Fit}}(0)+P_{\text{Fit}(1)}$ $P^{\prime}_{\text{Fit}}(0)$ $P^{\prime}_{\text{Fit}}(1)$ $P^{\prime}_{\text{Fit}}(0)+P^{\prime}_{\text{Fit}}(1)$
$60^{\circ}$ 4 0.92 0.08 1.00 0.07 0.93 1.00
32 0.99 0.01 1.00 0.22 0.78 1.00
64 0.99 0.01 1.00 0.24 0.76 1.00
1024 0.99 0.01 1.00 0.25 0.75 1.00
2048 0.99 0.01 1.00 0.25 0.75 1.00
$70^{\circ}$ 4 0.92 0.08 1.00 0.01 0.99 1.00
32 0.99 0.01 1.00 0.10 0.90 1.00
128 0.99 0.01 1.00 0.11 0.89 1.00
1024 0.99 0.01 1.00 0.12 0.88 1.00
2048 0.99 0.01 1.00 0.12 0.88 1.00
$80^{\circ}$ 4 0.92 0.08 1.00 0.02 0.98 1.00
32 0.99 0.01 1.00 0.03 0.97 1.00
128 0.99 0.01 1.00 0.04 0.96 1.00
1024 0.99 0.01 1.00 0.04 0.96 1.00
2048 0.99 0.01 1.00 0.04 0.96 1.00
Refer to caption Refer to caption Refer to caption
(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
Refer to caption Refer to caption Refer to caption
(d) $\alpha=60^{\circ}$ (e) $\alpha=70^{\circ}$ (f) $\alpha=80^{\circ}$
Figure A4: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $8\times 8$ matrices using $10^{5}$ shots ((a) through (c)) and using $10^{6}$ ((d) through (f))
Refer to caption Refer to caption Refer to caption
(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
Figure A5: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $16\times 16$ matrices and using $10^{6}$ shots throughout.
Refer to caption Refer to caption Refer to caption
(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
Figure A6: Data for $32\times 32$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
Refer to caption Refer to caption Refer to caption
(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
Figure A7: Data for $64\times 64$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
Figure A8: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) of HHL and Psi-HHL, as the condition number grows from $2^{2}$ to $2^{20}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$. Only the case where $\alpha=60^{\circ}$, $A$ not diagonal, $\vec{b}$ in an unequal superposition, and number of shots =$10^{6}$ is considered. The blue cells indicate cases where HHL gives outputs, and red denotes cases where HHL fails. The white cells denote those cases for which we did not carry a calculation, due to computational cost.
Table A5: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $8\times 8$ matrices.
HHL Psi-HHL
$\alpha=60^{\circ}$ $\alpha=70^{\circ}$ $\alpha=80^{\circ}$
# shots $\approx\mathcal{\kappa}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 10 -6.75(0.82, 0.18) 4.90(0.07, 0.93) 5.99(0.05, 0.95) 6.97(0.10, 0.90)
$2^{3}$ 11 -4.69(0.95, 0.05) -3.11(0.11, 0.89) -2.93(0.03, 0.97) -2.93(0.01, 0.99)
$2^{4}$ 12 -9.27(0.98, 0.02) -2.75(0.18, 0.82) -2.48(0.07, 0.93) 2.54(0.02, 0.98)
$2^{5}$ 13 1.87(0.99, 0.01) 1.60(0.21, 0.79) -0.90(0.10, 0.90) -0.09(0.02, 0.98)
$2^{6}$ 14 0.12(0.99, 0.01) -1.20(0.23, 0.77) -1.21(0.11, 0.89) -4.35(0.03, 0.97)
$2^{7}$ 15 0.09(0.99, 0.01) -1.96(0.24, 0.76) -3.83(0.12, 0.88) 0.40(0.03, 0.97)
$2^{8}$ 16 nan(0.99, 0.01) -0.33(0.25, 0.75) -1.74(0.12, 0.88) -4.53(0.03, 0.97)
$2^{9}$ 17 nan(0.99, 0.01) -13.83(0.25, 0.75) -3.82(0.12, 0.88) 10.35(0.04, 0.96)
$2^{10}$ 18 nan(0.99, 0.01) -1.66(0.25, 0.75) 0.94(0.12, 0.88) -7.44(0.04, 0.96)
$2^{11}$ 19 nan(0.99, 0.01) -0.82(0.25, 0.75) 4.90(0.12, 0.88) 3.13(0.04, 0.96)
$2^{12}$ 20 nan(0.99, 0.01) -0.91(0.25, 0.75) -7.52(0.12, 0.88) 10.89(0.04, 0.96)
$2^{13}$ 21 nan(0.99, 0.01) 2.62(0.25, 0.75) 8.41(0.12, 0.88) -0.26(0.04, 0.96)
$2^{14}$ 22 nan(0.99, 0.01) -3.20(0.25, 0.75) 7.12(0.12, 0.88) 8.53(0.04, 0.96)
$2^{15}$ 23 nan(0.99, 0.01) 5.74(0.25, 0.75) -7.34(0.12, 0.88) -7.60(0.04, 0.96)
$2^{16}$ 24 nan(0.99, 0.01) -5.89(0.25, 0.75) -4.65(0.12, 0.88) -13.89(0.04, 0.96)
$2^{17}$ 25 nan(0.99, 0.01) -6.58(0.25, 0.75) -10.10(0.12, 0.88) 2.78(0.04, 0.96)
$10^{5}$ $2^{18}$ 26 nan(0.99, 0.01) 1.03(0.25, 0.75) -3.87(0.12, 0.88) 4.50(0.04, 0.96)
$2^{2}$ 10 -6.40(0.82, 0.18) 5.23(0.10, 0.99) 6.13(0.05, 0.95) 7.15(0.10, 0.90)
$2^{3}$ 11 -4.69(0.95, 0.05) -3.14(0.10, 0.90) -3.02(0.03, 0.97) -2.83(0.01, 0.99)
$2^{4}$ 12 -8.84(0.98, 0.02) -2.60(0.18, 0.82) -2.58(0.07, 0.93) -2.71(0.02, 0.98)
$2^{5}$ 13 2.03(0.99, 0.01) 0.04(0.21, 0.79) 0.03(0.09, 0.91) -0.04(0.02, 0.98)
$2^{6}$ 14 -8.55(0.99, 0.01) -1.49(0.23, 0.77) -0.73(0.11, 0.89) -1.09(0.03, 0.97)
$2^{7}$ 15 -4.68(0.99, 0.01) -2.18(0.24, 0.76) 1.16(0.12, 0.88) 0.36(0.03, 0.97)
$2^{8}$ 16 -0.39(0.99, 0.01) 1.51(0.25, 0.75) -1.35(0.12, 0.88) -1.61(0.04, 0.96)
$2^{9}$ 17 -10.48(0.99, 0.01) 2.19(0.25, 0.75) 1.34(0.12, 0.88) -1.99(0.04, 0.96)
$2^{10}$ 18 -29.14(0.99, 0.01) 0.08(0.25, 0.75) -3.98(0.12, 0.88) 0.33(0.04, 0.96)
$2^{11}$ 19 nan(0.99, 0.01) 1.51(0.25, 0.75) -4.00(0.12, 0.88) -1.72(0.04, 0.96)
$2^{12}$ 20 nan(0.99, 0.01) 1.34(0.25, 0.75) 0.28(0.12, 0.88) 0.00(0.04, 0.96)
$2^{13}$ 21 nan(0.99, 0.01) -2.83(0.25, 0.75) 1.03(0.12, 0.88) 1.99(0.04, 0.96)
$2^{14}$ 22 nan(0.99, 0.01) 3.20(0.25, 0.75) -4.70(0.12, 0.88) 1.75(0.04, 0.96)
$2^{15}$ 23 nan(0.99, 0.01) -0.19(0.25, 0.75) -4.56(0.12, 0.88) -0.58(0.04, 0.96)
$2^{16}$ 24 nan(0.99, 0.01) 0.97(0.25, 0.75) 0.14(0.12, 0.88) -0.11(0.04, 0.96)
$2^{17}$ 25 nan(0.99, 0.01) 0.62(0.25, 0.75) 3.21(0.12, 0.88) 3.19(0.04, 0.96)
$10^{6}$ $2^{18}$ 26 nan(0.99, 0.01) -0.56(0.25, 0.75) -5.17(0.12, 0.88) 2.30(0.04, 0.96)
Table A6: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $16\times 16$ matrices.
HHL Psi-HHL
$60^{\circ}$ $70^{\circ}$ $80^{\circ}$
# shots $\approx\kappa$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 12 -5.90(0.83,0.17) 3.60(0.06, 0.94) 4.29(0.04,0.96) 5.11(0.08,0.92)
$2^{3}$ 13 -9.09(0.95,0.05) -3.42(0.11,0.89) -3.26(0.03,0.97) 3.19(0.01,0.99)
$2^{4}$ 14 -15.82(0.98,0.02) -2.09(0.18, 0.82) -2.17(0.07,0.93) -1.76(0.01,0.99)
$2^{5}$ 15 -8.45(0.99,0.01) -0.75(0.21, 0.79) -0.53(0.10,0.91) 0.92(0.02,0.98)
$2^{6}$ 16 -25.62(0.99,0.01) 0.10(0.23, 0.77) 0.60(0.11,0.89) -0.06(0.03,0.97)
$2^{7}$ 17 -24.21(0.99,0.01) -1.22(0.24, 0.76) -1.18(0.11,0.89,) -2.15(0.03,0.97)
$2^{8}$ 18 2.25(0.99,0.01) -1.30(0.25, 0.75) 0.47(0.12, 0.88) 2.68(0.03,0.97)
$2^{9}$ 19 nan(0.99,0.01) -0.85(0.25, 0.75) 3.64(0.12,0.88) -0.04(0.04,0.97)
$2^{10}$ 20 nan(0.99,0.01) 1.75(0.25, 0.75) -0.11(0.12,0.88) -5.97(0.04,0.96)
$2^{11}$ 21 nan(0.99,0.01) -2.02(0.25, 0.75) -2.59(0.12,0.88) 3.26(0.04,0.96)
$2^{12}$ 22 nan(0.99,0.01) -2.75(0.25, 0.75) -4.15(0.12,0.88) -7.93(0.04,0.96)
$2^{13}$ 23 nan(0.99,0.01) -4.07(0.25, 0.75) 0.29(0.12,0.88) 1.88(0.04,0.96)
$2^{14}$ 24 nan(0.99,0.01) -3.97(0.25, 0.75) -0.83(0.12,0.88) 1.30(0.04,0.96)
$2^{15}$ 25 nan(0.99,0.01) 7.01(0.25, 0.75) -1.46(0.12,0.88) 2.07(0.04,0.96)
$10^{6}$ $2^{16}$ 26 nan(0.99,0.01) -1.31(0.25, 0.75) -5.10(0.12,0.88) -11.71(0.04,0.96)
Table A7: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $32\times 32$ matrices.
HHL Psi-HHL
$60^{\circ}$ $70^{\circ}$ $80^{\circ}$
# shots $\approx\mathcal{\kappa}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 14 -0.38(0.17, 0.83) 1.73(0.06, 0.94) 1.72(0.04, 0.96) 1.68(0.08,0.92)
$2^{3}$ 15 -2.44(0.05, 0.95) -0.37(0.12, 0.88) -0.25(0.04, 0.96) -0.32(0.01,0.99)
$2^{4}$ 16 -9.53(0.02, 0.98) -1.73(0.18, 0.82) -1.71(0.07, 0.93) -1.75(0.01,0.99
$2^{5}$ 17 -3.12(0.01, 0.99) -0.71(0.21, 0.79) -1.30(0.09, 0.91) -0.64(0.02, 0.98)
$2^{6}$ 18 -18.33(0.01,0.99) 0.29(0.23, 0.97) -1.13(0.11, 0.89) 0.52(0.03, 0.97)
$2^{7}$ 19 -15.11(0.01, 0.99) -0.52(0.24, 0.76) -1.25(0.11,0.89) -1.00(0.03, 0.97
$2^{8}$ 20 -13.01(0.01, 0.99) -0.90(0.25, 0.75) -0.03(0.12, 0.88) -0.25(0.03, 0.97
$2^{9}$ 21 -11.12(0.01, 0.99) -0.15(0.25, 0.75) 1.12(0.12, 0.88) -1.40(0.04, 0.96)
$2^{10}$ 22 nan(0.01, 0.99) -0.44(0.25, 0.75) 2.37(0.12, 0.88) 6.12(0.04, 0.96)
$2^{11}$ 23 nan(0.01, 0.99) 4.36(0.25, 0.75) 1.54(0.12, 0.88) -0.60(0.04, 0.96)
$2^{12}$ 24 nan(0.01, 0.99) 0.72(0.25, 0.75) 0.23(0.12, 0.88) -1.07(0.04, 0.96)
$2^{13}$ 25 nan(0.01, 0.99) -2.31(0.25, 0.75) -1.56(0.12, 0.88) 1.31(0.04, 0.96)
$10^{6}$ $2^{14}$ 26 nan(0.01, 0.99) 0.82(0.25, 0.75) 4.95(0.12, 0.88) 1.70(0.04, 0.96)
Table A8: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $64\times 64$ matrices.
HHL Psi-HHL
$60^{\circ}$ $70^{\circ}$ $80^{\circ}$
# shots $\approx\mathcal{\kappa}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
$2^{2}$ 16 -3.13(0.82, 0.17) 0.01(0.12, 0.88) 0.21(0.10, 0.91) 0.82(0.10, 0.90)
$2^{3}$ 17 -2.26(0.94, 0.05) -1.06(0.012, 0.88) -0.95(0.04, 0.96) -1.01(0.02, 0.98)
$2^{4}$ 18 -14.67(0.98, 0.02) -1.82(0.17, 0.83) 1.47(0.03, 0.97) -1.56(0.02,0.98)
$2^{5}$ 19 -2.16(0.99, 0.01) -1.27(0.21, 0.79) -1.57(0.10, 0.91) -0.86(0.02,0.98)
$2^{6}$ 20 -1.28(1.00, 0.00) -0.55(0.23, 0.77) -1.26(0.10, 0.90) -1.06(0.02, 0.98)
$2^{7}$ 21 -6.18(1.00, 0.00) 0.16(0.24, 0.76) -0.10(0.11, 0.89) -0.10(0.03, 0.97)
$2^{8}$ 22 nan(1.00, 0.00) -0.86(0.25,0.75) -0.45(0.11, 0.89) 0.18(0.03, 0.97)
$2^{9}$ 23 nan(1.00, 0.00) -0.40(0.25, 0.75) 0.05(0.12, 0.88) -2.46(0.03, 0.97)
$10^{6}$ $2^{10}$ 24 nan(1.00, 0.00) 1.47(0.25, 0.75) 1.47(0.12, 0.88) 2.09(0.03, 0.97)
Appendix A2 Details on calculations involving quantum chemistry matrices
A2.1 $A$ matrix and $\vec{b}$ vector generation
We begin by explaining the procedure for generating the $A$ matrix of size $n$. We use the configuration state function-based Graphical Unitary Group Approach-Configuration Interaction Singles and Doubles (GUGA-CISD) method available in the GAMESS program [26] to construct the Hamiltonian for a given molecule in a specified single particle basis. GUGA-CI ensures that the Hamiltonian is constructed in an orthogonal basis; hence, it is Hermitian. Also, because it is a spin-adapted CI method, it avoids spin contamination of the wave function, thereby giving accurate results. We remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature). The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the Hartree-Fock (HF) energy and the correlation energy, with the latter being much lesser than the former). This is desirable when one executes HHL-LCC calculations on NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself. We also note that unlike the traditional approach where the HF energy is subtracted from the diagonal entries, we take away the MP2 energy, in order to make sure that $A$ is positive definite. In a sense, subtracting MP2 energy could be thought of as a level shift over and beyond subtracting the HF energy.
We now focus our attention on the elements of the $A$ matrix. The size of an $A$ matrix, $m$, generated from a quantum chemical calculation, which we can denote as $A_{mol}$, is not necessarily of the dimension $n\times n$. In such cases, we embed the $A_{mol}$ matrix of dimension ${m\times m}$ as a diagonal block in a larger matrix of dimension $n\times n$, with the other diagonal block being an identity matrix of size $(n-m)\times(n-m)$, denoted as $\mathcal{I}$. In such a case, we solve a system of linear equations given by
$A=\begin{bmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{bmatrix}\begin{bmatrix}\vec{b}\\ 0\end{bmatrix}=\begin{bmatrix}\vec{x}\\ 0\end{bmatrix}.$ (A25)
This approach ensures that the eigenvalues of $A_{mol}$ remain unaffected, as long as its maximum eigenvalue is $\approx$ 1. We choose our matrices carefully so that this condition is satisfied. In the event that the maximum diagonal entry of $A$ is less than 1, it is recommended to pad with the largest (or the smallest) diagonal entry of $A$ so that the number of shots expended is reduced. As an example, we picked a $3\times 3$ sparse matrix such that the largest diagonal entry is 0.70: $\begin{pmatrix}0.25&0.10&0.00\\ 0.10&0.45&0.20\\ 0.00&0.20&0.70\end{pmatrix}$, and considered two cases: padding with 1 and with 0.70. We found that kappa is larger in the latter (5.16) than in the former (4.21).
Since for the problems we consider, $\mathcal{\kappa}$ increases with system size monotonically, and further, going from an unpadded to a padded matrix always right shifts a data point, it is sufficient to carry out $\mathcal{\kappa}$ versus $m$ analysis for the unpadded case, as it upper bounds the complexity of the padded scenario of $\mathcal{\kappa}$ versus $n$.
Figure A9: Sub-figure (a) shows the increase in system size with number of virtuals, $n_{v}$, for RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b) and (c) show that the $A$ matrix increases as $n_{o}^{2}n_{v}^{2}$ for both cases, as one would expect.
A2.2 Orbital character analysis
We now explain the procedure that we adopted for selecting orbitals, based on orbital character. We begin with RbH, and then proceed to explain our procedure for CsH (with accompanying data provided in Table A9).
We fix the number of core orbitals to 13. We now describe the method adopted in choosing virtual molecular orbitals (MOs). Instead of taking the approach of ordering the virtual MOs by energy in ascending order and then imposing an energy cut-off to define our active space, we pick those MOs that are dominated by important atomic orbitals (AOs). We select the important AOs, and for the purposes of this work, we pick those AOs in the order of the principal quantum numbers. For example, in the case of RbH, we can decide the important AOs from the set that arise from our choice of basis, namely the Sapporo double zeta basis: $5s$ through $7p$. For the purposes of our analysis on condition number versus system size, we pick all the AOs that the basis allows in order to obtain more data points for fitting (details given in Table A9). We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 13 occupied MOs and 1 virtual one has a system size, $m$, of $40\times 40$, and subsequently upon padding, an $A$ matrix size, $n$, of $64\times 64$. For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5p_{y}$, leading to a $119\times 119$ matrix, which, after padding, becomes a matrix of size $128\times 128$. We continue to construct $A$ matrices in this fashion until we reach $m=9065$, for which $\mathcal{\kappa}$ is about 348.
For the case of CsH molecule, we fix the number of occupied orbitals to 18 while progressively increasing the number of valence orbitals. The order in which we add virtual MOs as well as each such virtual MO’s details are mentioned in Table A9.
Refer to caption Refer to caption
(a) (b)
Figure A10: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the RbH molecule (sub-figure (a)) and the CsH molecule (sub-figures (b)) in the Sapporo double zeta basis set. The system size is determined by increase in number of virtuals.
Refer to caption Refer to caption
(a) (b)
Figure A11: Plots showing the increase in sparsity, s, with $A$ matrix size, $m$, for the RbH (sub-figure (a)) and the CsH (sub-figure (b)) molecules in the Sapporo double zeta basis set. The system size is determined by increase in number of virtuals.
Table A9: Table providing the data for condition number ($\mathcal{\kappa}$) of the $n\times n$ $A$ matrix versus system size ($n$), for RbH and CsH molecules, where we vary the virtual orbitals to increase $n$. In the table, $n_{o}$ the number of active spin orbitals, and $n_{v}$ the number of unoccupied spin orbitals. The first row of the column, ‘MO character’, gives information on the atomic orbitals that contribute dominantly to the considered virtual molecular orbital for $m=40$ case, whereas the second row gives the same information about the second virtual added to the active space for $m=119$ case, and so on. The last column specifies the corresponding newly added virtual’s energy in units of Hartree.
Molecule $m$ $\mathcal{\kappa}(n_{o},n_{v})$ MO character Virtual orbital energy (in Hartree)
RbH 40 152.23 (13, 1) Rb $5s-5p_{z}$ hybrid -0.02
119 152.91 (13, 2) Rb $5p_{y}$ 0.04
228 153.18 (13, 3) Rb $5p_{x}$ 0.04
407 157.92 (13, 4) Rb $5s-5p_{z}$ hybrid 0.08
889 192.30 (13, 6) Rb $4d_{x^{2}-y^{2}}$, Rb $4d_{xy}$ 0.57, 0.57
1863 195.79 (13, 9) Rb $4d_{xz}$, Rb $4d_{yz}$, Rb $4d_{z^{2}}$ 0.57, 0.57
3866 228.12 (13, 13) Rb $6s$, Rb $6p_{x}$, Rb $6p_{y}$, Rb $6p_{z}$ 0.57, 0.06, 0.42, 0.42
4518 235.94 (13, 14) Rb $7s$ 0.44
5225 312.02 (13, 15) H $2s$ 0.83
5896 336.53 (13, 16) H $2p_{y}$ 0.87
6595 338.98 (13, 17) H $2p_{x}$ 1.69
7441 345.62 (13, 18) H $2p_{z}$ 1.69
8238 346.55 (13, 19) Rb $7p_{y}$ 1.71
9065 347.49 (13, 21) Rb $7p_{x}$, Rb $7p_{z}$ 2.12, 2.14
CsH 67 465.89 (18, 1) Cs $6p_{x}$ 0.42
210 466.59 (18, 2) Cs $6p_{y}$ 0.42
417 467.08 (18, 3) Cs $6p_{z}$ 0.48
744 487.04 (18, 4) Cs $6s$ 1.49
1171 541.58 (18, 5) Cs $8s$ 0.01
1698 546.70 (18, 6) Cs $7p_{y}$ 0.05
2175 551.19(18, 7) Cs $7p_{x}$ 0.05
2802 573.09(18, 8) Cs $7s$ 0.08
3493 593.31(18, 9) Cs $5d_{y^{2}}$ 0.38
4344 640.22 (18, 10) Cs $5d_{xy}$ 0.38
5215 646.10 (18, 11) Cs $5d_{xz}$ 0.39
6150 652.10 (18, 12) Cs $5d_{yz}$ 0.39
7261 714.59 (18, 13) Cs $5d_{z^{2}}$ 0.39
8473 956.37 (18, 14) H $2s$ 0.93
9784 974.45 (18, 15) H $2p_{x}$ 1.70
11058 1003.49 (18, 16) H $2p_{y}$ 1.70
Table A10: Table providing the minimum and maximum eigenvalues of the matrix $A_{mol}$, denoted in the table as $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ and $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$, for RbH and CsH systems. The energies are in units of milli-Hartree.
Molecule $m$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$
RbH 40 135.31 20599.95
119 135.18 20670.59
228 135.04 20685.98
407 131.94 20837.37
889 109.40 21038.25
1863 107.55 21058.36
3866 92.36 21071.00
4518 90.25 21294.40
5225 72.72 22690.04
5896 72.32 24341.29
6595 71.93 24386.59
7441 71.46 24701.91
8238 71.27 24702.28
9065 71.09 24702.65
CsH 67 128.09 59680.83
210 127.99 59722.33
417 127.90 59737.34
744 122.88 59847.85
1171 110.54 59867.65
1698 109.50 59867.64
2175 108.61 59867.63
2803 104.46 59867.67
3493 100.89 59867.69
4345 93.60 59916.79
5215 92.75 59929.31
6150 91.91 59939.41
7262 83.88 59948.78
8473 64.56 61759.13
9784 63.38 61761.97
11058 63.22 63448.99
Table A11: Average PFDs and Range of PFDs of HHL and Psi-HHL for Li2, RbH and CsH molecules resp. in STO-6G, Sapporo-DZP and Sapporo-DKH3-DZP-2012 basis sets. The PFDs are calculated over a range of 5 repetitions.
Li2 RbH CsH
# shots HHL(PFD, Range) Psi-HHL(PFD, Range) HHL(PFD, Range) Psi-HHL(Range) HHL(PFD, Range) Psi-HHL(PFD, Range)
$5\times 10^{2}$ (-7.68, 40.22) (-2.79, 29.58)
$10^{3}$ (-4.08, 19.80) (-1.46, 8.34) (1.15, 30.41) (2.51, 25.93) (-3.76, 55.34) (0.49, 23.56)
$5\times 10^{3}$ (-2.94, 8.89) (-2.11, 7.57)
$10^{4}$ (-0.58, 6.00) (1.07, 4.47) (3.64, 25.05) (-0.42, 6.18) (-11.83, 9.82) (2.24, 4.22)
$5\times 10^{4}$ (3.29, 16.90) (1.92, 3.14) (-12.38, 2.04) (1.63, 4.77)
$10^{5}$ (5.77, 4.03) (0.60, 1.79) (-11.74, 2.49) (-0.06, 2.29)
$5\times 10^{5}$ (5.29, 0.76) (0.99, 0.71) (-10.49, 2.29) (0.38, 1.77)
$10^{6}$ (5.46, 0.72) (1.07, 0.56) (-10.92, 2.01) (0.56, 1.05)

### v2

@@
 This has the important practical implication of having to use substantially fewer shots relative to the traditional HHL algorithm.
-We carry out two sets of simulations, where we go up to 26-qubit calculations, to demonstrate the ability of Psi-HHL to handle situations involving large $\mathcal{\kappa}$ matrices via: (a) a set of toy matrices, for which we go up to size $64\times 64$ and $\mathcal{\kappa}$ values of up to $\approx$ 1 million, and (b) a deep-dive into quantum chemistry, where we consider matrices up to size $256\times 256$ that reach $\mathcal{\kappa}$ of about 466.
-The molecular systems that we consider are Li2, RbH, and CsH.
+Additionally, we discuss the capability of Psi-HHL to address the cases where $A$ is singular.
+We carry out two sets of simulations, where we go up to 26-qubit calculations, to demonstrate the ability of Psi-HHL to handle situations involving large $\mathcal{\kappa}$ matrices via: (a) a set of toy matrices, for which we go up to size $64\times 64$ and $\mathcal{\kappa}$ values of up to $\approx$ 1 million, and (b) application to quantum chemistry, where we consider matrices up to size $256\times 256$ that reach $\mathcal{\kappa}$ of about 393.
+The molecular systems that we consider are Li2, KH, RbH, and CsH.
 Although the feature of |$\ket{x}$ considered in our examples is an overlap between the input and output states of the HHL algorithm, our approach is general and can be applied in principle to any transition matrix element involving |$\ket{x}$.
@@
 Despite possessing the aforementioned advantages, the HHL algorithm relies upon the probability of successfully post-selecting $1$ on the HHL ancilla qubit, that is, $P(1)$ [1].
-It is known that there is a substantial decrease in $P(1)$ with increase in the condition number of the problem matrix $A$, which we shall denote by $\mathcal{\kappa}$.
+It is known that there can be a substantial decrease in $P(1)$ with increase in the condition number of the problem matrix $A$, which we shall denote by $\mathcal{\kappa}$.
 In other words, the fraction of usable shots in the HHL algorithm becomes very low when large $\mathcal{\kappa}$ input matrices are involved.
@@
 For completeness, we add that the best known classical approach, conjugate gradient, scales as $\mathcal{O}(\sqrt{\mathcal{\kappa}})$ [13], but its scaling with respect to system size is, of course, exponentially more expensive than that for HHL.
-The paper is organized as follows: Section II presents the theory and methodology involved in our work by introducing the Psi-HHL approach, including a discussion on the complexity theoretic aspects (in $\mathcal{\kappa}$) of Psi-HHL.
+The paper is organized as follows: Section II presents the theory and methodology involved in our work by introducing the Psi-HHL approach, including a discussion on the complexity theoretic aspects (in $\mathcal{\kappa}$) of Psi-HHL and on the ability of the algorithm to handle instances where $A$ is singular.
 Section III presents our results for Psi-HHL applied to (a) suitably chosen sets of $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$ toy matrices, as well as for (b) molecular calculations ($64\times 64$ and $256\times 256$ $A$ matrix sizes) in the linearized coupled cluster (LCC) HHL and Psi-HHL frameworks.
@@
 $QPE$ refers to the quantum phase estimation while $CR(\theta)$ refers to the controlled-rotation module that inverts the eigenvalues of $A$.
+The $CR(\theta)$ circuit is built out of uniformly controlled rotations involving a set of angles $\{\theta_{i}\}$, denoted compactly as just $\theta$ in the figure.
 The dashed cambered rectangle around the CNOT gate in the HOM circuit module, and which is accompanied by $(i)$, indicates that the gate is controlled on the ith qubit of the |$\ket{b}$ register.
@@
 We briefly sketch the relevant details of the HHL algorithm (we refer the readers to Refs.
-[14, 15, 16, 17] for detailed descriptions of the algorithm) in Figs.
-1(a) and 1(b), where an amplitude-encoded input |$\arrowvert b\rangle=\sum_{i}b_{i}\arrowvert\nu_{i}\rangle$, such that $\sum_{i}|b_{i}|^{2}=1$ where {|$\{\arrowvert\nu_{i}\rangle\}$ refers to the eigenbasis of $A$, is fed into the HHL algorithm along with $n_{r}$ number of clock register qubits and one HHL ancillary qubit.
+[14, 15, 16, 17] for detailed descriptions of the algorithm) in Figures 1(a) and 1(b), where an amplitude-encoded input |$\arrowvert b\rangle=\sum_{i}b_{i}\arrowvert\nu_{i}\rangle$, such that $\sum_{i}|b_{i}|^{2}=1$ where {|$\{\arrowvert\nu_{i}\rangle\}$ refers to the eigenbasis of $A$, is fed into the HHL algorithm along with $n_{r}$ number of clock register qubits and one HHL ancillary qubit.
 The HHL algorithm itself involves QPE, followed by a non-trivial controlled-rotation module, and then a QPE module.
@@
 The probability of successfully post-selecting outcome ‘$1$’ is given as $P(1)=\||x\rangle_{\rm un}\|^{2}$.
-Finally, when one wants to predict a feature of the solution vector, say, the correlation energy of a molecular system (proportional to the overlap between |$\ket{x}$ and |$\ket{b}$), it can be done by appending an additional circuit (the Hong-Ou-Mandel module [18]), abbreviated as HOM in Fig.
-1(b) that inputs the output, |$\arrowvert x\rangle$, of the HHL algorithm and |$\arrowvert b\rangle$, to output the overlap between the two quantities.
+Finally, when one wants to predict a feature of the solution vector, say, the correlation energy of a molecular system (proportional to the overlap between |$\ket{x}$ and |$\ket{b}$), it can be done by appending an additional circuit (the Hong-Ou-Mandel module [18]), abbreviated as HOM in Figure 1(b) that inputs the output, |$\arrowvert x\rangle$, of the HHL algorithm and |$\arrowvert b\rangle$, to output the overlap between the two quantities.
 The correlation energy is then calculated as
@@
 II.2 The Psi-HHL approach
-In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Fig.
-1(c).
-The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the traditional HHL given in Fig.
-1(a) and Fig.
-1(b).
+In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Figure 1(c).
+The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the traditional HHL given in Figure 1(a) and Figure 1(b).
 The HHL2 entails traditional HHL module appended with a mixing $2\times 2$ unitary, given as $R_{Y}(2\alpha)$ and post-selecting 1 from the mixed signal.
@@
 II.2.2 HHL2: Computing the mixed signal
-The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancilla before the measurement as illustrated in the lower panel of Fig.
-1(c).
+The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancilla before the measurement as illustrated in the lower panel of Figure 1(c).
 This seemingly trivial addition to the algorithm introduces a new degree of freedom via the angle $\alpha$, enabling one to boost the probability of successful matrix inversion in the case of large $\mathcal{\kappa}$.
@@
 $=$ |$-\|\ket{b}_{\rm un}\|^{2}\sum_{i}|b_{i}|^{2}\frac{C}{\tilde{\lambda}_{i}}\rm cot(\alpha).$
-By comparing the right hand sides of Eq.
-II.1 and Eq.
-14, we see that the expressions for $E_{\mathrm{c}}^{\rm HHL}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ differ only by a factor of $\rm cot(\alpha)$ as shown below.
+By comparing the right hand sides of Eq. II.1 and Eq. 14, we see that the expressions for $E_{\mathrm{c}}^{\rm HHL}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the same as shown below:
 $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}=\frac{\frac{E_{\mathrm{c}}^{\mathrm{m}}}{\rm sin(\alpha)}-E_{\mathrm{c}}^{\mathrm{w}}}{\rm cot(\alpha)}=E_{\mathrm{c}}^{\rm HHL}.$ (15)
@@
 $\approx$ $\|\sqrt{(\mathbb{I}-C^{2}A^{-2})}|b\rangle\|^{2},$
-where $\mathbb{I}$ is the identity operator written in the eigenbasis of operator $A$, constant $C=O(1/\mathcal{\kappa})$, where the condition number $\mathcal{\kappa}$ is taken as $\lambda_{\rm max}/\lambda_{\rm min}$ and $|b\rangle$ is the input to the HHL circuit.
+where $\mathbb{I}$ is the identity operator written in the eigenbasis of operator $A$, constant $C=\mathcal{O}(1/\mathcal{\kappa})$, where the condition number $\mathcal{\kappa}$ is taken as $\lambda_{\rm max}/\lambda_{\rm min}$ and $|b\rangle$ is the input to the HHL circuit.
 The set of values $\bigg{\{}\sqrt{1-\frac{C^{2}}{\lambda_{i}^{2}}}\bigg{\}}$ are the eigenvalues of the operator $\sqrt{\mathbb{I}-C^{2}A^{-2}}$, with the largest eigenvalue being $\sqrt{1-\frac{C^{2}}{\lambda_{\rm max}^{2}}}$, and therefore
@@
 One could circumvent this issue by appropriately choosing a scaled $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, leading to $P(0)\geq 1-\gamma^{2}$.
-This idea has been illustrated via an example as shown in Fig.
-2, where we choose the scaling $\gamma=0.5$.
+This idea has been illustrated via an example as shown in Figure 2, where we choose the scaling $\gamma=0.5$.
 Refer to caption Refer to caption
@@
 11 as follows,
-$P^{\prime}(1)$$\approx$ $\rm\|[\sqrt{(\mathbb{I}-C^{2}A^{-2}})\mathrm{sin}(\alpha)$ (18)
+$P^{\prime}(1)$$\approx$ $\|[\sqrt{(\mathbb{I}-C^{2}A^{-2}})\mathrm{sin}(\alpha)$ (18)
 $+$ $(CA^{-1})\mathrm{cos}(\alpha)]|b\rangle\|^{2}.$
@@
 Using once again the sub-multiplicativity of matrix norms in Eq. 18, we can write
-$P^{\prime}(1)$$\leq$ $\rm\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
+$P^{\prime}(1)$$\leq$ $\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
 $+$ $CA^{-1}\cos(\alpha)]\|^{2}_{\rm op}\||b\rangle\|^{2},$
@@
 For the choice of $C=\gamma\lambda_{\min}$, where $\gamma\in(0,1)$, one can push the lower bound for the choice of given $\alpha$.
-This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Fig.
-1(a).
-We show this in Fig.
-2, where for a given $\alpha=60^{\circ}$, we obtain
+This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Figure 1(a).
+We show this in Figure 2, where for a given $\alpha=60^{\circ}$, we obtain
 $\rm\bigg{|}\bigg{[}\bigg{(}\sqrt{1-\frac{\gamma^{2}}{\mathcal{\kappa}^{2}}}\bigg{)}sin(\alpha)+\frac{\gamma}{\mathcal{\kappa}}cos(\alpha)\bigg{]}\bigg{|}^{2}$$\leq$ $P^{\prime}(1)$ (23)
@@
 III.
+II.4 Singular matrices
+When the HHL algorithm is executed for a large system size for some target application, the user ideally should not possess knowledge of some crucial properties of $A$, including its condition number and invertibility.
+In the extreme event that $A$ happens to be singular and its condition number is not known a priori, the algorithm always outputs 00, that is, $P(1)=0$, and a user has to supply a rather large number of shots before perhaps drawing on the realization that the problem may not produce a $1$.
+However, Psi-HHL, by construction, circumvents this issue that HHL suffers from.
+For illustration, we pick two small but extreme examples, as follows,
+  • A $2\times 2$ matrix given by $A=\begin{pmatrix}0.25&0.00\\ 0.00&0.00\end{pmatrix}$ and the $\vec{b}$ is $\begin{pmatrix}0.00\\ 1.00\end{pmatrix}$,
+  • A $4\times 4$ matrix $A=\begin{pmatrix}1.00&2.00&0.00&0.00\\ 2.00&4.00&0.00&0.00\\ 0.00&0.00&0.00&0.00\\ 0.00&0.00&0.00&2.00\end{pmatrix}$ and with $\vec{b}=\begin{pmatrix}0.00\\ 0.00\\ 1.00\\ 0.00\end{pmatrix}$.
+In both the cases above $\vec{b}\in N(A)$, such that $N(A)$ is the null space of $A$ matrix.
+We execute HHL as well as Psi-HHL for these two cases by setting $C/\lambda_{i}$=00 whenever the eigenvalues of $A$ are zero, that is, $\lambda_{i}=0$.
+The traditional HHL algorithm would expend all the supplied shots without being able to post-select outcome $1$, since $P(1)=0$.
+The HHL1 of Psi-HHL successfully post-selects outcome 00 with probability $P(0)=1$.
+On the other hand, the HHL2 module of Psi-HHL post-selects outcome $1$ with probability $\sin^{2}(\alpha)$, see Eq.
+10.
+The data points presented in Figure 3 demonstrate these probabilities for the two examples considered.
+Both these HHLs yield the solution $|x\rangle=|b\rangle$.
+The overlap $|\langle x|b\rangle|$ from the HOM module in HHL1 and HHL2 both yield a value of 1, with all of the additional $\sin({\alpha})$ factors cancelling out (see Figure 1).
+Subtracting the two overlap values results in 0, as expected.
+Thus, the Psi-HHL approach proves powerful when $P(0)\gg P(1)$.
+We now consider the case when $\vec{b}$ is an eigenvector of $A$.
+Here too, one could still get the same result via Psi-HHL by choosing the evolution time, $t$, as a multiple of $2\pi$, which could be incorrect.
+Therefore, in order to reliably calculate the outputs from both the cases using Psi-HHL, one chooses a non-trivial $t$ in the QPE module, such that it is not an integral multiple of $2\pi$, in order to avoid the possibility of obtaining $\lambda_{i}=0$ for both the cases.
+Choosing a non-trivial $t$ will resolve this issue, in the following way:
+  • In the case where $\vec{b}$ is an eigenvector of $A$, one obtains an eigenvalue $\lambda_{i}\neq 0$, yielding an associated nontrivial controlled-rotation angle, $\theta_{i}$.
+  • However, when $\vec{b}$ is a null space vector, one would still obtain an eigenvalue $\lambda_{i}=0$ leading to a controlled-rotation angle of 00.
+          (a)                             (b)
+Figure 3: Figure showing the performance of HHL and Psi-HHL for singular matrices of size $2\times 2$ (sub-figure (a)) and $4\times 4$ (sub-figure (b)).
 III Results and discussions
@@
   • Quantum chemistry: Since quantum chemistry is one of the main applications of interest with quantum computers, it is worth exploring this domain for our analysis [2].
-However, since we no longer are considering toy examples, we begin by qualitatively assessing the applicability of HHL to the LCC problem by:
-    1. 1.
-      Checking how the $A$ matrices grow with system size (and along the way, defining the system size), thereby providing a handle on the classical hardness of the problem, and
-    2. 2.
-      Checking how $\mathcal{\kappa}$ itself grows as a function of what we set to be the system size. Only after this qualitative analysis do we move to our main results for this problem domain.
+However, since we no longer are considering toy examples, we begin by qualitatively assessing the applicability of HHL to the LCC problem by checking how the $A$ matrices grow with system size (and along the way, defining the system size), thereby providing a handle on the classical hardness of the problem.
+We also check the growth of $\mathcal{\kappa}$ itself with what we set to be the system size.
+Following this analysis, we move to our main results for this problem domain.
   • We carry out the following analyses:
@@
       For the toy matrices, we compare HHL and Psi-HHL, and demonstrate the ability of the latter in predicting correlation energies across a range of condition number values.
-As we shall explain later in this section, we find that since $\mathcal{\kappa}$ grows slowly for the quantum chemistry case, it is not computationally feasible to carry out the analysis.
 We quantify the quality of our results for the correlation energies through percentage fraction difference, abbreviated as PFD, given by $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{HHL}}}{E_{\mathrm{c}}}\times 100$ for HHL and $\frac{E_{\mathrm{c}}-E_{\mathrm{c}}^{\mathrm{Psi-HHL}}}{E_{\mathrm{c}}}\times 100$ for Psi-HHL.
-We denote the correlation energy obtained on a traditional computer by $E_{\mathrm{c}}$, whereas $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the values for the quantity obtained from HHL and Psi-HHL respectively.
+We denote the correlation energy obtained using a suitable classical algorithm (we term it as ‘classical value’ hereafter) by $E_{\mathrm{c}}$, whereas $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the values for the quantity obtained from HHL and Psi-HHL respectively.
+As we shall explain later in this section, we find that since $\mathcal{\kappa}$ grows polylogarithmically for the quantum chemistry case, it is not computationally feasible for us to simulate system sizes where HHL breaks down.
     2. 2.
       For both the toy matrices as well as the quantum chemistry matrices, we study the PFDs of HHL and Psi-HHL approaches versus the number of shots.
+    3. 3.
+      Additionally, for the chemistry case, we compare the performance of HHL and Psi-HHL across a limited range of geometries on a potential energy curve (PEC) for a suitably chosen molecule.
+All of our simulation results are obtained by building on the Qiskit program [21].
 Refer to caption Refer to caption Refer to caption
@@
 Our choices for the other matrices, $8\times 8$ through $64\times 64$ can be found in Figure A1 of the Appendix.
-We have set the following values for sparsity: 3 for $8\times 8$, 4 for $16\times 16$, 17 for $32\times 32$, and 7 for $64\times 64$ matrices.
+We have set the following values for sparsity: 3 for $8\times 8$, 4 for $16\times 16$, 5 for $32\times 32$, and 7 for $64\times 64$ matrices.
 Throughout, we have ensured that the eigenvalues of the $A$ matrix lies between $1/\mathcal{\kappa}$ and 1.
 We provide more examples for the $4\times 4$ matrix size in Section A1 of the Appendix.
-Figures 3(a) through 3(e) provides our findings for $4\times 4$ through the $64\times 64$ examples considered.
+Figures 4(a) through 4(e) provides our findings for $4\times 4$ through the $64\times 64$ examples considered.
 Each data point in the sub-figures is the average result obtained over 10 repetitions, with each repetition involving 1 million shots.
+We verify that the observed trends do not change with 50 repetitions by picking the $4\times 4$ matrix as a representative example (Figures A2 and A3 of Supplemental Material).
 We pick $\alpha=60^{\circ}$ for the Psi-HHL calculations.
@@
 Figures A4 through A7 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A4 also giving the information for $10^{5}$ shots.
+Tables A5 through A8 in the Appendix provide the accompanying data.
 We observe that broadly, the trends are similar to what we find for the smaller $4\times 4$ cases.
@@
 In view of the computational cost involved, we did not go all the way up to a condition number of $2^{20}$ for matrices above $4\times 4$ size, and we stopped much earlier for the $64\times 64$ case.
+Refer to caption Refer to caption Refer to caption
+(a) (b) (c)
+Figure 6: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the case of 6 (sub-figure (a)), 8 (sub-figure (b)) and 10 (sub-figure (c)) fixed occupied orbitals in the Sapporo-DKH3-DZP-2012 basis set.
+The left-most legend on top of sub-figure (a) shows black circles whose areas increase proportionally to the number of virtual orbitals considered.
 III.2 Quantum chemistry matrices
@@
 Therefore, it is interesting to consider the performance of Psi-HHL in this scenario.
-In particular, we consider quantum chemistry problems via the LCC approach [2], for this purpose.
-Quantum chemistry on quantum computers can involve different approaches to selecting systems of interest depending on the application domain, including but not limited to performing calculations on numerous candidate molecules of progressively increasing size, picking one molecule and systematically improving the precision in one’s calculations via suitable active space choices (as well as better quality single particle bases, incorporating more physical effects in the molecular non-relativistic Hamiltonian, etc), etc.
-For the purposes of this work, since we only want to compare performance of the HHL and Psi-HHL approaches, we opt for the substantially simpler second approach.
-Refer to caption Refer to caption Refer to caption
-(a) (b) (c)
-Figure 5: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for three representative molecules.
-PFDs of Li2 (sub-figure (a)), RbH (sub-figure (b)), and CsH (sub-figure (c)) molecules versus number of shots (denoted in the figure as ‘No.
-of shots’).
-The centre of each unfilled circle represents the average PFD over a range of 5 repetitions.
-Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$, while for RbH, we evaluate a 25-qubit circuit involving an $A$ matrix size of $256\times 256$.
-For CsH, we carry out 26-qubit calculations where the $A$ matrix size is $256\times 256$.
-The zero PFD value is marked with a horizontal line for ease of visualization.
-III.2.1 The $A$ matrix dimension
-Since we are now going beyond toy examples, we begin with some preliminary qualitative remarks.
-We begin with the growth of the LCC $A_{mol}$ matrix.
-We make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$, where we recall that $n=2^{N}$.
+In particular, we consider quantum chemistry problems via the LCC approach [2].
+We start with preliminary qualitative remarks on the suitability of HHL/Psi-HHL for carrying out LCC calculations, and to that end, discuss the growth of the $A$ matrix dimension and $\mathcal{\kappa}$ with system size.
+This in turn requires us to define system size.
+Since our focus is to demonstrate the performance of Psi-HHL and not a deep-dive into chemistry, we take the approach of: a.
+choosing a few simple diatomic closed-shell molecules in their ground state and in their equilibrium bond lengths, b.
+picking a suitable single particle basis, c.
+deciding on the number of occupied orbitals, and d.
+systematically increasing the number of unoccupied orbitals (virtuals) for system size.
+We choose four representative molecules: the light Li2, the moderately heavy KH and RbH, and the heavy CsH neutral diatomic molecules, all in their equilibrium bond lengths (Li2: 5.0600 Bohr, KH: 4.2300 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr) [23, 24].
+The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule.
+We pick the Sapporo double zeta basis sets for KH, RbH, and CsH (K, Rb, and Cs: Sapporo-DKH3-DZP-2012, and H: Sapporo-DZP) [25], while for Li2, we choose the STO-6G basis [26].
+We employ the $C_{2v}$ point group symmetry for KH, RbH and CsH computations, while for the case of Li2, we pick the $D_{2h}$ symmetry.
+We do not include relativistic effects as our aim is to only compare HHL and Psi-HHL and not account for as many relevant physical effects as possible.
+We now proceed to discuss the dimension of the $A$ matrix and the condition number scaling in LCC for these chosen molecules.
+III.2.1 Dimension of the $A$ matrix
+Since the LCC matrix is not guaranteed to be of dimension $2^{N}\equiv n$, we make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$.
 We go from $A_{mol}$ to $A$ via padding, which we explain in Section A2 of the Appendix.
-  • The number of rows (or columns), $m$, of the $A_{mol}$ matrix, which we shall hereafter think of as the system size, is the number of particle-hole excitations that one considers for a calculation, which in turn is decided by choice of single particle basis and active space size.
+We note that $m$, which we shall hereafter think of as the system size, is the number of particle-hole excitations that one considers for an LCC calculation, which in turn is decided by choice of single particle basis and active space size.
 We denote the number of occupied orbitals as $n_{o}$ and virtuals by $n_{v}$.
-It is common to employ the LCC approach by only considering single and double excitations (termed LCCSD), in which case $A_{mol}$ matrix grows as $m\sim n_{o}^{2}n_{v}^{2}$ for fixed values of $n_{o}$ and $n_{v}$.
-As we include more excitations systematically, such as triples, quadruples, etc, while keeping the number of orbitals fixed, we observe that the $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$, which is exponential.
+The $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$ (the number of orbitals are kept fixed), which is exponential.
 $\mathcal{E}$ refers to the excitation rank.
-  • For a given excitation rank, $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially, for example, see Figure A9 in the Appendix.
-For the purposes of this work, we restrict ourselves to LCCSD approximation ($\mathcal{E}=2$), and study the size of $A_{mol}$ with increase in number of unoccupied orbitals, $n_{v}$, in a molecule, while keeping $n_{o}$ fixed at a reasonable number.
-For our analyses, we do not attempt to consider a molecule in a given active space size to study the variation of $m$ with choice of single particle basis, and defer the study for a future work.
-We choose three examples: the light Li2, the moderately heavy RbH, and the heavy CsH neutral diatomic molecules, all in their equilibrium bond lengths (Li2: 5.06 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr)  [22, 23].
-The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule.
-We pick the Sapporo double zeta basis sets for RbH and CsH (Rb: Sapporo-DZP, Cs: Sapporo-DKH3-DZP-2012, and H: Sapporo-DZP) [24], while for Li2, we choose the STO-6G basis [25].
-We employ the $C_{2v}$ point group symmetry for RbH and CsH computations, while for the case of Li2, we pick the $D_{2h}$ symmetry.
-We fix $n_{o}$ to 3, 13 , and 18 for Li2, RbH, and CsH respectively.
-The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector are discussed in Section A2.A of the Appendix.
-III.2.2 Condition number of the $A$ matrix
-Although HHL offers an exponential advantage in terms of system size, it is worth checking if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the systems that we considered.
+For example, $\mathcal{E}=2$ is the LCCSD approximation, where S stands for single excitations and D for double excitations.
+For a given $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially, for example, see Figure A9 in the Appendix.
+For the purposes of this work, we study the size of $A_{mol}$ with increase in number of unoccupied orbitals, $n_{v}$, in a molecule, while keeping $n_{o}$ fixed at a reasonable number, all within the LCCSD approximation.
+The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector for our computations are discussed in Section A2.A of the Appendix.
+III.2.2 Condition number scaling in LCC
+Although HHL offers an exponential advantage in terms of system size, it is important to check if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the systems that we considered.
 We point out towards the end of Section A2.A of the Appendix that it is sufficient to analyze $\mathcal{\kappa}$ versus $m$; this curve upper bounds $\mathcal{\kappa}$ versus $n$.
-We increase the active space, and thus $m$, in a systematic manner by carrying out an orbital character analysis, whose details we explain in Section A2.B of the Appendix.
-Furthermore, for this analysis, we only consider RbH and CsH, and not Li2 as the system is too small (in the STO-6G basis) to carry out condition number versus system size analysis.
-For RbH, we consider $n_{o}=13$ for our calculations, and vary the number of virtuals from 1 to 21, which corresponds to $A_{mol}$ matrix size of $40\times 40$ to $9065\times 9065$ respectively.
-For CsH, $n_{o}=18$ and virtuals are varied from 1 to 16, while the $A_{mol}$ matrix size goes from $67\times 67$ to $11058\times 11058$.
-Our results shown in Figure A10 of the Appendix indicate that for RbH as well as for CsH, as we go from $m$ of 40 to 9065 and 67 to 11059 respectively, $\mathcal{\kappa}$ grows polylogarithmically (fourth and sixth order respectively).
-This behaviour is also consistent with what one may expect from chemistry intuition: we begin by referring the reader to Table A10 of the Appendix, where we list the largest and smallest eigenvalues of the $A$ matrix for each value of $m$, for both RbH and CsH.
+We increase the size of the active space (via $n_{v}$), and thus $m$ in a systematic manner, by carrying out an orbital character analysis whose details we explain in Section A2.B of the Appendix and also accompanying Table A10.
+Furthermore, for this analysis, we consider KH, RbH and CsH, but not Li2, as the system is too small (in the STO-6G basis) to carry out a $\mathcal{\kappa}$ versus $m$ analysis.
+We start with $(n_{o},n_{v})$ of (6, 6), (8, 8), and (10, 10) for all the three molecules, and for each of these cases, we systematically increase the number of virtuals.
+Figure 6 presents our results.
+We see that the condition number varies as a polylogarithmic function of system size.
+This behaviour is also consistent with what one may expect from chemistry intuition: we begin by referring the reader to Table A11 of the Appendix, where we list the largest and smallest eigenvalues of the $A$ matrix for each value of $m$, for KH, RbH, and CsH.
 The data shows that the increase in $\mathcal{\kappa}$ as we add virtual orbitals is driven by the decrease of the lowest eigenvalue of the $A$ matrix.
 We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock configuration.
-The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the HF configuration).
+The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the Hartree-Fock (HF) configuration).
 Thus, we expect that the decrease in the lowest eigenvalue of $A$ saturates as we keep adding virtual orbitals, and eventually, the lowest eigenvalue of the $A$ matrix will converge to a certain value.
 Thus, the increase of $\mathcal{\kappa}$ saturates too, thus leading one to expect that the condition number would vary as a polylogarithmic function of system size.
-We add at this juncture that while considering only two molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may a reasonable direction to consider.
+We also observe in each sub-figure (that is, for a fixed number of occupied orbitals), for large $m$, the condition number of KH is the largest, followed by RbH, and then CsH.
+This trend can be explained qualitatively using chemical intuition, by looking at the considered orbitals for the (6, 6) case as a representative example.
+Each eigenvalue of $A$ can be thought of as an excited state energy, given that $A$ is sparse and diagonally dominant.
+The highest excited state arises from electron excitations from the lowest occupied orbitals, which are $2p$ and $3d$ for KH and RbH cases respectively.
+It is clear that $2p$ of K has much lower orbital energy than the $3d$ orbital of Rb.
+Thus, we can expect that the $2p$longrightarrow virtual excitation in KH has a higher excitation energy associated with it than the 3d $\longrightarrow$ virtual excitation in RbH.
+Similarly, we can expect that the $4d$longrightarrow virtual excited state of CsH is lower in energy than the $3d$longrightarrow virtual excited state of RbH.
+In summary, we can expect that the excitation energies decrease monotonically as we go from KH to CsH.
+We combine this observation with our data from Table A11 where the maximum eigenvalue grows much faster than the decrease in minimum eigenvalue from KH to RbH to CsH for a given choice of $n_{o}$ and $n_{v}$, to infer that $\mathcal{\kappa}$ decreases monotonically from KH to CsH.
+Furthermore, in Figure 6(c), KH’s condition number reaches a staggering 4000 with all the virtuals considered, whereas in Figures 6(a) and (b), $\mathcal{\kappa}$ is substantially lower.
+This is not surprising, as for sub-figure 6(c), we consider 10 occupied orbitals, and KH has only 10 occupied orbitals and thus the active space contains the $1s$ orbital of K.
+We can invoke the fact that the orbital energy gap is larger between the inner s orbitals, for example, and the density of orbitals becomes higher in the vicinity of highest occupied molecular orbital and the lowest unoccupied molecular orbital.
+Thus, if we include inner core orbitals in our active space, we expect that the excitation energy of the highest state, namely the largest eigenvalue of the $A$ matrix, to be larger.
+In contrast, for the 6, 8, and the 10 occupied orbitals’ cases, RbH and CsH have comparable condition numbers.
+This can be attributed to the fact that the lowest orbital is always the $3d$ of Rb in all the three cases.
+Since the orbital character of the lowest orbital is the same among the three active spaces, they yield very similar condition numbers.
+Figure 7: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for two representative molecules, via PFDs of Li2 (sub-figure (a)) and KH (sub-figure (b)) molecules versus number of shots (denoted in the figure as ‘No.
+of shots’).
+The centre of each unfilled circle represents the average PFD over a range of 30 repetitions.
+Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$ ($\mathcal{\kappa}$ of 64.07), while for KH, we evaluate a 26-qubit circuit involving an $A$ matrix size of $256\times 256$ ($\mathcal{\kappa}$ of 365.40).
+The zero PFD value is marked with a horizontal line for ease of visualization.
+Figure 8: Figure illustrating the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the classical value of the LCCSD total energy (sub-figure (a)) and PFDs (sub-figure (b)) vs bond lengths of the KH molecule obtained from HHL and Psi-HHL methods (26-qubit calculations involving $A$ matrix size of $256\times 256$; red and blue circles indicate HHL and Psi-HHL calculations respectively, averaged over 30 repetitions and with 50000 shots) and those computed with the classical LCC value.
+The associated value of $\mathcal{\kappa}$ is shown just above or below each data point.
+We add at this juncture that while considering only three molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may a reasonable direction to consider.
 However, further analysis would be required to assess what one would expect for $\mathcal{\kappa}(m)$ in general for an HHL-LCC computation.
-We also note that we carried out a similar analysis for sparsity, and found that it too grows polylogarithmically for RbH and CsH (see Figure A11 of the Appendix).
 III.2.3 HHL and Psi-HHL results
@@
 It is not practical to carry out an analysis similar to that in Figure 4, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon.
-For the purposes of our analysis in Figure 4, we vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 5 repetitions.
-For this analysis, we consider Li2, RbH and CsH.
+We next move to Figure 7 (with the accompanying data given in Table A12 of the Appendix), where we vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 30 repetitions.
+For this analysis, we consider Li2 and KH.
 The details are:
   1. 1.
-    Li2: 21-qubit circuit with an $A$ matrix size of $64\times 64$ (built out of single and double excitations arising from 3 occupied and 8 virtual orbitals) in the STO-6G basis, with $\mathcal{\kappa}=64.07$.
+    Li2: 21-qubit circuit with an $A$ matrix size of $64\times 64$ (built out of single and double excitations arising from 3 occupied and 7 virtual orbitals) in the STO-6G basis, with $\mathcal{\kappa}=64.07$.
   2. 2.
-    RbH: 25-qubit circuit with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 13 occupied and 3 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=153.17$.
-  3. 3.
-    CsH: 26-qubit calculations with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 18 occupied and 2 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=466.59$.
-Although RbH and CsH have the same $A$ matrix size, the latter involves a 26-qubit calculation because we set an additional ancilla for CsH due to its larger atomic number.
-We note that for RbH, the same calculation would have incurred more qubits for VQE-UCCSD (32-qubit circuit) and QPE-CASCI (with $n_{r}=8$, 40-qubit circuit) computations, while we carry out a 25-qubit circuit evaluation in the case of HHL-LCC.
-Similarly, for CsH, VQE-UCCSD: 40 qubits, QPE-CASCI: 48 qubits, and HHL-LCC: 26 qubits.
-Our results are presented in Figure 5, with the accompanying data given in Table A11 of the Appendix.
+    KH: 26-qubit circuit with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 6 occupied and 6 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=365.40$.
+We pick these two systems, as in one case, $\mathcal{\kappa}$ is less than 100, and in the other much larger at over 300.
+Repeating the calculations for RbH and CsH may not give any new insights, and hence we restrict the results only to Li2 and KH.
 For Li2, the condition number is very small, and hence we only study the performance of HHL and Psi-HHL between 500 and 10000 shots.
-The data indicates that for Li2, the Psi-HHL approach gives better average PFDs, especially in the regime of fewer number of shots (500 and 1000 shots).
-In the same regime, the spread in PFDs is also lesser for Psi-HHL.
-For the cases of RbH and CsH, as we increase the number of shots, the error bars for both HHL and Psi-HHL are seen to become much smaller.
+The data clearly shows that for Li2, the Psi-HHL approach always has a smaller spread in PFDs when compared to HHL.
+For KH, we observe that at 1000 shots, HHL fails, while Psi-HHL does not, but comes with large error bars.
+As we increase the number of shots, the error bars for both HHL and Psi-HHL are seen to become much smaller, but Psi-HHL consistently has smaller error bars than HHL.
+We now move to our next set of results, namely the PEC for the KH molecule.
+Since it is computationally not feasible to pick the larger system sizes from Figure 6, we choose the 6 occupied and 6 virtual orbitals case.
+This choice corresponds to a 26-qubit calculation with an $A$ matrix size of $256\times 256$.
+We consider geometries in the neighbourhood of the equilibrium bond length (3 geometries in steps of 0.47 Bohr on either side of the equilibrium bond length).
+With such a choice, the correlation energies required to be captured ranges from about 0.3 mHa (milli-Hartree) for 5.65 Bohr geometry to 3 mHa for the 2.82 Bohr geometry (the two extreme points in the 7 geometries considered), thus allowing us to check the performance of HHL and Psi-HHL across a range of correlation energy values.
+Calculating energies across a range of bond lengths further allows us to check across a (limited) range of condition numbers (the smallest being 365.40 and the largest 393.21) the behaviour of HHL and Psi-HHL.
+Each calculation involves 50000 shots and our reported results are averaged over 30 repetitions.
+For our Psi-HHL computations, we pick $\alpha$ of 60 degrees.
+Figure 8(a) (and the accompanying Table A13 of the Appendix) shows the results for $\Delta E$, which is the difference between HHL (Psi-HHL) total energy and and the classical LCC value for the considered geometries.
+We find that Psi-HHL performs better than HHL in predicting correlation energies, especially on the extreme ends of our PEC when $\mathcal{\kappa}$ values are relatively larger than for those points around the equilibrium geometry.
+When we examine Figure 8(b), we see that Psi-HHL performs consistently better than HHL.
+Furthermore, the spread in results across 30 repetitions follows a trend globally: it increases from 2.82 through 5.65 Bohr calculations.
+We recall that the correlation energies themselves decrease monotonically and by as much as an order of magnitude as we move from 2.82 Bohr to 5.65 Bohr geometries.
+Thus, although the condition numbers are comparable (for example, 374.93 and 371.38 for the 3.29 and the 4.71 Bohr cases respectively), since the quantity to be captured (the correlation energy) itself is much smaller for the latter, we need to supply more shots to restrict the spread in PFD across repeats.
+However, since we keep the number of shots fixed throughout the PEC, we observe a larger spread in PFDs.
+We verify this with a simple toy calculation, where we calculate the overlap $|\langle 0|(a|0\rangle+b|1\rangle)|$ using HHL, with 1000 shots and 20 repeats.
+We average our results over 20 repetitions.
+$a$ is varied from 0.26 to 0.99 (below 0.26, we get nan).
+We observe the same trend as in Figure 8(b) (see Figure A10 of Appendix).
 IV Conclusion
-We propose the Psi-HHL algorithm to efficiently handle problems involving $A$ matrices with large condition numbers.
+We propose the Psi-HHL algorithm to efficiently handle problems involving $A$ matrices with large condition numbers ($\mathcal{\kappa}$).
 The approach involves post-selecting 0 in HHL (the ‘wrong signal’), followed by post-selecting 1 from an HHL-like circuit that contains an additional mixing $2\times 2$ unitary (the ‘mixed signal’), and taking the difference between the wrong and the mixed signals to extract the correct signal.
-We present arguments to show that the complexity of Psi-HHL is optimal in condition number for large $\mathcal{\kappa}$ regimes; the complexity of the algorithm scales as $\mathcal{\kappa}$.
+We present arguments to show that the complexity of Psi-HHL is optimal in condition number for large $\mathcal{\kappa}$ regimes; the complexity of the algorithm scales as $\mathcal{O}(\mathcal{\kappa})$.
 This is an improvement over earlier approaches in literature, such as HHL with amplitude amplification (scales quadratically, and incurs several additional gates due to its iterative nature) and HHL with variable time amplitude amplification (scales as $\mathcal{\kappa}\mathrm{log}^{3}\mathcal{\kappa}$).
 It is also important to note that while the optimal $\mathcal{\kappa}$ behaviour can be recovered using a discrete quantum adiabatic approach based quantum linear solver algorithm [12], we achieve the same result using Psi-HHL, which is a simple modification to the HHL algorithm.
-We apply our method to different toy matrices, where we systematically vary the condition number all the way to about 1 million while keeping the number of shots fixed to 1 million, and show that while HHL fails beyond a certain value, Psi-HHL continues to predict good results.
-We also demonstrate the performance of our proposed Psi-HHL approach for quantum chemistry examples, where we once again find that Psi-HHL yields better quality results in much fewer number of shots.
+We also discuss the case of applying Psi-HHL to singular matrices.
+We apply our method to different toy matrices, where we systematically vary the condition number all the way to about 1 million and keeping the number of shots fixed to 1 million for each of the computations, and show that while HHL fails beyond a certain value of $\mathcal{\kappa}$, Psi-HHL continues to predict good results that we quantify via percentage fraction difference (PFD) relative to the classical value as well as via spread in the PFD across many repetitions.
+We also demonstrate the performance of our proposed Psi-HHL approach for quantum chemistry examples with up to 26-qubit computations, and we once again find that Psi-HHL yields better quality results in much fewer number of shots.
 En route, we also carry out a qualitative analysis to check suitability of treating LCC equations using the HHL algorithm.
@@
 Shibdas Roy for nice discussions on complexity.
-VSP and PBT thank Prof.
+PBT thank Prof.
 Debashis Mukherjee for useful discussions on the quantum chemical sections of the work.
@@
 (c) (d)
-Figure A1: Our choices for (a) $8\times 8$, (b) $16\times 16$, (c) $32\times 32$, and (d) $64\times 64$ toy matrices for comparing the performance of HHL and the Psi-HHL algorithms.
-The cell with a star symbol denotes the matrix element with the variable, $n_{r}$ (the number of clock register qubits from the HHL circuit), which sets $\mathcal{\kappa}$.
+Figure A1: Our choices for (a) $8\times 8$, (b) $16\times 16$, (c) $32\times 32$, and (d) $64\times 64$ toy $A$ matrices and un-normalized $\vec{b}$ for comparing the performance of HHL and the Psi-HHL algorithms.
+The cell with a star symbol denotes the matrix element with the variable $n_{r}$ (the number of clock register qubits from the HHL circuit), which sets $\mathcal{\kappa}$.
+A white cell denotes a zero matrix element.
 In our Results section in the main text, we discussed the case of $A$ not diagonal and the amplitude encoded state corresponding to $\vec{b}$ in an unequal superposition.
@@
     $A$ is diagonal and the amplitude encoded state corresponding to $\vec{b}$ is in an equal superposition.
-In particular, $A=\begin{pmatrix}2^{-(n_{r}-1)}&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, $\vec{b}=\begin{pmatrix}1.00\\ 1.00\\ 1.00\\ 1.00\end{pmatrix}$.
+In particular, $A=\begin{pmatrix}2^{-(n_{r}-1)}&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, and the un-normalized $\vec{b}=\begin{pmatrix}1.00\\ 1.00\\ 1.00\\ 1.00\end{pmatrix}$.
 The sparsity, $s$, of $A$ is set to 1.
@@
 Sub-figures (a) through (c) consider the case where $A$ is diagonal and $\vec{b}$ is in equal superposition.
-Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Subfigures (g) through (i) consider the case where $A$ is not diagonal and $\vec{b}$ is in unequal superposition.
+Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal and $\vec{b}$ is in unequal superposition.
+Each data point is generated by taking the average over 50 repetitions.
 Refer to caption Refer to caption Refer to caption
@@
 Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal, and $\vec{b}$ is in unequal superposition.
+Each data point is generated by taking the average over 50 repetitions.
 Table A1: Table presenting the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$) for $4\times 4$ matrices, and in particular, for three cases: $A$ diagonal and $\vec{b}$ in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ equal’), $A$ diagonal and $\vec{b}$ not in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ unequal’), and $A$ not diagonal and $\vec{b}$ in an unequal superposition (denoted in the table as ‘$A$ not diag, $\vec{b}$ unequal’).
 The string ‘nan’ refers to situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result.
-$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
-[…]
-$2^{2}$ 8 -0.44(0.65, 0.35) -0.22(0.21, 0.79) -1.01(0.92, 0.08) -0.10(0.06, 0.94) -0.87(0.92, 0.08) -0.04(0.08, 0.92)
-$2^{3}$ 9 -0.55(0.72, 0.28) -0.07(0.27, 0.73) -5.82(0.97, 0.03) -0.32(0.15, 0.85) -5.72(0.97, 0.03) 0.08(0.15, 0.85)
-$2^{4}$ 10 -0.81(0.74, 0.26) 0.32(0.32, 0.68) 14.39(0.99, 0.01) -0.69(0.2, 0.80) 17.20(0.99, 0.01) 0.26(0.2, 0.80)
-$2^{5}$ 11 -0.17(0.75, 0.25) -0.02(0.35, 0.65) 3.96(0.99, 0.01) 2.85(0.23, 0.77) -0.12(0.99, 0.01) 0.32(0.22, 0.78)
-$2^{6}$ 12 0.30(0.75, 0.25) 0.05(0.37, 0.63) -6.73(0.99, 0.01) 1.66(0.24, 0.76) -6.73(0.99, 0.01) -1.56(0.24, 0.76)
-$2^{7}$ 13 nan(0.75, 0.25) -0.45(0.37, 0.63) nan(0.99, 0.01) 1.91(0.25, 0.75) nan(0.99, 0.01) -2.62(0.25, 0.75)
-$2^{8}$ 14 nan(0.75, 0.25) -0.01(0.37, 0.63) nan(0.99, 0.01) -2.00(0.25, 0.75) nan(0.99, 0.01) -3.87(0.25, 0.75)
-$2^{9}$ 15 nan(0.75, 0.25) -0.22(0.37, 0.63) nan(0.99, 0.01) -7.50(0.25, 0.75) nan(0.99, 0.01) 8.02(0.25, 0.75)
-$2^{10}$ 16 nan(0.75, 0.25) -0.26(0.37, 0.63) nan(0.99, 0.01) -2.33(0.25, 0.75) nan(0.99, 0.01) -7.77(0.25, 0.75)
-$2^{11}$ 17 nan(0.75, 0.25) 0.08(0.37, 0.63) nan(0.99, 0.01) 2.65(0.25, 0.75) nan(0.99, 0.01) -0.55(0.25, 0.75)
-$2^{12}$ 18 nan(0.75, 0.25) 0.03(0.37, 0.63) nan(0.99, 0.01) 7.93(0.25, 0.75) nan(0.99, 0.01) -2.41(0.25, 0.75)
-$2^{13}$ 19 nan(0.75, 0.25) 0.30(0.38, 0.62) nan(0.99, 0.01) 4.67(0.25, 0.75) nan(0.99, 0.01) 3.12(0.25, 0.75)
-$2^{14}$ 20 nan(0.75, 0.25) 0.41(0.38, 0.62) nan(0.99, 0.01) 9.11(0.25, 0.75) nan(0.99, 0.01) -5.60(0.25, 0.75)
-$2^{15}$ 21 nan(0.75, 0.25) 0.13(0.38, 0.62) nan(0.99, 0.01) 9.42(0.26,0.74) nan(0.99, 0.01) -1.11(0.25, 0.75)
-$2^{16}$ 22 nan(0.75, 0.25) 0.54(0.38, 0.62) nan(0.99, 0.01) -0.25(0.25, 0.75) nan(0.99, 0.01) 2.43(0.25, 0.75)
-$2^{17}$ 23 nan(0.75, 0.25) -0.03(0.38, 0.62) nan(0.99, 0.01) -0.40(0.26, 0.74) nan(0.99, 0.01) -7.87(0.25, 0.75)
-$2^{18}$ 24 nan(0.75, 0.25) 0.26(0.37, 0.63) nan(0.99, 0.01) -8.56(0.26, 0.74) nan(0.99, 0.01) 9.69(0.26, 0.74)
-$2^{19}$ 25 nan(0.75, 0.25) 0.19(0.37, 0.63) nan(0.99, 0.01) 6.91(0.26, 0.74) nan(0.99, 0.01) 1.06(0.25, 0.75)
-$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -0.58(0.37, 0.63) nan(0.99, 0.01) 0.30(0.25, 0.75) nan(0.99, 0.01) 6.15(0.25, 0.75)
-$2^{2}$ 8 -0.57(0.65, 0.35) 0.00(0.21, 0.79) -0.87(0.92, 0.08) 0.03(0.08, 0.92) -0.88(0.92, 0.08) 0.04(0.08, 0.92)
-$2^{3}$ 9 -0.61(0.72, 0.28) -0.14(0.27, 0.73) -5.82(0.97, 0.03) -0.04(0.15, 0.85) -5.82(0.97, 0.03) 0.00(0.15, 0.85)
-$2^{4}$ 10 -0.63(0.74, 0.26) -0.04(0.31, 0.69) 14.66(0.99, 0.01) 0.11(0.20, 0.8) 14.95(0.99, 0.01) -0.15(0.20, 0.80)
-$2^{5}$ 11 0.27(0.75, 0.25) -0.06(0.34, 0.66) 1.84(0.99, 0.01) 0.16(0.22, 0.78) 0.61(0.99, 0.01) 0.44(0.22, 0.78)
-$2^{6}$ 12 0.04(0.75, 0.25) 0.00(0.36, 0.64) -0.72(0.99, 0.01) 0.01(0.24, 0.76) 1.68(0.99, 0.01) -0.70(0.24, 0.76)
-$2^{7}$ 13 -0.07(0.75, 0.25) 0.20(0.37, 0.63) 1.35(0.99, 0.01) -1.56(0.25, 0.75) -4.87(0.99, 0.01) -0.11(0.25, 0.75)
-$2^{8}$ 14 -0.13(0.75, 0.25) -0.02(0.37, 0.63) 0.09(0.99, 0.01) -0.04(0.25, 0.75) -4.04(0.99, 0.01) -1.15(0.25, 0.75)
-$2^{9}$ 15 0.05(0.75, 0.25) 0.06(0.37, 0.63) -2.82(0.99, 0.01) 2.82(0.25, 0.75) nan(0.99, 0.01) 2.26(0.25, 0.75)
-$2^{10}$ 16 0.04(0.75, 0.25) -0.04(0.37, 0.63) nan(0.99, 0.01) -0.12(0.25, 0.75) nan(0.99, 0.01) -0.59(0.25, 0.75)
-$2^{11}$ 17 0.19(0.75, 0.25) -0.05(0.37, 0.63) nan(0.99, 0.01) -1.49(0.25, 0.75) nan(0.99, 0.01) 1.41(0.25, 0.75)
-$2^{12}$ 18 0.08(0.75, 0.25) -0.20(0.37, 0.63) nan(0.99, 0.01) 0.11(0.25, 0.75) nan(0.99, 0.01) 1.26(0.25, 0.75)
-$2^{13}$ 19 0.12(0.75, 0.25) 0.01(0.37, 0.63) nan(0.99, 0.01) 0.13(0.25, 0.75) nan(0.99, 0.01) -3.37(0.25, 0.75)
-$2^{14}$ 20 -0.03(0.75, 0.25) -0.00(0.37, 0.63) nan(0.99, 0.01) -4.52(0.25, 0.75) nan(0.99, 0.01) 1.47(0.25, 0.75)
-$2^{15}$ 21 0.05(0.75, 0.25) -0.05(0.37, 0.63) nan(0.99, 0.01) -0.70(0.25, 0.75) nan(0.99, 0.01) -0.79(0.25, 0.75)
-$2^{16}$ 22 0.10(0.75, 0.25) 0.12(0.38, 0.63) nan(0.99, 0.01) 1.57(0.25, 0.75) nan(0.99, 0.01) 0.84(0.25, 0.75)
-$2^{17}$ 23 -0.03(0.75, 0.25) -0.15(0.38, 0.63) nan(0.99, 0.01) -0.40(0.25, 0.75) nan(0.99, 0.01) 0.05(0.25, 0.75)
-$2^{18}$ 24 0.11(0.25, 0.75) 0.00(0.37, 0.63) nan(0.99, 0.01) -1.22(0.25, 0.75) nan(0.99, 0.01) -3.58(0.25, 0.75)
-$2^{19}$ 25 0.08(0.75, 0.25) -0.03(0.37, 0.63) nan(0.99, 0.01) 0.78(0.25, 0.75) nan(0.99, 0.01) 1.41(0.25, 0.75)
-$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.19(0.37, 0.63) nan(0.99, 0.01) -1.52(0.25, 0.75) nan(0.99, 0.01) 0.01(0.25, 0.75)
+$A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
+# shots $\approx\kappa$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
+$2^{2}$ 8 -0.08(0.64, 0.36) 0.01(0.21, 0.79) 0.03(0.92, 0.08) -0.11(0.08, 0.92) 0.07(0.92, 0.08) 0.01(0.08, 0.92)
+$2^{3}$ 9 0.01(0.72, 0.28) -0.17(0.27, 0.73) -0.05(0.97, 0.03) -0.09(0.15, 0.85) -0.02(0.97, 0.03) -0.05(0.15, 0.85)
+$2^{4}$ 10 -0.01(0.74, 0.26) 0.06(0.32, 0.68) 0.31(0.99, 0.01) 0.47(0.20, 0.80) 0.48(0.99, 0.01) 0.39(0.20, 0.80)
+$2^{5}$ 11 -0.06(0.75, 0.25) -0.14(0.35, 0.65) 1.85(0.99, 0.01) -0.04(0.23, 0.77) -0.10(0.99, 0.01) -0.15(0.22, 0.78)
+$2^{6}$ 12 0.08(0.75, 0.25) -0.38(0.37, 0.63) nan(0.99, 0.01) -1.41(0.24, 0.76) nan(0.99, 0.01) -1.75(0.24, 0.76)
+$2^{7}$ 13 0.00(0.75, 0.25) 0.10(0.37, 0.63) nan(0.99, 0.01) 1.39(0.25, 0.75) nan(0.99, 0.01) 1.85(0.25, 0.75)
+$2^{8}$ 14 0.14(0.75, 0.25) -0.10(0.37, 0.63) nan(0.99, 0.01) 0.42(0.25, 0.75) nan(0.99, 0.01) 1.92(0.25, 0.75)
+$2^{9}$ 15 0.12(0.75, 0.25) 0.14(0.37, 0.63) nan(0.99, 0.01) 3.21(0.25, 0.75) nan(0.99, 0.01) -1.05(0.25, 0.75)
+$2^{10}$ 16 -0.01(0.75, 0.25) -0.08(0.37, 0.63) nan(0.99, 0.01) 0.87(0.25, 0.75) nan(0.99, 0.01) 1.00(0.25, 0.75)
+$2^{11}$ 17 -0.05(0.75, 0.25) -0.28(0.37, 0.63) nan(0.99, 0.01) -0.23(0.25, 0.75) nan(0.99, 0.01) 2.19(0.25, 0.75)
+$2^{12}$ 18 0.02(0.75, 0.25) -0.03(0.37, 0.63) nan(0.99, 0.01) 1.38(0.25, 0.75) nan(0.99, 0.01) 0.96(0.25, 0.75)
+$2^{13}$ 19 0.09(0.75, 0.25) -0.38(0.38, 0.62) nan(0.99, 0.01) -0.59(0.25, 0.75) nan(0.99, 0.01) -2.71(0.25, 0.75)
+$2^{14}$ 20 0.27(0.75, 0.25) 0.03(0.38, 0.62) nan(0.99, 0.01) 2.97(0.25, 0.75) nan(0.99, 0.01) -0.36(0.25, 0.75)
+$2^{15}$ 21 0.21(0.75, 0.25) -0.09(0.38, 0.62) nan(0.99, 0.01) -3.51(0.26,0.74) nan(0.99, 0.01) 2.68(0.25, 0.75)
+$2^{16}$ 22 -0.11(0.75, 0.25) 0.05(0.38, 0.62) nan(0.99, 0.01) -1.74(0.25, 0.75) nan(0.99, 0.01) 1.15(0.25, 0.75)
+$10^{5}$ $2^{17}$ 23 0.29(0.75,0.25) 0.05(0.38, 0.62) nan(0.99, 0.01) 3.06(0.26, 0.74) nan(0.99, 0.01) 2.17(0.25, 0.75)
+$2^{2}$ 8 0.00(0.64, 0.36) 0.05(0.21, 0.79) 0.03(0.92, 0.08) 0.02(0.08, 0.92) 0.05(0.92, 0.08) -0.03(0.08, 0.92)
+$2^{3}$ 9 0.00(0.72, 0.28) -0.05(0.27, 0.73) 0.02(0.97, 0.03) 0.00(0.15, 0.85) -0.08(0.97, 0.03) 0.01(0.15, 0.85)
+$2^{4}$ 10 0.00(0.74, 0.26) 0.04(0.27, 0.73) -0.07(0.99, 0.01) -0.10(0.20, 0.8) 0.12(0.99, 0.01) 0.14(0.20, 0.80)
+$2^{5}$ 11 0.06(0.75, 0.25) -0.02(0.31, 0.69) 0.12(0.99, 0.01) 0.07(0.22, 0.78) 0.16(0.99, 0.01) 0.37(0.22, 0.78)
+$2^{6}$ 12 0.06(0.75, 0.25) 0.05(0.34, 0.66) 0.99(0.99, 0.01) 0.29(0.24, 0.76) 0.02(0.99, 0.01) -0.03(0.24, 0.76)
+$2^{7}$ 13 0.10(0.75, 0.25) -0.03(0.36, 0.64) 1.36(0.99, 0.01) 0.17(0.25, 0.75) 2.19(0.99, 0.01) 0.50(0.25, 0.75)
+$2^{8}$ 14 -0.07(0.75, 0.25) 0.07(0.37, 0.63) 8.58(0.99, 0.01) -0.03(0.25, 0.75) 3.63(0.99, 0.01) -1.34(0.25, 0.75)
+$2^{9}$ 15 0.04(0.75, 0.25) 0.04(0.37, 0.63) nan(0.99, 0.01) -0.05(0.25, 0.75) nan(0.99, 0.01) 0.25(0.25, 0.75)
+$2^{10}$ 16 -0.14(0.75, 0.25) 0.05(0.37, 0.63) nan(0.99, 0.01) -0.25(0.25, 0.75) nan(0.99, 0.01) 0.71(0.25, 0.75)
+$2^{11}$ 17 -0.09(0.75, 0.25) 0.10(0.37, 0.63) nan(0.99, 0.01) -0.03(0.25, 0.75) nan(0.99, 0.01) -0.31(0.25, 0.75)
+$2^{12}$ 18 -0.03(0.75, 0.25) 0.06(0.37, 0.63) nan(0.99, 0.01) 0.78(0.25, 0.75) nan(0.99, 0.01) -0.82(0.25, 0.75)
+$2^{13}$ 19 0.04(0.75, 0.25) -0.02(0.37, 0.63) nan(0.99, 0.01) 0.75(0.25, 0.75) nan(0.99, 0.01) -0.87(0.25, 0.75)
+$2^{14}$ 20 -0.12(0.75, 0.25) 0.09(0.37, 0.63) nan(0.99, 0.01) 0.07(0.25, 0.75) nan(0.99, 0.01) -0.05(0.25, 0.75)
+$2^{15}$ 21 0.02(0.75, 0.25) -0.01(0.37, 0.63) nan(0.99, 0.01) 1.00(0.25, 0.75) nan(0.99, 0.01) 0.14(0.25, 0.75)
+$2^{16}$ 22 0.09(0.75, 0.25) 0.02(0.37, 0.63) nan(0.99, 0.01) -0.09(0.25, 0.75) nan(0.99, 0.01) 0.21(0.25, 0.75)
+$10^{6}$ $2^{17}$ 23 0.02(0.75, 0.25) -0.01(0.37, 0.63) nan(0.99, 0.01) -0.66(0.25, 0.75) nan(0.99, 0.01) 0.45(0.25, 0.75)
 Table A2: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=70^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Table A1.
-$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
-# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
-$2^{2}$ 8 -0.04(0.65, 0.35) -0.24(0.23, 0.77) -1.01(0.92, 0.08) -0.18(0.02, 0.98) -0.87(0.92, 0.08) 0.58(0.02, 0.98)
-$2^{3}$ 9 -0.55(0.72, 0.28) -0.28(0.25, 0.75) -5.82(0.97, 0.03) -0.50(0.06, 0.94) -5.72(0.97, 0.03) -0.41(0.06, 0.94)
-$2^{4}$ 10 -0.81(0.74, 0.26) 0.51(0.27, 0.73) 14.39(0.99, 0.01) 0.33(0.12, 0.88) 17.20(0.99, 0.01) 0.44(0.09, 0.91)
-$2^{5}$ 11 -0.17(0.75, 0.25) 0.05(0.29, 0.71) 3.96(0.99, 0.01) -0.17(0.1, 0.90) -0.12(0.99, 0.01) 1.17(0.10, 0.90)
-$2^{6}$ 12 0.30(0.75, 0.25) 0.54(0.30, 0.70) nan(0.99, 0.01) -1.16( 0.12, 0.88) -6.73(0.99, 0.01) 0.99(0.11, 0.89)
-$2^{7}$ 13 nan(0.75, 0.25) 0.74(0.30,0.70) nan(0.99, 0.01) 0.30(0.12, 0.88) nan(0.99, 0.01) 2.20(0.02, 0.88)
-$2^{8}$ 14 nan(0.75, 0.25) 0.40(0.30, 0.69) nan(0.99, 0.01) -7.03(0.12, 0.88) nan(0.99, 0.01) 3.54(0.02, 0.88)
-$2^{9}$ 15 nan(0.75, 0.25) -0.31(0.31, 0.69) nan(0.99, 0.01) 2.88(0.12, 0.88) nan(0.99, 0.01) 1.96(0.02, 0.88)
-$2^{10}$ 16 nan(0.75, 0.25) 0.02(0.31, 0.69) nan(0.99, 0.01) 5.28(0.12, 0.88) nan(0.99, 0.01) 4.93(0.02, 0.88)
-$2^{11}$ 17 nan(0.75, 0.25) -0.30(0.31, 0.69) nan(0.99, 0.01) 2.52(0.12, 0.88) nan(0.99, 0.01) -2.36(0.02, 0.88)
-$2^{12}$ 18 nan(0.75, 0.25) 0.48(0.31, 0.69) nan(0.99, 0.01) -3.50(0.12, 0.88) nan(0.99, 0.01) 0.81(0.02, 0.88)
-$2^{13}$ 19 nan(0.75, 0.25) 0.38(0.31, 0.69) nan(0.99, 0.01) -9.08(0.12, 0.88) nan(0.99, 0.01) -2.22(0.02, 0.88)
-$2^{14}$ 20 nan(0.75, 0.25) 0.15(0.31, 0.69) nan(0.99, 0.01) -0.46(0.12, 0.88) nan(0.99, 0.01) 0.33(0.02, 0.88)
-$2^{15}$ 21 nan(0.75, 0.25) 1.26(0.31, 0.69) nan(0.99, 0.01) 3.66(0.12, 0.88) nan(0.99, 0.01) 6.98(0.02, 0.88)
-$2^{16}$ 22 nan(0.75, 0.25) -0.34(0.31, 0.69) nan(0.99, 0.01) 5.62(0.12, 0.88) nan(0.99, 0.01) -17.07(0.02, 0.88)
-$2^{17}$ 23 nan(0.75, 0.25) 1.31(0.31, 0.69) nan(0.99, 0.01) 5.56(0.12, 0.88) nan(0.99, 0.01) 2.02(0.02, 0.88)
-$2^{18}$ 24 nan(0.75, 0.25) 0.00(0.69, 0.31) nan(0.99, 0.01) 0.00(0.12, 0.88) nan(0.99, 0.01) 18.88(0.13, 0.87)
-$2^{19}$ 25 nan(0.75, 0.25) 0.33(0.31, 0.69) nan(0.99, 0.01) 0.33(0.12, 0.88) nan(0.99, 0.01) 6.42( 0.12, 0.88)
-$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -0.27(0.31, 0.69) nan(0.99, 0.01) -6.81(0.12, 0.88) nan(0.99, 0.01) 6.75(0.12, 0.88)
-$2^{2}$ 8 -0.57(0.64, 0.36) -0.01(0.23, 0.77) -0.87(0.92, 0.08) -0.02(0.02, 0.98) -0.88(0.92, 0.08) -0.01(0.02, 0.98)
-$2^{3}$ 9 -0.61(0.72, 0.28) -0.16(0.22, 0.76) -5.82(0.97, 0.03) 0.04(0.06,0.94) -5.82(0.97, 0.03) 0.00(0.06, 0.94)
-$2^{4}$ 10 -0.63(0.74, 0.26) -0.05(0.27, 0.73) 14.66(0.99, 0.01) -0.26(0.09, 0.91) 14.95(0.99, 0.01) -0.27(0.09, 0.91)
-$2^{5}$ 11 0.27(0.75, 0.25) -0.08(0.29, 0.71) 1.84(0.99, 0.01) -0.26(0.10, 0.90) 0.61(0.99, 0.01) 0.50(0.11, 0.89)
-$2^{6}$ 12 0.04(0.75, 0.25) 0.02(0.3, 0.70) -0.72(0.99, 0.01) 1.32(0.11, 0.89) 1.68(0.99, 0.01) -0.27(0.12, 0.88)
-$2^{7}$ 13 -0.07(0.75, 0.25) -0.07(0.31, 0.69) 1.35(0.99, 0.01) -0.24(0.12, 0.88) -4.87(0.99, 0.01) 0.40(0.12, 0.88)
-$2^{8}$ 14 -0.13(0.75, 0.25) 0.12(0.31, 0.69) 0.09(0.99, 0.01) 0.43(0.12, 0.88) -4.04(0.99, 0.01) 1.91(0.12, 0.88)
-$2^{9}$ 15 0.05(0.75, 0.25) 0.19(0.31, 0.69) -2.82(0.99, 0.01) -1.81(0.12, 0.88) nan(0.99, 0.01) 1.12(0.12, 0.88)
-$2^{10}$ 16 -0.04(0.75, 0.25) -0.26(0.31, 0.69) nan(0.99, 0.01) 1.19(0.12, 0.88) nan(0.99, 0.01) -2.59(0.12, 0.88)
-$2^{11}$ 17 0.18(0.75, 0.25) 0.04(0.31, 0.69) nan(0.99, 0.01) -1.68(0.12, 0.88) nan(0.99, 0.01) -0.85(0.12, 0.88)
-$2^{12}$ 18 0.07(0.75, 0.25) -0.44(0.31, 0.69) nan(0.99, 0.01) 2.36(0.12, 0.88) nan(0.99, 0.01) -5.96(0.12, 0.88)
-$2^{13}$ 19 0.12(0.75, 0.25) 0.15(0.31, 0.69) nan(0.99, 0.01) 5.19(0.12, 0.88) nan(0.99, 0.01) 1.52(0.12, 0.88)
-$2^{14}$ 20 -0.04(0.75, 0.25) 0.25(0.31, 0.69) nan(0.99, 0.01) 2.50(0.12, 0.88) nan(0.99, 0.01) -1.89(0.12, 0.88)
-$2^{15}$ 21 0.06(0.75, 0.25) -0.33(0.31, 0.69) nan(0.99, 0.01) 0.41(0.12, 0.88) nan(0.99, 0.01) -3.65(0.12, 0.88)
-$2^{16}$ 22 0.10(0.75, 0.25) -0.01(0.31, 0.69) nan(0.99, 0.01) -3.09(0.12, 0.88) nan(0.99, 0.01) -1.88(0.12, 0.88)
-$2^{17}$ 23 -0.03(0.75, 0.25) -0.07(0.31, 0.69) nan(0.99, 0.01) 2.54(0.12, 0.88) nan(0.99, 0.01) -0.92(0.12, 0.88)
-$2^{18}$ 24 0.11(0.75, 0.25) 0.07(0.31, 0.69) nan(0.99, 0.01) -0.47(0.88, 0.12) nan(0.99, 0.01) -0.46(0.88, 0.12)
-$2^{19}$ 25 -0.08(0.75, 0.25) 0.09(0.31, 0.69) nan(0.99, 0.01) 1.81(0.12, 0.88) nan(0.99, 0.01) -0.89(0.12, 0.88)
-$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.02(0.31, 0.69) nan(0.99, 0.01) -2.00(0.12, 0.88) nan(0.99, 0.01) 3.65(0.12, 0.88)
+$A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
+# shots $\approx\kappa$ qubits PFD(P(0), P(1)) PFD(P’(0), P’(1)) PFD(P(0), P(1)) PFD(P’(0), P’(1)) PFD(P(0), P(1)) PFD(P’(0), P’(1))
+$2^{2}$ 8 -0.08(0.64, 0.36) -0.03(0.23, 0.77) 0.03(0.92, 0.08) -0.09(0.02, 0.98) 0.07(0.92, 0.08) -0.05(0.02, 0.98)
+$2^{3}$ 9 0.01(0.72, 0.28) -0.17(0.24, 0.76) -0.05(0.97, 0.03) 0.12(0.06, 0.94) -0.02(0.97, 0.03) -0.22(0.06, 0.94)
+$2^{4}$ 10 -0.02(0.74, 0.26) -0.26(0.27, 0.73) 0.31(0.99, 0.01) -0.48(0.12, 0.88) 0.48(0.99, 0.01) -0.17(0.09, 0.91)
+$2^{5}$ 11 -0.06(0.75, 0.25) 0.07(0.29, 0.71) 1.85(0.99, 0.01) 0.70(0.10, 0.90) -0.10(0.99, 0.01) -0.68(0.10, 0.90)
+$2^{6}$ 12 0.09(0.75, 0.25) 0.23(0.30, 0.70) nan(0.99, 0.01) -0.55(0.12, 0.88) nan(0.99, 0.01) 0.18(0.11, 0.89)
+$2^{7}$ 13 -0.01(0.75, 0.25) 0.28(0.30, 0.70) nan(0.99, 0.01) 0.74(0.12, 0.88) nan(0.99, 0.01) 1.88(0.11, 0.88)
+$2^{8}$ 14 0.15(0.75, 0.25) -0.09(0.3, 0.69) nan(0.99, 0.01) 2.95(0.12, 0.88) nan(0.99, 0.01) -1.04(0.12, 0.88)
+$2^{9}$ 15 0.12(0.75, 0.25) 0.57(0.31, 0.69) nan(0.99, 0.01) 0.33(0.12, 0.88) nan(0.99, 0.01) 2.59( 0.12, 0.88)
+$2^{10}$ 16 -0.01(0.75, 0.25) -0.03(0.31, 0.69) nan(0.99, 0.01) -0.69(0.12, 0.88) nan(0.99, 0.01) 1.22(0.12, 0.88)
+$2^{11}$ 17 -0.05(0.75, 0.25) -0.51(0.31, 0.69) nan(0.99, 0.01) 1.78(0.12, 0.88) nan(0.99, 0.01) -1.75(0.12, 0.88)
+$2^{12}$ 18 0.02(0.75, 0.25) 0.06(0.31, 0.69) nan(0.99, 0.01) 4.37(0.12, 0.88) nan(0.99, 0.01) -0.00( 0.12, 0.88)
+$2^{13}$ 19 0.10(0.75, 0.25) -0.38(0.31, 0.69) nan(0.99, 0.01) -2.56(0.12, 0.88) nan(0.99, 0.01) -0.03(0.12, 0.88)
+$2^{14}$ 20 0.28(0.75, 0.25) 0.14(0.31, 0.69) nan(0.99, 0.01) 2.95(0.12, 0.88) nan(0.99, 0.01) -3.04(0.12, 0.88)
+$2^{15}$ 21 0.21(0.75, 0.25) -0.26(0.31, 0.69) nan(0.99, 0.01) -0.37(0.12, 0.88) nan(0.99, 0.01) 3.27(0.12, 0.88)
+$2^{16}$ 22 -0.11(0.75, 0.25) -0.24(0.31, 0.69) nan(0.99, 0.01) -1.35(0.12, 0.88) nan(0.99, 0.01) 2.36(0.12, 0.88)
+$10^{5}$ $2^{17}$ 23 0.29(0.75, 0.25) -0.36(0.31, 0.69) nan(0.99, 0.01) 0.15(0.12, 0.88) nan(0.99, 0.01) 2.02(0.12, 0.88)
+$2^{2}$ 8 -0.01(0.64, 0.36) -0.01(0.23, 0.77) 0.03(0.92, 0.08) 0.00(0.02, 0.98) 0.05(0.92, 0.08) -0.02(0.02, 0.98)
+$2^{3}$ 9 0.00(0.72, 0.28) 0.07(0.24, 0.76) 0.02(0.97, 0.03) 0.01(0.06,0.94) -0.07(0.97, 0.03) 0.01(0.06, 0.94)
+$2^{4}$ 10 0.00(0.74, 0.26) 0.02(0.27, 0.73) -0.07(0.99, 0.01) -0.06(0.09, 0.91) 0.12(0.99, 0.01) 0.14(0.09, 0.91)
+$2^{5}$ 11 0.05(0.74, 0.26) -0.03(0.29, 0.71) 0.13(0.99, 0.01) 0.03(0.10, 0.90) 0.16(0.99, 0.01) 0.37(0.11, 0.89)
+$2^{6}$ 12 0.07(0.75, 0.25) 0.01(0.30, 0.70) 0.99(0.99, 0.01) 0.68(0.11, 0.89) 0.03(0.99, 0.01) -0.03(0.12, 0.88)
+$2^{7}$ 13 0.10(0.75, 0.25) -0.09(0.30, 0.70) 1.36(0.99, 0.01) -0.07(0.12, 0.88) 2.18(0.99, 0.01) 0.51(0.12, 0.88)
+$2^{8}$ 14 -0.07(0.75, 0.25) 0.00(0.31, 0.69) 8.58(0.99, 0.01) -0.12(0.12, 0.88) 3.63(0.99, 0.01) -1.34(0.12, 0.88)
+$2^{9}$ 15 0.05(0.75, 0.25) -0.08(0.31, 0.69) nan(0.99, 0.01) 0.23(0.12, 0.88) nan(0.99, 0.01) 0.25(0.12, 0.88)
+$2^{10}$ 16 -0.14(0.75, 0.25) -0.12(0.31, 0.69) nan(0.99, 0.01) -0.31(0.12, 0.88) nan(0.99, 0.01) 0.71(0.12, 0.88)
+$2^{11}$ 17 -0.09(0.75, 0.25) 0.03(0.31, 0.69) nan(0.99, 0.01) 0.88(0.12, 0.88) nan(0.99, 0.01) -0.31(0.12, 0.88)
+$2^{12}$ 18 -0.04(0.75, 0.25) -0.11(0.31, 0.69) nan(0.99, 0.01) -0.32(0.12, 0.88) nan(0.99, 0.01) -0.82(0.12, 0.88)
+$2^{13}$ 19 0.04(0.75, 0.25) 0.00(0.31, 0.69) nan(0.99, 0.01) 0.33(0.12, 0.88) nan(0.99, 0.01) -0.86(0.12, 0.88)
+$2^{14}$ 20 -0.12(0.75, 0.25) -0.07(0.31, 0.69) nan(0.99, 0.01) 0.84(0.12, 0.88) nan(0.99, 0.01) -0.06(0.12, 0.88)
+$2^{15}$ 21 0.02(0.75, 0.25) 0.01(0.31, 0.69) nan(0.99, 0.01) 0.07(0.12, 0.88) nan(0.99, 0.01) 0.14(0.12, 0.88)
+$2^{16}$ 22 0.09(0.75, 0.25) -0.05(0.31, 0.69) nan(0.99, 0.01) 0.75(0.12, 0.88) nan(0.99, 0.01) 0.21(0.12, 0.88)
+$10^{6}$ $2^{17}$ 23 0.03(0.75, 0.25) 0.06(0.31, 0.69) nan(0.99, 0.01) 1.02(0.12, 0.88) nan(0.99, 0.01) 0.45(0.12, 0.88)
 Table A3: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=80^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Tables A1 and A2.
-$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
-# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
-$2^{2}$ 8 -0.44(0.65, 0.35) 0.43(0.21, 0.79) -1.01(0.92, 0.08) 0.06(0.01, 0.99) -0.87(0.92, 0.08) 0.31(0.02, 0.98)
-$2^{3}$ 9 -0.55(0.72, 0.28) 0.29(0.27, 0.73) -5.82(0.97, 0.03) -0.59(0.02, 0.98) -5.72(0.97, 0.03) -0.92(0.02, 0.98)
-$2^{4}$ 10 -0.81(0.74, 0.26) 2.47(0.32, 0.68) 14.39(0.99, 0.01) 2.59(0.02, 0.98) 17.20(0.99, 0.01) 1.55(0.02, 0.98)
-$2^{5}$ 11 -0.17(0.75, 0.25) 0.67(0.35, 0.65) 3.96(0.99, 0.01) 1.63(0.03, 0.97) -0.12(0.99, 0.01) 1.89(0.03, 0.97)
-$2^{6}$ 12 0.30(0.75, 0.25) -0.17(0.36, 0.64) -6.73(0.99, 0.01) 5.34(0.04, 0.96) -6.73(0.99, 0.01) -4.59(0.03, 0.97)
-$2^{7}$ 13 nan(0.75, 0.25) -0.30(0.37, 0.63) nan(0.99, 0.01) 0.55(0.04, 0.96) nan(0.99, 0.01) 0.37(0.04, 0.96)
-$2^{8}$ 14 nan(0.75, 0.25) -2.17(0.37, 0.63) nan(0.99, 0.01) 6.89(0.04, 0.96) nan(0.99, 0.01) 0.46(0.04, 0.96)
-$2^{9}$ 15 nan(0.75, 0.25) 0.04(0.37, 0.63) nan(0.99, 0.01) -2.64(0.04, 0.96) nan(0.99, 0.01) 4.65(0.04, 0.96)
-$2^{10}$ 16 nan(0.75, 0.25) 1.01(0.37, 0.63) nan(0.99, 0.01) 3.21(0.04, 0.96) nan(0.99, 0.01) -4.06(0.04, 0.96)
-$2^{11}$ 17 nan(0.75, 0.25) -0.20(0.37, 0.63) nan(0.99, 0.01) -5.46(0.04, 0.96) nan(0.99, 0.01) -0.93(0.04, 0.96)
-$2^{12}$ 18 nan(0.75, 0.25) -0.55(0.37, 0.63) nan(0.99, 0.01) -4.80(0.04, 0.96) nan(0.99, 0.01) 9.42(0.04, 0.96)
-$2^{13}$ 19 nan(0.75, 0.25) -0.17(0.37, 0.63) nan(0.99, 0.01) 2.18(0.04, 0.96) nan(0.99, 0.01) 7.83(0.04, 0.96)
-$2^{14}$ 20 nan(0.75, 0.25) 0.56(0.37, 0.63) nan(0.99, 0.01) 5.32(0.04, 0.96) nan(0.99, 0.01) -0.86(0.04, 0.96)
-$2^{15}$ 21 nan(0.75, 0.25) -0.11(0.37, 0.63) nan(0.99, 0.01) 5.89(0.04, 0.96) nan(0.99, 0.01) 3.92(0.04, 0.96)
-$2^{16}$ 22 nan(0.75, 0.25) 0.24(0.37, 0.63) nan(0.99, 0.01) -12.26(0.04, 0.96) nan(0.99, 0.01) 9.60(0.04, 0.96)
-$2^{17}$ 23 nan(0.75, 0.25) -0.94(0.37, 0.63) nan(0.99, 0.01) -2.94(0.04, 0.96) nan(0.99, 0.01) -5.61(0.04, 0.96)
-$2^{18}$ 24 nan(0.75, 0.25) -1.64(0.26, 0.74) nan(0.99, 0.01) -0.60(0.04, 0.96) nan(0.99, 0.01) 5.92(0.04, 0.96)
-$2^{19}$ 25 nan(0.75, 0.25) 1.47(0.27, 0.73) nan(0.99, 0.01) -0.60(0.04, 0.96) nan(0.99, 0.01) -4.92(0.04, 0.96)
-$10^{5}$ $2^{20}$ 26 nan(0.75, 0.25) -1.70(0.26, 0.74) nan(0.99, 0.01) -17.20(0.04, 0.96) nan(0.99, 0.01) 5.03(0.04, 0.96)
-$2^{2}$ 8 -0.57(0.64, 0.36) -0.13(0.28, 0.72) -0.87(0.92, 0.08) 0.00(0.02, 0.98) -0.88(0.92, 0.08) 0.03(0.02, 0.98)
-$2^{3}$ 9 -0.61(0.72, 0.28) 0.14(0.24, 0.76) -5.82(0.97, 0.03) -0.19(0.01, 0.99) -5.82(0.97, 0.03) -0.06(0.01, 0.99)
-$2^{4}$ 10 -0.63(0.74, 0.26) -0.03(0.25, 0.75) 14.66(0.99, 0.01) -0.22(0.02, 0.98) 14.95(0.99, 0.01) -0.02(0.02, 0.98)
-$2^{5}$ 11 0.27(0.75, 0.25) 0.64(0.26, 0.74) 1.84(0.99, 0.01) 0.03(0.03, 0.97) 0.61(0.99, 0.01) 1.11(0.03, 0.97)
-$2^{6}$ 12 0.04(0.75, 0.25) -0.05(0.26, 0.74) -0.72(0.99, 0.01) 1.21(0.04, 0.96) 1.68(0.99, 0.01) 0.68(0.03, 0.97)
-$2^{7}$ 13 -0.07(0.75, 0.25) 0.17(0.26, 0.74) 1.35(0.99, 0.01) -1.22(0.04, 0.96) -4.87(0.99, 0.01) 0.24(0.03, 0.97)
-$2^{8}$ 14 -0.13(0.75, 0.25) -0.17(0.26, 0.74) 0.09(0.99, 0.01) -3.19(0.04, 0.96) -4.04(0.99, 0.01) 0.32(0.04, 0.96)
-$2^{9}$ 15 0.05(0.75, 0.25) -0.50(0.26, 0.74) 2.82(0.99, 0.01) -2.13(0.04, 0.96) nan(0.99, 0.01) -2.48(0.04, 0.96)
-$2^{10}$ 16 -0.03(0.75, 0.25) -0.30(0.26, 0.74) nan(0.99, 0.01) -0.59(0.04, 0.96) nan(0.99, 0.01) -0.24(0.04, 0.96)
-$2^{11}$ 17 0.18(0.75, 0.25) 0.47(0.26, 0.74) nan(0.99, 0.01) -2.93(0.04, 0.96) nan(0.99, 0.01) 0.05(0.04, 0.96)
-$2^{12}$ 18 0.07(0.75, 0.25) 0.40(0.26, 0.74) nan(0.99, 0.01) 0.62(0.04, 0.96) nan(0.99, 0.01) -2.73(0.04, 0.96)
-$2^{13}$ 19 0.11(0.75, 0.25) 0.14(0.26, 0.74) nan(0.99, 0.01) 0.12(0.04, 0.96) nan(0.99, 0.01) 6.03(0.04, 0.96)
-$2^{14}$ 20 -0.03(0.75, 0.25) 0.62(0.26, 0.74) nan(0.99, 0.01) 2.52(0.04, 0.96) nan(0.99, 0.01) 0.28(0.04, 0.96)
-$2^{15}$ 21 0.05(0.75, 0.25) -0.34(0.26, 0.74) nan(0.99, 0.01) 1.03(0.04, 0.96) nan(0.99, 0.01) -6.30(0.04, 0.96)
-$2^{16}$ 22 0.10(0.75, 0.25) -0.59(0.26, 0.74) nan(0.99, 0.01) -3.21(0.04, 0.96) nan(0.99, 0.01) -2.40(0.04, 0.96)
-$2^{17}$ 23 -0.03(0.75, 0.25) 0.49(0.26, 0.74) nan(0.99, 0.01) 0.45(0.04, 0.96) nan(0.99, 0.01) 2.72(0.04, 0.96)
-$2^{18}$ 24 0.11(0.75, 0.25) 0.14(0.27, 0.73) nan(0.99, 0.01) -2.46(0.04, 0.96) nan(0.99, 0.01) -1.52(0.04, 0.96)
-$2^{19}$ 25 0.08(0.75, 0.25) 0.01(0.27, 0.73) nan(0.99, 0.01) -1.95(0.04, 0.96) nan(0.99, 0.01) -0.72(0.04, 0.96)
-$10^{6}$ $2^{20}$ 26 0.16(0.75, 0.25) -0.33(0.27, 0.73) nan(0.99, 0.01) -2.07(0.04, 0.96) nan(0.99, 0.01) -2.46(0.04, 0.96)
+$A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
+# shots $\approx\kappa$ qubits PFD(P(0), P(1)) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
+$2^{2}$ 8 -0.08(0.64, 0.36) -0.08(0.28, 0.72) 0.04(0.92, 0.08) 0.27(0.02, 0.98) 0.07(0.92, 0.08) 0.26(0.02, 0.98)
+$2^{3}$ 9 0.01(0.72, 0.28) -0.14(0.24, 0.76) -0.06(0.97, 0.03) 0.14(0.01, 0.99) -0.02(0.97, 0.03) 0.05(0.02, 0.98)
+$2^{4}$ 10 -0.01(0.74, 0.26) 0.13(0.24, 0.76) 0.31(0.99, 0.01) 0.14(0.02, 0.98) 0.48(0.99, 0.01) -0.37(0.02, 0.98)
+$2^{5}$ 11 -0.06(0.75, 0.25) -0.41(0.26, 0.74) 1.85(0.99, 0.01) 0.27(0.03, 0.97) -0.10(0.99, 0.01) -0.33(0.03, 0.97)
+$2^{6}$ 12 0.08(0.75, 0.25) -0.62(0.26, 0.74) nan(0.99, 0.01) -0.48(0.04, 0.96) nan(0.99, 0.01) -0.28(0.03, 0.97)
+$2^{7}$ 13 -0.01(0.75, 0.25) 0.76(0.26, 0.74) nan(0.99, 0.01) -1.63(0.04, 0.96) nan(0.99, 0.01) 1.86(0.04, 0.96)
+$2^{8}$ 14 0.15(0.75, 0.25) 0.18(0.26, 0.74) nan(0.99, 0.01) -1.02(0.04, 0.96) nan(0.99, 0.01) 1.42(0.04, 0.96)
+$2^{9}$ 15 0.12(0.75, 0.25) 0.20(0.27, 0.73) nan(0.99, 0.01) 5.05(0.04, 0.96) nan(0.99, 0.01) 4.95(0.04, 0.96)
+$2^{10}$ 16 -0.01(0.75, 0.25) 0.25(0.26, 0.74) nan(0.99, 0.01) 2.12(0.04, 0.96) nan(0.99, 0.01) -3.28(0.04, 0.96)
+$2^{11}$ 17 -0.05(0.75, 0.25) 0.85(0.26, 0.74) nan(0.99, 0.01) 0.70(0.04, 0.96) nan(0.99, 0.01) 0.53(0.04, 0.96)
+$2^{12}$ 18 0.02(0.75, 0.25) -0.29(0.26, 0.74) nan(0.99, 0.01) 6.62(0.04, 0.96) nan(0.99, 0.01) 2.93(0.04, 0.96)
+$2^{13}$ 19 0.10(0.75, 0.25) 1.14(0.27, 0.73) nan(0.99, 0.01) 5.23(0.04, 0.96) nan(0.99, 0.01) 0.42(0.04, 0.96)
+$2^{14}$ 20 0.27(0.75, 0.25) -0.46(0.26, 0.74) nan(0.99, 0.01) 6.99(0.04, 0.96) nan(0.99, 0.01) 2.82(0.04, 0.96)
+$2^{15}$ 21 0.21(0.75, 0.25) 0.50(0.26, 0.74) nan(0.99, 0.01) 1.87(0.04, 0.96) nan(0.99, 0.01) 1.49(0.04, 0.96)
+$2^{16}$ 22 -0.11(0.75, 0.25) -0.46(0.27, 0.73) nan(0.99, 0.01) -1.35(0.04, 0.96) nan(0.99, 0.01) 0.76(0.04, 0.96)
+$10^{5}$ $2^{17}$ 23 0.30(0.75, 0.25) -0.36(0.26, 0.74) nan(0.99, 0.01) 0.16(0.04, 0.96) nan(0.99, 0.01) 1.32(0.04, 0.96)
+$2^{2}$ 8 -0.01(0.64, 0.36) 0.08(0.28, 0.72) 0.03(0.92, 0.08) 0.00(0.02, 0.98) 0.05(0.92, 0.08) -0.02(0.02, 0.98)
+$2^{3}$ 9 0.00(0.72, 0.28) -0.21(0.24, 0.76) 0.02(0.97, 0.03) -0.02(0.01, 0.99) -0.07(0.97, 0.03) -0.11(0.01, 0.99)
+$2^{4}$ 10 0.00(0.74, 0.26) -0.20(0.25, 0.75) -0.07(0.99, 0.01) -0.01(0.02, 0.98) 0.12(0.99, 0.01) 0.05(0.02, 0.98)
+$2^{5}$ 11 0.06(0.75, 0.25) 0.19(0.26, 0.74) 0.12(0.99, 0.01) 0.49(0.03, 0.97) 0.16(0.99, 0.01) 0.17(0.03, 0.97)
+$2^{6}$ 12 0.07(0.75, 0.25) -0.40(0.26, 0.74) 0.99(0.99, 0.01) 0.25 (0.04, 0.96) 0.03(0.99, 0.01) 0.23(0.03, 0.97)
+$2^{7}$ 13 0.10(0.75, 0.25) 0.01(0.26, 0.74) 1.36(0.99, 0.01) 0.18(0.04, 0.96) 2.18(0.99, 0.01) 0.18(0.03, 0.97)
+$2^{8}$ 14 -0.07(0.75, 0.25) -0.08(0.26, 0.74) 8.58(0.99, 0.01) -0.09(0.04, 0.96) 3.63(0.99, 0.01) -0.61(0.04, 0.96)
+$2^{9}$ 15 0.05(0.75, 0.25) 0.24(0.26, 0.74) nan(0.99, 0.01) 0.84(0.04, 0.96) nan(0.99, 0.01) 0.01(0.04, 0.96)
+$2^{10}$ 16 -0.14(0.75, 0.25) -0.21(0.26, 0.74) nan(0.99, 0.01) -0.93(0.04, 0.96) nan(0.99, 0.01) 0.63(0.04, 0.96)
+$2^{11}$ 17 -0.09(0.75, 0.25) 0.00(0.26, 0.74) nan(0.99, 0.01) 1.95(0.04, 0.96) nan(0.99, 0.01) -0.37(0.04, 0.96)
+$2^{12}$ 18 -0.04(0.75, 0.25) -0.14(0.26, 0.74) nan(0.99, 0.01) 0.54(0.04, 0.96) nan(0.99, 0.01) 2.01(0.04, 0.96)
+$2^{13}$ 19 0.04(0.75, 0.25) 0.15(0.26, 0.74) nan(0.99, 0.01) -0.75(0.04, 0.96) nan(0.99, 0.01) -1.04(0.04, 0.96)
+$2^{14}$ 20 -0.12(0.75, 0.25) 0.13(0.26, 0.74) nan(0.99, 0.01) 0.32(0.04, 0.96) nan(0.99, 0.01) 0.21(0.04, 0.96)
+$2^{15}$ 21 0.02(0.75, 0.25) 0.01(0.26, 0.74) nan(0.99, 0.01) -0.66(0.04, 0.96) nan(0.99, 0.01) -0.37(0.04, 0.96)
+$2^{16}$ 22 0.08(0.75, 0.25) 0.13(0.26, 0.74) nan(0.99, 0.01) 0.65(0.04, 0.96) nan(0.99, 0.01) 0.45(0.04, 0.96)
+$10^{6}$ $2^{17}$ 23 0.02(0.75, 0.25) -0.24(0.26, 0.74) nan(0.99, 0.01) -2.53(0.04, 0.96) nan(0.99, 0.01) 0.52(0.04, 0.96)
 Table A4: Summary of the fitted probabilities of HHL and Psi-HHL for the 4x4 case $A$ not diagonal and $\vec{b}$ in an unequal superposition of states, $\alpha=60^{\circ},70^{\circ},80^{\circ}$ and the number of shots is fixed to $10^{6}$.
 $\alpha$ $\approx\mathcal{\kappa}$ $P_{\text{Fit}}(0)$ $P_{\text{Fit}}(1)$ $P_{\text{Fit}}(0)+P_{\text{Fit}}(1)$ $P^{\prime}_{\text{Fit}}(0)$ $P^{\prime}_{\text{Fit}}(1)$ $P^{\prime}_{\text{Fit}}(0)+P^{\prime}_{\text{Fit}}(1)$
-$60^{\circ}$ 4 0.92 0.08 1.00 0.07 0.93 1.00
+$60^{\circ}$ 4 0.92 0.08 1.00 0.08 0.92 1.00
 32 0.99 0.01 1.00 0.22 0.78 1.00
@@
 2048 0.99 0.01 1.00 0.25 0.75 1.00
-$70^{\circ}$ 4 0.92 0.08 1.00 0.01 0.99 1.00
+$70^{\circ}$ 4 0.92 0.08 1.00 0.02 0.98 1.00
 32 0.99 0.01 1.00 0.10 0.90 1.00
@@
 32 0.99 0.01 1.00 0.03 0.97 1.00
-128 0.99 0.01 1.00 0.04 0.96 1.00
+128 0.99 0.01 1.00 0.03 0.97 1.00
 1024 0.99 0.01 1.00 0.04 0.96 1.00
@@
 Figure A7: Data for $64\times 64$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
-Figure A8: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) of HHL and Psi-HHL, as the condition number grows from $2^{2}$ to $2^{20}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$.
+Figure A8: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) of HHL and Psi-HHL, as the condition number grows from $2^{2}$ to $2^{17}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$.
 Only the case where $\alpha=60^{\circ}$, $A$ not diagonal, $\vec{b}$ in an unequal superposition, and number of shots = $10^{6}$ is considered.
@@
 $2^{3}$ 11 -4.69(0.95, 0.05) -3.11(0.11, 0.89) -2.93(0.03, 0.97) -2.93(0.01, 0.99)
-$2^{4}$ 12 -9.27(0.98, 0.02) -2.75(0.18, 0.82) -2.48(0.07, 0.93) 2.54(0.02, 0.98)
+$2^{4}$ 12 -9.27(0.98, 0.02) -2.75(0.18, 0.82) -2.48(0.07, 0.93) -2.54(0.02, 0.98)
 $2^{5}$ 13 1.87(0.99, 0.01) 1.60(0.21, 0.79) -0.90(0.10, 0.90) -0.09(0.02, 0.98)
@@
 $2^{16}$ 24 nan(0.99, 0.01) -5.89(0.25, 0.75) -4.65(0.12, 0.88) -13.89(0.04, 0.96)
-$2^{17}$ 25 nan(0.99, 0.01) -6.58(0.25, 0.75) -10.10(0.12, 0.88) 2.78(0.04, 0.96)
-$10^{5}$ $2^{18}$ 26 nan(0.99, 0.01) 1.03(0.25, 0.75) -3.87(0.12, 0.88) 4.50(0.04, 0.96)
+$10^{5}$ $2^{17}$ 25 nan(0.99, 0.01) -6.58(0.25, 0.75) -10.10(0.12, 0.88) 2.78(0.04, 0.96)
 $2^{2}$ 10 -6.40(0.82, 0.18) 5.23(0.10, 0.99) 6.13(0.05, 0.95) 7.15(0.10, 0.90)
@@
 $2^{16}$ 24 nan(0.99, 0.01) 0.97(0.25, 0.75) 0.14(0.12, 0.88) -0.11(0.04, 0.96)
-$2^{17}$ 25 nan(0.99, 0.01) 0.62(0.25, 0.75) 3.21(0.12, 0.88) 3.19(0.04, 0.96)
-$10^{6}$ $2^{18}$ 26 nan(0.99, 0.01) -0.56(0.25, 0.75) -5.17(0.12, 0.88) 2.30(0.04, 0.96)
+$10^{6}$ $2^{17}$ 25 nan(0.99, 0.01) 0.62(0.25, 0.75) 3.21(0.12, 0.88) 3.19(0.04, 0.96)
 Table A6: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $16\times 16$ matrices.
@@
 $2^{2}$ 12 -5.90(0.83,0.17) 3.60(0.06, 0.94) 4.29(0.04,0.96) 5.11(0.08,0.92)
-$2^{3}$ 13 -9.09(0.95,0.05) -3.42(0.11,0.89) -3.26(0.03,0.97) 3.19(0.01,0.99)
+$2^{3}$ 13 -9.09(0.95,0.05) -3.42(0.11,0.89) -3.26(0.03,0.97) -3.19(0.01,0.99)
 $2^{4}$ 14 -15.82(0.98,0.02) -2.09(0.18, 0.82) -2.17(0.07,0.93) -1.76(0.01,0.99)
@@
 $2^{6}$ 16 -25.62(0.99,0.01) 0.10(0.23, 0.77) 0.60(0.11,0.89) -0.06(0.03,0.97)
-$2^{7}$ 17 -24.21(0.99,0.01) -1.22(0.24, 0.76) -1.18(0.11,0.89,) -2.15(0.03,0.97)
+$2^{7}$ 17 -24.21(0.99,0.01) -1.22(0.24, 0.76) -1.18(0.11,0.89) -2.15(0.03,0.97)
 $2^{8}$ 18 2.25(0.99,0.01) -1.30(0.25, 0.75) 0.47(0.12, 0.88) 2.68(0.03,0.97)
@@
 # shots $\approx\mathcal{\kappa}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
-$2^{2}$ 14 -0.38(0.17, 0.83) 1.73(0.06, 0.94) 1.72(0.04, 0.96) 1.68(0.08,0.92)
-$2^{3}$ 15 -2.44(0.05, 0.95) -0.37(0.12, 0.88) -0.25(0.04, 0.96) -0.32(0.01,0.99)
-$2^{4}$ 16 -9.53(0.02, 0.98) -1.73(0.18, 0.82) -1.71(0.07, 0.93) -1.75(0.01,0.99
-$2^{5}$ 17 -3.12(0.01, 0.99) -0.71(0.21, 0.79) -1.30(0.09, 0.91) -0.64(0.02, 0.98)
-$2^{6}$ 18 -18.33(0.01,0.99) 0.29(0.23, 0.97) -1.13(0.11, 0.89) 0.52(0.03, 0.97)
-$2^{7}$ 19 -15.11(0.01, 0.99) -0.52(0.24, 0.76) -1.25(0.11,0.89) -1.00(0.03, 0.97
-$2^{8}$ 20 -13.01(0.01, 0.99) -0.90(0.25, 0.75) -0.03(0.12, 0.88) -0.25(0.03, 0.97
-$2^{9}$ 21 -11.12(0.01, 0.99) -0.15(0.25, 0.75) 1.12(0.12, 0.88) -1.40(0.04, 0.96)
-$2^{10}$ 22 nan(0.01, 0.99) -0.44(0.25, 0.75) 2.37(0.12, 0.88) 6.12(0.04, 0.96)
-$2^{11}$ 23 nan(0.01, 0.99) 4.36(0.25, 0.75) 1.54(0.12, 0.88) -0.60(0.04, 0.96)
-$2^{12}$ 24 nan(0.01, 0.99) 0.72(0.25, 0.75) 0.23(0.12, 0.88) -1.07(0.04, 0.96)
-$2^{13}$ 25 nan(0.01, 0.99) -2.31(0.25, 0.75) -1.56(0.12, 0.88) 1.31(0.04, 0.96)
-$10^{6}$ $2^{14}$ 26 nan(0.01, 0.99) 0.82(0.25, 0.75) 4.95(0.12, 0.88) 1.70(0.04, 0.96)
+$2^{2}$ 14 -5.23(0.17, 0.83) -3.73(0.11, 0.89) -3.56(0.08, 0.92) -3.42(0.10,0.90)
+$2^{3}$ 15 -2.27(0.05, 0.95) -1.28(0.13, 0.87) -1.16(0.04, 0.96) -1.09(0.02,0.98)
+$2^{4}$ 16 -3.27(0.02, 0.98) -1.90(0.18, 0.82) -1.70(0.07, 0.93) -1.47(0.02,0.98)
+$2^{5}$ 17 -0.87(0.01, 0.99) -0.95(0.21, 0.79) -0.45(0.10, 0.90) -0.49(0.02, 0.98)
+$2^{6}$ 18 -2.71(0.00,1.00) -0.56(0.23, 0.77) -0.80(0.11, 0.89) 0.49(0.03, 0.97)
+$2^{7}$ 19 -0.10(0.00, 1.00) -0.93(0.24, 0.76) -0.05(0.11,0.89) -0.61(0.03, 0.97)
+$2^{8}$ 20 5.48(0.00, 1.00) -0.91(0.25, 0.75) 1.97(0.12, 0.88) -4.47(0.03, 0.97)
+$2^{9}$ 21 nan(0.00, 1.00) -3.74(0.25, 0.75) 2.07(0.12, 0.88) -2.92(0.03, 0.97)
+$2^{10}$ 22 nan(0.00, 1.00) -1.17(0.25, 0.75) 2.07(0.12, 0.88) -0.50(0.03, 0.97)
+$2^{11}$ 23 nan(0.00, 1.00) 4.90(0.25, 0.75) 7.32(0.12, 0.88) -3.07(0.03, 0.97)
+$2^{12}$ 24 nan(0.00, 1.00) 6.11(0.25, 0.75) -1.09(0.12, 0.88) 8.85(0.03, 0.97)
+$2^{13}$ 25 nan(0.00, 1.00) 0.76(0.25, 0.75) -4.61(0.12, 0.88) -0.63(0.03, 0.97)
+$10^{6}$ $2^{14}$ 26 nan(0.00, 1.00) 8.74(0.25, 0.75) 7.76(0.12, 0.88) 1.40(0.03, 0.97)
 Table A8: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $64\times 64$ matrices.
@@
 # shots $\approx\mathcal{\kappa}$ qubits PFD($P(0),P(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$) PFD($P^{\prime}(0),P^{\prime}(1)$)
-$2^{2}$ 16 -3.13(0.82, 0.17) 0.01(0.12, 0.88) 0.21(0.10, 0.91) 0.82(0.10, 0.90)
+$2^{2}$ 16 -3.13(0.82, 0.17) 0.01(0.12, 0.88) 0.21(0.09, 0.91) 0.82(0.10, 0.90)
 $2^{3}$ 17 -2.26(0.94, 0.05) -1.06(0.012, 0.88) -0.95(0.04, 0.96) -1.01(0.02, 0.98)
@@
 We remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature).
-The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the Hartree-Fock (HF) energy and the correlation energy, with the latter being much lesser than the former).
+The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the HF energy and the correlation energy, with the latter being much lesser than the former).
 This is desirable when one executes HHL-LCC calculations on NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself.
@@
 In such a case, we solve a system of linear equations given by
-$A=\begin{bmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{bmatrix}\begin{bmatrix}\vec{b}\\ 0\end{bmatrix}=\begin{bmatrix}\vec{x}\\ 0\end{bmatrix}.$ (A25)
+$A=\begin{bmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{bmatrix}\begin{bmatrix}\vec{x}\\ 0\end{bmatrix}=\begin{bmatrix}\vec{b}\\ 0\end{bmatrix}.$ (A25)
 This approach ensures that the eigenvalues of $A_{mol}$ remain unaffected, as long as its maximum eigenvalue is $\approx$ 1.
@@
 Since for the problems we consider, $\mathcal{\kappa}$ increases with system size monotonically, and further, going from an unpadded to a padded matrix always right shifts a data point, it is sufficient to carry out $\mathcal{\kappa}$ versus $m$ analysis for the unpadded case, as it upper bounds the complexity of the padded scenario of $\mathcal{\kappa}$ versus $n$.
-Figure A9: Sub-figure (a) shows the increase in system size with number of virtuals, $n_{v}$, for RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b) and (c) show that the $A$ matrix increases as $n_{o}^{2}n_{v}^{2}$ for both cases, as one would expect.
+Refer to caption Refer to caption
+(a) (b)
+Refer to caption Refer to caption
+(c) (d)
+Figure A9: Sub-figure (a) shows the increase in system size with number of virtuals, $n_{v}$, for KH, RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b), (c) and (d) show that the $A$ matrix of respectively KH, RbH and CsH increases as $n_{o}^{2}n_{v}^{2}$ for both cases, as one would expect.
 A2.2 Orbital character analysis
@@
 We begin with RbH, and then proceed to explain our procedure for CsH (with accompanying data provided in Table A9).
-We fix the number of core orbitals to 13.
+We fix the number of core orbitals to 6, 8 and 10.
 We now describe the method adopted in choosing virtual molecular orbitals (MOs).
@@
 For the purposes of our analysis on condition number versus system size, we pick all the AOs that the basis allows in order to obtain more data points for fitting (details given in Table A9).
-We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 13 occupied MOs and 1 virtual one has a system size, $m$, of $40\times 40$, and subsequently upon padding, an $A$ matrix size, $n$, of $64\times 64$.
-For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5p_{y}$, leading to a $119\times 119$ matrix, which, after padding, becomes a matrix of size $128\times 128$.
-We continue to construct $A$ matrices in this fashion until we reach $m=9065$, for which $\mathcal{\kappa}$ is about 348.
-For the case of CsH molecule, we fix the number of occupied orbitals to 18 while progressively increasing the number of valence orbitals.
+We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 6 occupied MOs and 1 virtual one has a system size, $m$, of $209\times 209$, and subsequently upon padding, an $A$ matrix size, $n$, of $256\times 256$.
+For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5s$, leading to a $279\times 279$ matrix, which, after padding, becomes a matrix of size $512\times 512$.
+We continue to construct $A$ matrices in this fashion until we reach $m=2296$, for which $\mathcal{\kappa}$ is about 389.76.
+We do the same for KH and CsH molecules, we fix the number of occupied orbitals to 6, 8 and 10 occupied oritals while progressively increasing the number of valence orbitals.
 The order in which we add virtual MOs as well as each such virtual MO’s details are mentioned in Table A9.
-Refer to caption Refer to caption
-(a) (b)
-Figure A10: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the RbH molecule (sub-figure (a)) and the CsH molecule (sub-figures (b)) in the Sapporo double zeta basis set.
-The system size is determined by increase in number of virtuals.
-Refer to caption Refer to caption
-(a) (b)
-Figure A11: Plots showing the increase in sparsity, s, with $A$ matrix size, $m$, for the RbH (sub-figure (a)) and the CsH (sub-figure (b)) molecules in the Sapporo double zeta basis set.
-The system size is determined by increase in number of virtuals.
-Table A9: Table providing the data for condition number ($\mathcal{\kappa}$) of the $n\times n$ $A$ matrix versus system size ($n$), for RbH and CsH molecules, where we vary the virtual orbitals to increase $n$.
-In the table, $n_{o}$ the number of active spin orbitals, and $n_{v}$ the number of unoccupied spin orbitals.
-The first row of the column, ‘MO character’, gives information on the atomic orbitals that contribute dominantly to the considered virtual molecular orbital for $m=40$ case, whereas the second row gives the same information about the second virtual added to the active space for $m=119$ case, and so on.
-The last column specifies the corresponding newly added virtual’s energy in units of Hartree.
-Molecule $m$ $\mathcal{\kappa}(n_{o},n_{v})$ MO character Virtual orbital energy (in Hartree)
-RbH 40 152.23 (13, 1) Rb $5s-5p_{z}$ hybrid -0.02
-119 152.91 (13, 2) Rb $5p_{y}$ 0.04
-228 153.18 (13, 3) Rb $5p_{x}$ 0.04
-407 157.92 (13, 4) Rb $5s-5p_{z}$ hybrid 0.08
-889 192.30 (13, 6) Rb $4d_{x^{2}-y^{2}}$, Rb $4d_{xy}$ 0.57, 0.57
-1863 195.79 (13, 9) Rb $4d_{xz}$, Rb $4d_{yz}$, Rb $4d_{z^{2}}$ 0.57, 0.57
-3866 228.12 (13, 13) Rb $6s$, Rb $6p_{x}$, Rb $6p_{y}$, Rb $6p_{z}$ 0.57, 0.06, 0.42, 0.42
-4518 235.94 (13, 14) Rb $7s$ 0.44
-5225 312.02 (13, 15) H $2s$ 0.83
-5896 336.53 (13, 16) H $2p_{y}$ 0.87
-6595 338.98 (13, 17) H $2p_{x}$ 1.69
-7441 345.62 (13, 18) H $2p_{z}$ 1.69
-8238 346.55 (13, 19) Rb $7p_{y}$ 1.71
-9065 347.49 (13, 21) Rb $7p_{x}$, Rb $7p_{z}$ 2.12, 2.14
-CsH 67 465.89 (18, 1) Cs $6p_{x}$ 0.42
-210 466.59 (18, 2) Cs $6p_{y}$ 0.42
-417 467.08 (18, 3) Cs $6p_{z}$ 0.48
-744 487.04 (18, 4) Cs $6s$ 1.49
-1171 541.58 (18, 5) Cs $8s$ 0.01
-1698 546.70 (18, 6) Cs $7p_{y}$ 0.05
-2175 551.19(18, 7) Cs $7p_{x}$ 0.05
-2802 573.09(18, 8) Cs $7s$ 0.08
-3493 593.31(18, 9) Cs $5d_{y^{2}}$ 0.38
-4344 640.22 (18, 10) Cs $5d_{xy}$ 0.38
-5215 646.10 (18, 11) Cs $5d_{xz}$ 0.39
-6150 652.10 (18, 12) Cs $5d_{yz}$ 0.39
-7261 714.59 (18, 13) Cs $5d_{z^{2}}$ 0.39
-8473 956.37 (18, 14) H $2s$ 0.93
-9784 974.45 (18, 15) H $2p_{x}$ 1.70
-11058 1003.49 (18, 16) H $2p_{y}$ 1.70
-Table A10: Table providing the minimum and maximum eigenvalues of the matrix $A_{mol}$, denoted in the table as $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ and $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$, for RbH and CsH systems.
+Table A9: Table providing the data for condition number ($\mathcal{\kappa}$) of the $m\times m$ $A$ matrix versus system size ($m$), for KH, RbH and CsH molecules, where we vary the virtual orbitals to increase $m$.
+In the table, $n_{o}$ is the number of active spin orbitals, and $n_{v}$ is the number of unoccupied spin orbitals.
+The column labelled ‘Reordered virtual MOs’ represents the atomic orbitals reordered according to the ones that contribute dominantly to the virtual molecular orbital.
+The column labelled ‘MO Energies’ gives the molecular virtual orbital energies in units of milli-Hartree.
+Molecule Reordered virtual MOs MO Energies $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
+m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$
+K $4s$ 1151.70 235 233.65(6, 6) 580 41.03(8, 8) 1484 2195.80(10, 10)
+K $4p_{y}$ 3609.70 313 234.44(6, 7) 745 41.57(8, 9) 1761 2206.59(10, 11)
+K $4p_{x}$ 3609.70 370 235.12(6, 8) 932 236.57(8, 10) 2046 2217.42(10, 12)
+K $4p_{z}$ 3663.70 432 236.80(6, 9) 1110 237.74(8, 11) 2451 2323.69(10, 13)
+K $3d_{xy}$ 743.00 541 238.28(6, 10) 1296 238.90(8, 12) 2900 2849.98(10, 14)
+K $3dy^{2}$ 743.00 630 244.66(6, 11) 1540 250.34(8, 13) 3393 2904.40(10, 15)
+K $3d_{yz}$ 747.40 748 250.45(6, 12) 1808 307.03(8, 14) 3794 2933.74(10, 16)
+K $3d_{xz}$ 747.40 889 265.31(6, 13) 2100 312.83(8, 15) 4203 2964.04(10, 17)
+K $3dz^{2}$ 755.30 1044 340.59(6, 14) 2358 316.09(8, 16) 4788 3610.73(10, 18)
+K $6s$ 3.70 1213 348.68(6, 15) 2624 319.44(8,17) 5265 3629.57(10,19)
+K $6p_{y}$ 72.70 1338 371.54(6, 16) 2972 389.11(8,18) 5750 3648.68(10,20)
+K $6p_{x}$ 72.70 1510 374.07(6, 17) 3278 391.14(8, 19) 6427 3674.83(10, 21)
+K $6p_{z}$ 82.80 1711 380.25(6, 18) 3592 393.20(8, 20)
+K $5s$ 147.20 1860 385.51(6, 19) 3996 395.77(8, 21)
+K $5p_{z}$ 692.70 2062 386.30(6, 20)
+K $5p_{y}$ 702.90 2296 389.76(6, 21)
+K $5p_{x}$ 702.90
+H $2s$ 934.00
+H $2p_{y}$ 1668.70
+H $2p_{x}$ 1668.70
+KH H $2p_{z}$ 1914.10
+Rb $5s-5p_{z}$ 940.00 209 102.73(6, 6) 632 104.88(8, 8) 1401 108.16(10,10)
+Rb $5s$ 925.80 279 103.00(6, 7) 744 105.81(8, 9) 1673 111.85(10, 11)
+Rb $5p_{x}$ 2283.00 370 103.74(6, 8) 932 106.63(8, 10) 1963 115.27(10, 12)
+Rb $5p_{y}$ 2283.00 432 104.65(6, 9) 1110 110.27(8, 11) 2321 125.33(10, 13)
+Rb $5p_{z}$ 2305.80 541 105.46(6, 10) 1296 113.64(8, 12) 2713 156.32(10, 14)
+Rb $4d_{xy}$ 635.40 630 109.05(6, 11) 1540 123.51(8, 13) 3141 182.23(10, 15)
+Rb $4dx^{2}$ 635.40 748 112.38(6, 12) 1808 154.29(8, 14) 3537 206.22(10, 16)
+Rb $4d_{xz}$ 641.20 889 122.12(6, 13) 2100 179.90(8, 15) 3951 207.93(10, 17)
+Rb $4d_{yz}$ 641.20 1044 152.70(6, 14) 2358 203.86(8, 16) 4460 213.88(10, 18)
+Rb $4dz^{2}$ 666.90 1213 178.09(6, 15) 2624 205.55(8, 17) 4932 214.42(10, 19)
+Rb $7s$ 3.20 1381 202.02(6, 16) 2972 211.53(8, 18) 5422 214.96(10, 20)
+Rb $7p_{x}$ 64.20 1510 203.70(6, 17) 3278 212.06(8, 19) 6013 216.27(10, 21)
+Rb $7p_{y}$ 64.20 1711 209.67(6, 18) 3592 212.60(8, 20)
+Rb $6s$ 78.40 1860 210.19(6, 19) 3996 213.87(8, 21)
+Rb $6s$ 130.80 2062 210.72(6, 20)
+Rb $6p_{x}$ 477.70 2296 211.94(6, 21)
+Rb $6p_{y}$ 477.70
+Rb $6p_{z}$ 487.50
+H $2p_{x}$ 1678.20
+[…]
+[…]
+Table A10: (Continued.)
+Molecule Reordered virtual MOs MO Energies $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
+m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$ m $\kappa(n_{o},n_{v})$
+Cs $6p_{x}$ 364.80 209 77.05(6, 6) 632 81.29(8, 8) 1484 94.54(10, 10)
+Cs $6p_{y}$ 364.80 279 77.18(6, 7) 812 92.18(8, 9) 1711 95.68(10, 11)
+Cs $6p_{z}$ 382.00 370 80.54(6, 8) 1016 93.40(8, 10) 2007 100.02(10, 12)
+Cs $6s$ 1455.30 475 91.32(6, 9) 1044 94.51(8, 11) 2321 104.14(10, 13)
+Cs $8s$ 172.00 595 92.54(6, 10) 1339 98.79(8, 12) 2714 114.07(10, 14)
+[…]
+Cs $7p_{x}$ 25.50 759 97.83(6, 12) 1808 112.65(8, 14) 3537 184.86(10, 16)
+Cs $7s$ 75.60 890 101.83(6, 13) 2100 156.01(8, 15) 3951 186.03(10,17)
+Cs $7s$ 85.50 1044 111.49(6, 14) 2358 182.96(8, 16) 4461 192.15(10, 18)
+Cs $5dy^{2}$ 307.30 1213 154.45(6, 15) 2624 184.12(8, 17)
+Cs $5d_{xy}$ 307.30 1338 181.34(6, 16) 2973 190.17(8, 18)
+Cs $5d_{xz}$ 309.70 1510 182.49(6, 17)
+Cs $5d_{yz}$ 309.70 1711 188.52(6, 18)
+Cs $5dz^{2}$ 310.90
+H $2s$ 882.50
+H $2p_{x}$ 1716.80
+H $2p_{y}$ 1716.80
+CsH H $2p_{z}$ 1716.80
+Table A11: Table providing the minimum and maximum eigenvalues of the matrix $\mathrm{A}_{mol}$ of size $m$, denoted in the table as $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ and $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ respectively, for KH, RbH and CsH systems.
+$n_{o}$ and $n_{v}$ denote the number of occupied and virtual orbitals respectively.
 The energies are in units of milli-Hartree.
-Molecule $m$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$
-RbH 40 135.31 20599.95
-119 135.18 20670.59
-228 135.04 20685.98
-407 131.94 20837.37
-889 109.40 21038.25
-1863 107.55 21058.36
-3866 92.36 21071.00
-4518 90.25 21294.40
-5225 72.72 22690.04
-5896 72.32 24341.29
-6595 71.93 24386.59
-7441 71.46 24701.91
-8238 71.27 24702.28
-9065 71.09 24702.65
-CsH 67 128.09 59680.83
-210 127.99 59722.33
-417 127.90 59737.34
-744 122.88 59847.85
-1171 110.54 59867.65
-1698 109.50 59867.64
-2175 108.61 59867.63
-2803 104.46 59867.67
-3493 100.89 59867.69
-4345 93.60 59916.79
-5215 92.75 59929.31
-6150 91.91 59939.41
-7262 83.88 59948.78
-8473 64.56 61759.13
-9784 63.38 61761.97
-11058 63.22 63448.99
-Table A11: Average PFDs and Range of PFDs of HHL and Psi-HHL for Li2, RbH and CsH molecules resp.
-in STO-6G, Sapporo-DZP and Sapporo-DKH3-DZP-2012 basis sets.
-The PFDs are calculated over a range of 5 repetitions.
-Li2 RbH CsH
-# shots HHL(PFD, Range) Psi-HHL(PFD, Range) HHL(PFD, Range) Psi-HHL(Range) HHL(PFD, Range) Psi-HHL(PFD, Range)
+Molecule $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
+m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$
+202 6 733.73 25918.07 580 8 730.05 29960.23 1484 10 127.91 280908.57
+274 7 731.91 29570.15 745 9 728.20 30271.74 1761 11 127.29 280909.73
+337 8 730.17 29604.25 932 10 127.95 30272.05 2046 12 126.64 280910.95
+432 9 728.35 29727.55 1110 11 127.33 30272.33 2451 13 120.90 280914.12
+541 10 128.01 29729.71 1296 12 126.71 30272.59 2900 14 98.56 280914.37
+655 11 127.38 29730.03 1540 13 120.92 30272.79 3393 15 96.78 281150.00
+748 12 126.76 29730.35 1808 14 98.59 30272.81 3794 16 95.80 281161.10
+889 13 120.98 29731.63 2100 15 96.83 30293.88 4203 17 94.84 281171.57
+1043 14 98.65 29731.69 2358 16 95.87 30304.25 4788 18 77.86 281200.60
+1213 15 96.89 29830.59 2624 17 94.89 30314.00 5265 19 77.45 281200.93
+1381 16 95.93 29834.52 2972 18 77.91 30317.70 5750 20 76.99 281201.42
+1510 17 94.96 29839.08 3278 19 77.50 30317.87 6427 21 76.56 281469.15
+1711 18 77.98 29854.11 3592 20 77.10 30318.05
+1909 19 77.57 29854.20 3996 21 76.62 30328.12
+2062 20 77.17 29854.30
+KH 2295 21 76.69 29894.66
+209 6 108.95 11193.43 632 8 108.02 11330.03 1401 10 106.22 11489.85
+279 7 108.73 11200.46 744 9 107.07 11330.11 1673 11 102.72 11490.43
+370 8 108.02 11206.98 932 10 106.25 11330.19 1963 12 99.68 11491.02
+432 9 107.08 11207.05 1110 11 102.75 11330.84 2321 13 91.77 11503.07
+541 10 106.26 11207.13 1296 12 99.70 11331.49 2713 14 81.20 12694.32
+630 11 102.77 11207.95 1540 13 91.80 11339.56 3141 15 72.02 13125.74
+748 12 99.73 11208.41 1808 14 81.23 12533.58 3537 16 71.65 14777.36
+889 13 91.84 11216.40 2100 15 72.05 12963.58 3951 17 71.28 14822.69
+1044 14 81.27 12410.75 2358 16 71.68 14614.13 4460 18 70.82 15148.61
+1213 15 72.10 12840.73 2624 17 71.31 14659.47 4932 19 70.64 15148.86
+1381 16 71.72 14490.85 2972 18 70.85 14988.70 5422 20 70.46 15149.12
+1510 17 71.35 14536.14 3278 19 70.67 14988.95 6013 21 70.05 15151.35
+1711 18 70.90 14865.97 3592 20 70.50 14989.21
+1860 19 70.72 14866.22 3996 21 70.09 14991.44
+2062 20 70.55 14866.47
+RbH 2295 21 70.15 14868.69
+209 6 119.60 9216.72 632 8 114.40 9306.7 1484 10 99.94 9447.99
+279 7 119.41 9216.73 812 9 101.3 9338.37 1711 11 98.74 9448.39
+370 8 114.52 9224.27 1016 10 99.97 9338.71 2007 12 94.46 9448.85
+475 9 101.35 9255.84 1044 11 98.81 9339.05 2321 13 90.73 9449.29
+595 10 100.02 9256.34 1339 12 94.53 9339.23 2714 14 82.84 9449.94
+664 11 98.86 9256.39 1540 13 90.80 9339.41 3141 15 63.54 10033.15
+759 12 94.61 9256.47 1808 14 82.91 9340.27 3537 16 63.38 11718.57
+890 13 90.89 9256.59 2100 15 63.60 9923.04 3951 17 63.23 11763.91
+1044 14 83.03 9257.34 2358 16 63.44 11609.02 4461 18 62.86 12079.72
+1213 15 63.71 9840.84 2624 17 63.29 11654.36
+1338 16 63.56 11526.12 2973 18 62.92 11966.65
+1510 17 63.40 11571.52
+CsH 1711 18 63.04 11884.73
+Table A12: Average PFDs and Range of PFDs of HHL and Psi-HHL for Li2 and KH molecules, both in Sapporo-DKH3-DZP-2012 basis set. The PFDs are calculated over a range of 30 repetitions.
+Li2 KH
+# shots HHL(PFD, Range) Psi-HHL(PFD, Range) HHL(PFD, Range) Psi-HHL(PFD, Range)
 $5\times 10^{2}$ (-7.68, 40.22) (-2.79, 29.58)
-$10^{3}$ (-4.08, 19.80) (-1.46, 8.34) (1.15, 30.41) (2.51, 25.93) (-3.76, 55.34) (0.49, 23.56)
-$5\times 10^{3}$ (-2.94, 8.89) (-2.11, 7.57)
-$10^{4}$ (-0.58, 6.00) (1.07, 4.47) (3.64, 25.05) (-0.42, 6.18) (-11.83, 9.82) (2.24, 4.22)
-$5\times 10^{4}$ (3.29, 16.90) (1.92, 3.14) (-12.38, 2.04) (1.63, 4.77)
-$10^{5}$ (5.77, 4.03) (0.60, 1.79) (-11.74, 2.49) (-0.06, 2.29)
-$5\times 10^{5}$ (5.29, 0.76) (0.99, 0.71) (-10.49, 2.29) (0.38, 1.77)
-$10^{6}$ (5.46, 0.72) (1.07, 0.56) (-10.92, 2.01) (0.56, 1.05)
+$10^{3}$ (-4.08, 19.80) (-1.46, 8.34) (nan, nan) (-11.98, 106.87)
+$5\times 10^{3}$ (-2.94, 8.89) (-2.11, 7.56)
+$10^{4}$ (-0.58, 6.00) (1.07, 4.47) (0.34, 68.36) (1.29, 44.70)
+$5\times 10^{4}$ (0.73, 32.51) (0.68, 16.78)
+$10^{5}$ (0.73, 17.00) (0.72, 9.02)
+$5\times 10^{5}$ (2.65, 10.59) (1.11, 5.65)
+$10^{6}$ (3.99, 5.09) (0.22, 4.05)
+Table A13: Table providing the details on the condition numbers ($\kappa$), the average total energy, the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the LCCSD total energy and the PFDs over 30 repetitions with respect to the bond length of KH molecule.
+All the energies are in units of milli-Hartree.
+Bond length (Bohr) $\kappa$ HHL Psi-HHL
+Total energy $\Delta E$ PFD Total energy $\Delta E$ PFD
+2.82 393.21 -598878.025 -0.124 -3.92 -598877.968 -0.068 -2.48
+3.29 374.93 -598940.871 -0.086 -3.84 -598940.844 -0.059 2.31
+3.76 367.32 -598967.447 -0.003 -1.96 -598967.447 -0.004 -0.26
+4.23 365.40 -598975.800 0.007 1.46 -598975.802 0.004 0.34
+4.71 371.38 -598974.517 -0.010 -10.00 -598974.509 -0.007 1.39
+5.18 385.16 -598968.421 -0.001 -0.44 -598968.419 $0$0.08
+5.65 373.56 -598960.199 -0.001 -0.41 -598960.196 0.000 0.26
+Figure A10: Figure showing average PFDs and range of PFDs (over 20 repetitions) for different values of $a$ for the overlap, $|\langle 0|(a|0\rangle+b|1\rangle)|$.
+We fix the number of shots to 1000 for generating all of the data.

### v3

@@
 The wrong signal is obtained by successfully post-selecting outcome bit ‘0’ instead of ‘1’ (we call this approach HHL1).
-The mixed signal is generated in an HHL quantum circuit by introducing an appropriately chosen $2\times 2$ unitary just before measurement of the HHL ancilla (we name this procedure HHL2).
+The mixed signal is generated from an HHL quantum circuit by introducing an appropriately chosen $2\times 2$ unitary just before measurement of the HHL ancilla (we name this procedure HHL2).
 Figure 1 presents schematic diagrams outlining the workings of the HHL and the Psi-HHL algorithms.
@@
 |$\ket{b}$ is the input state register with $n_{b}$ qubits, while $n_{r}$ refers to the number of clock register qubits.
-The Hong-Ou-Mandel (HOM) module is a controlled-SWAP test without an additional ancilla qubit, and we use it to extract |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ | (interpreted as the correlation energy, $E_{\mathrm{c}}^{\mathrm{HHL}}$, when HHL is applied to molecular calculations), which is the feature of |$\ket{x}$ that we are interested in.
+The Hong-Ou-Mandel (HOM) module is a destructive version of the controlled-SWAP test without an additional ancilla qubit, and we use it to extract $|\langle b|x\rangle|$, using which we evaluate |$-\|\ket{b}_{\rm un}\|^{2}\ \|\ket{x}_{\rm un}\|\ \ |\langle b|x\rangle|$ | (interpreted as the correlation energy, $E_{\mathrm{c}}^{\mathrm{HHL}}$, when HHL is applied to molecular calculations), which is the feature of |$\ket{x}$ that we are interested in.
+The subscript ‘un’ stands for un-normalized.
 Sub-figure (b) provides a block-diagram description of the HHL and the HOM modules.
@@
 In our notation, $E_{\mathrm{c}}^{\mathrm{HHL}}$ and $E_{\mathrm{c}}^{\mathrm{Psi-HHL}}$ are the correlation energies obtained from HHL and Psi-HHL respectively.
-Further, $a_{i},b_{i},c_{i},d_{i}\in\mathbb{R}$ are constants.
+Further, $a,b,c,d\in\mathbb{R}$ are constants.
 II Theory and Methodology
@@
 In the above expression, $\{\tilde{\lambda}_{i}\}$ refers to the estimates of the actual eigenvalues of $A$ ($1\geq\tilde{\lambda}_{i}\geq 1/\mathcal{\kappa}$) obtained via the QPE module of HHL.
-Typically, $C$ is chosen to be the minimum eigenvalue of $A$, but it is worth noting that incurring such a large classical overhead can always be circumvented via approaches such as AdaptHHL [2].
+Typically, $C$ is chosen to be the minimum eigenvalue of $A$, but it is worth noting that incurring the large classical overhead associated with computing it can always be circumvented via approaches such as AdaptHHL [2].
 When one post-selects the outcome ‘1’ on the HHL ancilla qubit, one can think of it as choosing the ‘correct’ signal.
@@
 In this section, we discuss the two essential ingredients— HHL1 and HHL2, which make up Psi-HHL as illustrated in Figure 1(c).
-The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the traditional HHL given in Figure 1(a) and Figure 1(b).
-The HHL2 entails traditional HHL module appended with a mixing $2\times 2$ unitary, given as $R_{Y}(2\alpha)$ and post-selecting 1 from the mixed signal.
+The HHL1 module depicts the case where the ‘wrong’ signal is post-selected from the HHL circuit, as shown in Figure 1(b) (Figure 1(a) illustrates the common case of post-selecting `$1$’).
+The HHL2 entails the traditional HHL module appended with a mixing $2\times 2$ unitary ($R_{Y}(2\alpha)$) on the HHL ancillary qubit just before measurement, and post-selecting `$1$’ from the ‘mixed’ signal.
 The role of this unitary is to mix the coefficients of $|0\rangle$ and $|1\rangle$ in the traditional HHL (see Eq.
@@
 The first HHL module, HHL1, serves to post-select the ‘wrong’ signal of the HHL ancilla qubit, that is, the state attached to |$\ket{0}$ in Eq.1.
-We re-write the state at the output of all the registers $|\psi_{\rm out}\rangle$ except the clock qubits register, as follows,
+We re-write the state, $|\psi_{\rm out}\rangle$, at the output of all the registers except the clock qubits register, as follows,
 […]
@@
 $E_{\rm c}^{\rm w}=-\||b\rangle_{\rm un}\|^{2}\ \||x_{\rm w}\rangle\|\ |\langle b|x_{0}\rangle|.$ (7)
-Here, $|b\rangle_{\rm un}$ is the unnormalized input vector after amplitude encoding, and
+In the expression given above, $|b\rangle_{\rm un}$ is the unnormalized input vector after amplitude encoding, and
 $|\langle b|x_{0}\rangle|=\sum_{i}|b_{i}|^{2}\sqrt{1-\frac{C^{2}}{\tilde{\lambda}_{i}^{2}}}.$ (8)
 II.2.2 HHL2: Computing the mixed signal
-The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancilla before the measurement as illustrated in the lower panel of Figure 1(c).
+The second HHL execution involves attaching the rotation gate $R_{Y}(2\alpha)$ on the HHL ancillary qubit before the measurement as illustrated in the lower panel of Figure 1(c).
 This seemingly trivial addition to the algorithm introduces a new degree of freedom via the angle $\alpha$, enabling one to boost the probability of successful matrix inversion in the case of large $\mathcal{\kappa}$.
@@
 Since for the purposes of deriving condition number behaviour at the measurement-post-selection part of HHL, the distinction between the estimated eigenvalues of $A$, that is, $\tilde{\lambda}_{i}$, and the actual eigenvalues, $\lambda_{i}$, does not matter, we denote all estimated eigenvalues in this section as simply $\lambda_{i}$ for the sake of convenience.
-In order to elaborate on the second part of our Proposition, we invoke the probability with which one post selects the desired outcome successfully in terms of $\kappa$.
+In order to elaborate on the second part of our Proposition, we invoke the probability with which one post-selects the desired outcome successfully in terms of $\kappa$.
 a) Success probability of the post-selection process in HHL1: The probability of successfully post-selecting the outcome ‘0’, $P(0)$, in HHL1 as given in Eq. 6 can be expanded as [20]
@@
 Using once again the sub-multiplicativity of matrix norms in Eq. 18, we can write
-$P^{\prime}(1)$$\leq$ $\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
+$P^{\prime}(1)$$\leq$ $\rm\|[\sqrt{\mathbb{I}-C^{2}A^{-2}}\ sin(\alpha)$ (21)
 $+$ $CA^{-1}\cos(\alpha)]\|^{2}_{\rm op}\||b\rangle\|^{2},$
@@
 This implies that fewer measurements are required to invert matrices with large condition numbers, as compared to the traditional HHL technique illustrated in Figure 1(a).
-We show this in Figure 2, where for a given $\alpha=60^{\circ}$, we obtain
+We obtain
 $\rm\bigg{|}\bigg{[}\bigg{(}\sqrt{1-\frac{\gamma^{2}}{\mathcal{\kappa}^{2}}}\bigg{)}sin(\alpha)+\frac{\gamma}{\mathcal{\kappa}}cos(\alpha)\bigg{]}\bigg{|}^{2}$$\leq$ $P^{\prime}(1)$ (23)
@@
 $+$ ${\gamma}\mathrm{cos}(\alpha)\bigg{]}\bigg{|}^{2}.$
+We show this in Figure 2, where we pick $\alpha=60^{\circ}$, the choice of $A$ matrix is $A=\begin{pmatrix}0.25&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, and with six different choices of $\vec{b}$ (the eigenvector associated with the minimum eigenvalue of $A$, the eigenvector associated with the maximum eigenvalue of $A$, eigenvector associated with two other eigenvalues of $A$, equal superposition of all eigenstates of $A$, and an arbitrary superposition of eigenstates of $A$).
 While the $A$ and $B$ matrices share the same basis {|$\{\ket{\nu_{i}}\}$, an eigenstate of $B$ matrix that achieves the lower bound for $P^{\prime}(1)$ may not often correspond to the largest/smallest eigenvalue of $A$ matrix, for a given choice of angle $\alpha$.
@@
 The HHL1 of Psi-HHL successfully post-selects outcome 00 with probability $P(0)=1$.
-On the other hand, the HHL2 module of Psi-HHL post-selects outcome $1$ with probability $\sin^{2}(\alpha)$, see Eq.
-10.
+On the other hand, the HHL2 module of Psi-HHL post-selects outcome $1$ with probability $\sin^{2}(\alpha)$ (see Eq.
+10).
 The data points presented in Figure 3 demonstrate these probabilities for the two examples considered.
@@
 In this section, we consider two sets of examples to compare the performance of Psi-HHL with that of HHL:
-  • Toy $A$ matrices of sizes $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$: For practical reasons associated with computational resources, we do not choose matrices whose $\mathcal{\kappa}$ grows with the $A$ matrix size.
+  • Toy $A$ matrices of sizes $4\times 4$, $8\times 8$, $16\times 16$, $32\times 32$, and $64\times 64$: For practical reasons associated with (classical) computational resources, we do not choose matrices whose $\mathcal{\kappa}$ grows with the $A$ matrix size.
 Instead, we generate a sequence of $A$ matrices with systematically increasing condition numbers for every system size.
@@
       Additionally, for the chemistry case, we compare the performance of HHL and Psi-HHL across a limited range of geometries on a potential energy curve (PEC) for a suitably chosen molecule.
-All of our simulation results are obtained by building on the Qiskit program [21].
+All of our simulations are developed and executed on the Qiskit software development kit [21].
 Refer to caption Refer to caption Refer to caption
@@
 Figure 4: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for toy matrices.
-Sub-figures whose top panels present the average probabilities of HHL and Psi-HHL with varying values of $\mathcal{\kappa}$ for different $A$ matrix sizes: $4\times 4$ (sub-figure (a)), $8\times 8$ (sub-figure (b)), $16\times 16$ (sub-figure (c)), $32\times 32$ (sub-figure (d)), and $64\times 64$ (sub-figure (e)).
+The top panels in each sub-figure present the average probabilities of HHL and Psi-HHL with varying values of $\mathcal{\kappa}$ for different $A$ matrix sizes: $4\times 4$ (sub-figure (a)), $8\times 8$ (sub-figure (b)), $16\times 16$ (sub-figure (c)), $32\times 32$ (sub-figure (d)), and $64\times 64$ (sub-figure (e)).
 The bottom panel of each sub-figure represents the average Percentage Fraction Differences (PFD) of the overlap, $E_{c}$, calculated with HHL and Psi-HHL with respect to the classical value versus $\mathcal{\kappa}$.
-Each data point results from 10 repeated runs of $10^{6}$ shots each.
-Also, all the computations have been done with $A$ not diagonal and $\vec{b}$ in unequal superposition, as described in the main text.
+Each data point results from 10 repeated runs of $10^{6}$ shots.
+Also, all the computations have been carried out with $A$ not diagonal and $\vec{b}$ in unequal superposition, as described in the main text.
 We also note here that the approximate equality is due to the fit coefficients in the figure being rounded off to two decimal places.
+For the Psi-HHL results, we set $\alpha$ to be 60 degrees.
 III.1 Toy matrices: $4\times 4$ through $64\times 64$
@@
 and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$.
-The sparsity, $s$, of $A$ is defined as the number of non-zero entries in the row with the most number of non-zero elements and is set to 2.
+The sparsity, $s$, of $A$ is defined as the number of non-zero entries in the row with the most number of non-zero elements, and is set to 2.
 Our choices for the other matrices, $8\times 8$ through $64\times 64$ can be found in Figure A1 of the Appendix.
@@
 Each data point in the sub-figures is the average result obtained over 10 repetitions, with each repetition involving 1 million shots.
-We verify that the observed trends do not change with 50 repetitions by picking the $4\times 4$ matrix as a representative example (Figures A2 and A3 of Supplemental Material).
-We pick $\alpha=60^{\circ}$ for the Psi-HHL calculations.
+For the Psi-HHL results, we set $\alpha$ to be 60 degrees.
+Each sub-figure provides two sets of data: one which compares the probabilities, $P(0)$, $P(1)$, $P^{\prime}(0)$, and $P^{\prime}(1)$, obtained by executing HHL and Psi-HHL, and the other providing the PFDs for the two methods.
+We choose condition numbers all the way till about 1 million.
+The figures show that our data fits for HHL align with the expected behaviour of $P(1)\propto\mathcal{\kappa}^{-2}$.
+We also find that $P^{\prime}(1)$ scales as $\propto 1$ for large condition numbers (with the fit parameter $a_{1}=0.75$ matching with the Y-axis intercept for large condition number values).
+We also consider two more cases, where $A$ is diagonal but $\vec{b}$ is in an equal superposition, and $A$ diagonal and $\vec{b}$ in an unequal superposition.
+Furthermore, we also consider the situations where $\alpha$ is set to 70 and 80 degrees.
+The data for all of these cases are presented in Figures A2 and A3 of the Appendix, accompanied by Tables A1 through A3.
+We also verify that the observed trends do not change with 50 repetitions, for which we pick the $4\times 4$ matrix (with $\alpha=60^{\circ}$ for the Psi-HHL calculations) from Figure 4 as a representative example (see Figures A4 and A5 of the Appendix).
 The corresponding data is given in Table A4 of the Appendix.
-Our results obtained using other choices of angles, $70^{\circ}$ and $80^{\circ}$, are provided in Tables A2 and A3 respectively.
-We now return to Figure 4, where each sub-figure provides two sets of data: one which compares the probabilities, $P(0)$, $P(1)$, $P^{\prime}(0)$, and $P^{\prime}(1)$, obtained by executing HHL and Psi-HHL, and the other providing the PFDs for the two methods.
-We choose condition numbers all the way till about 1 million.
-Figures A2 and A3 of the Appendix shows these results for $\alpha$ of $60^{\circ}$, $70^{\circ}$, and $80^{\circ}$, with $10^{5}$ and $10^{6}$ shots respectively.
-The figures show that our data fits for HHL align with the expected behaviour of $P(1)\sim\mathcal{O}(\mathcal{\kappa}^{-2})$.
-We also find that $P^{\prime}(1)$ scales as $\sim\mathcal{O}(1)$ for large condition numbers (with the fit parameter $a=0.75$ matching with the Y-axis intercept for large condition number values).
-For completeness, we verify that the fit probabilities add to one (see Table A4 of the Appendix).
+Results obtained using other choices of angles, $70^{\circ}$ and $80^{\circ}$, are provided in Tables A5 and A6 respectively.
+For completeness, we also verify that the fit probabilities from Figure 4 add to one (see Table A7 of the Appendix).
 Refer to caption Refer to caption
@@
 Figure 5: Figure showing a comparison of the spread in PFDs as well as the average PFD (over 10 repetitions), obtained using HHL (sub-figure (a)) and Psi-HHL (sub-figure (b)), against a range of three different numbers of shots and for the case where $A$ is of size $4\times 4$, and for which $\mathcal{\kappa}\approx 2^{12}=4096$.
-We now comment on the PFDs.
-We immediately see that HHL fails at some value of $\mathcal{\kappa}$ for the chosen number of shots (indicated by ‘nan’ values in Tables A1, A2, and A3 of the Appendix).
+We now comment on the observed PFDs from Figure 4.
+We immediately see that HHL fails at some value of $\mathcal{\kappa}$ for the chosen number of shots (indicated by ‘nan’ values in Tables A1 through A6 of the Appendix).
 This is due to the fact that as $P(1)$ becomes very small, the fraction of the total number of shots that yield 1 upon measurement tend to very small values.
 Therefore, the number of shots available for the HOM module become too small to give any reasonable result beyond a certain value of $\mathcal{\kappa}$.
-The bottom panel of each sub-figure of Figure 4 shows that the error bars, in which the upper and lower points of the bars reflect the largest and the smallest PFD values obtained using HHL, increases as the condition number increases, till it reaches a critical value of $\mathcal{\kappa}$, beyond which HHL fails.
+Our data also show that the error bars, in which the upper and lower points of the bars reflect the largest and the smallest PFD values obtained using HHL, increases as the condition number increases, till it reaches a critical value of $\mathcal{\kappa}$, beyond which HHL fails.
 In contrast, we see that for Psi-HHL, even at $\mathcal{\kappa}\ \approx$ 1 million, the length of the error bar is about 17 percent.
-We now verify that HHL does not fail upon increasing the number of shots, but the number requires to be about 1 billion in order to get a comparable range of PFDs as Psi-HHL, while Psi-HHL only incurs 10 million shots, as Figure 5 shows.
-In particular, HHL gives a maximum PFD of 1.51 percent and a minimum of $-2.98$ percent with a billion shots, whereas Psi-HHL gives a maximum and minimum PFD of 2.33 and $-2.97$ percent respectively with only 10 million shots.
 We finally comment on our results from Figure 4 on the larger size matrices ($8\times 8$ through $64\times 64$), where for Psi-HHL, we consider $\alpha=60^{\circ}$.
-Figures A4 through A7 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A4 also giving the information for $10^{5}$ shots.
-Tables A5 through A8 in the Appendix provide the accompanying data.
+Figures A6 through A9 provides the data for $70^{\circ}$ and $80^{\circ}$ cases, with Figure A6 also giving the information for $10^{5}$ shots.
+Tables A8 through A11 in the Appendix provide the accompanying data.
 We observe that broadly, the trends are similar to what we find for the smaller $4\times 4$ cases.
@@
 In view of the computational cost involved, we did not go all the way up to a condition number of $2^{20}$ for matrices above $4\times 4$ size, and we stopped much earlier for the $64\times 64$ case.
-Refer to caption Refer to caption Refer to caption
+We now turn our attention to Figure 5, where we see that HHL does not fail upon increasing the number of shots, but the number requires to be about 1 billion in order to get a comparable range of PFDs as Psi-HHL, while Psi-HHL only incurs 10 million shots.
+In particular, HHL gives a maximum PFD of 1.51 percent and a minimum of $-2.98$ percent with a billion shots, whereas Psi-HHL gives a maximum and minimum PFD of 2.33 and $-2.97$ percent respectively with only 10 million shots.
       (a)                                               (b)                                                 (c)
 Figure 6: Plots showing the increase in $\mathcal{\kappa}$ with $A_{mol}$ matrix size, denoted as $m$, for the case of 6 (sub-figure (a)), 8 (sub-figure (b)) and 10 (sub-figure (c)) fixed occupied orbitals in the Sapporo-DKH3-DZP-2012 basis set.
-The left-most legend on top of sub-figure (a) shows black circles whose areas increase proportionally to the number of virtual orbitals considered.
+The legend to the left at the top of the figure shows solid circles whose areas increase proportionally to the number of virtual orbitals considered.
 III.2 Quantum chemistry matrices
@@
 This in turn requires us to define system size.
-Since our focus is to demonstrate the performance of Psi-HHL and not a deep-dive into chemistry, we take the approach of: a.
+Since our focus is to demonstrate the performance of Psi-HHL and not a deep-dive into chemistry, we follow the following steps: a.
 choosing a few simple diatomic closed-shell molecules in their ground state and in their equilibrium bond lengths, b.
@@
 systematically increasing the number of unoccupied orbitals (virtuals) for system size.
-We choose four representative molecules: the light Li2, the moderately heavy KH and RbH, and the heavy CsH neutral diatomic molecules, all in their equilibrium bond lengths (Li2: 5.0600 Bohr, KH: 4.2300 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr) [23, 24].
+We choose four representative molecules: the light Li2, the moderately heavy KH and RbH, and the heavy CsH, all in their equilibrium bond lengths (Li2: 5.0600 Bohr, KH: 4.2300 Bohr, RbH: 4.4726 Bohr, and CsH: 4.7135 Bohr) [23, 24].
 The adjective ‘light’ and ‘heavy’ refer to the atomic number of the heavier atom in a diatomic molecule.
@@
 III.2.1 Dimension of the $A$ matrix
-Since the LCC matrix is not guaranteed to be of dimension $2^{N}\equiv n$, we make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$.
+Since the LCC matrix is not guaranteed to be of dimension $2^{N}\times 2^{N}\equiv n\times n$, we make a distinction between the dimension of the LCC matrix, $A_{mol}$, which is $m\times m$, and the dimension of the associated $A$ matrix, $n\times n$.
 We go from $A_{mol}$ to $A$ via padding, which we explain in Section A2 of the Appendix.
@@
 We denote the number of occupied orbitals as $n_{o}$ and virtuals by $n_{v}$.
-The $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$ (the number of orbitals are kept fixed), which is exponential.
+The number of rows (or columns) of the $A_{mol}$ matrix grows as $\sim n_{o}^{\mathcal{E}}n_{v}^{\mathcal{E}}$ (the number of orbitals are kept fixed), which is exponential.
 $\mathcal{E}$ refers to the excitation rank.
 For example, $\mathcal{E}=2$ is the LCCSD approximation, where S stands for single excitations and D for double excitations.
-For a given $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially, for example, see Figure A9 in the Appendix.
-For the purposes of this work, we study the size of $A_{mol}$ with increase in number of unoccupied orbitals, $n_{v}$, in a molecule, while keeping $n_{o}$ fixed at a reasonable number, all within the LCCSD approximation.
+For a given $\mathcal{E}$, as we increase the number of orbitals, the system size, $m$, grows further polynomially (for example, see Figure A11 in the Appendix).
+For the purposes of this work, we fix $n_{o}$ and increase $n_{v}$ for a given molecule, all within the LCCSD approximation.
 The technical details involved in the generation of $A$ matrix as well as the $\vec{b}$ vector for our computations are discussed in Section A2.A of the Appendix.
 III.2.2 Condition number scaling in LCC
-Although HHL offers an exponential advantage in terms of system size, it is important to check if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the systems that we considered.
+Although HHL offers an exponential advantage in terms of system size, it is important to check if for the LCCSD problem, the condition number itself scales favourably with system size, $m$, for the molecules that we consider.
 We point out towards the end of Section A2.A of the Appendix that it is sufficient to analyze $\mathcal{\kappa}$ versus $m$; this curve upper bounds $\mathcal{\kappa}$ versus $n$.
@@
 Furthermore, for this analysis, we consider KH, RbH and CsH, but not Li2, as the system is too small (in the STO-6G basis) to carry out a $\mathcal{\kappa}$ versus $m$ analysis.
-We start with $(n_{o},n_{v})$ of (6, 6), (8, 8), and (10, 10) for all the three molecules, and for each of these cases, we systematically increase the number of virtuals.
+We start with $(n_{o},n_{v})$ of (6, 6), (8, 8), and (10, 10) for all the three molecules, and for each of these cases, we systematically increase $n_{v}$.
 Figure 6 presents our results.
@@
 The data shows that the increase in $\mathcal{\kappa}$ as we add virtual orbitals is driven by the decrease of the lowest eigenvalue of the $A$ matrix.
-We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock configuration.
-The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the Hartree-Fock (HF) configuration).
+We note that every matrix element of the $A$ matrix can be expressed as $\langle\Phi_{p}|H|\Phi_{q}\rangle$, where $H$ is the Hamiltonian and |$\ket{\Phi_{p}}$ and |$\ket{\Phi_{q}}$ are the excited determinants built from particle-hole excitations arising from the Hartree-Fock (HF) configuration.
+The smallest eigenvalue of the sparse and diagonally dominant $A$ matrix, which we saw as being important in determining the growth of $\mathcal{\kappa}$, can be qualitatively thought of as corresponding to the energy of the first excited state (assuming that the wave function of the first excited state has small contribution from the HF configuration).
 Thus, we expect that the decrease in the lowest eigenvalue of $A$ saturates as we keep adding virtual orbitals, and eventually, the lowest eigenvalue of the $A$ matrix will converge to a certain value.
 Thus, the increase of $\mathcal{\kappa}$ saturates too, thus leading one to expect that the condition number would vary as a polylogarithmic function of system size.
-We also observe in each sub-figure (that is, for a fixed number of occupied orbitals), for large $m$, the condition number of KH is the largest, followed by RbH, and then CsH.
+Figure 7: Figure showing a comparison of the performance of HHL and Psi-HHL approaches, via PFDs versus number of shots (denoted in the figure as ‘No.
+of shots’), for two representative molecules, Li2 (sub-figure (a)) and KH (sub-figure (b)).
+The centre of each unfilled circle represents the average PFD over a range of 30 repetitions.
+Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$ ($\mathcal{\kappa}$ of 64.07), while for KH, we evaluate a 26-qubit circuit involving an $A$ matrix size of $256\times 256$ ($\mathcal{\kappa}$ of 365.40).
+The zero PFD value is marked with a horizontal line for ease of visualization.
+We also observe in each sub-figure (that is, for a fixed number of occupied orbitals) that for large $m$, the condition number of KH is the largest, followed by RbH, and then CsH.
 This trend can be explained qualitatively using chemical intuition, by looking at the considered orbitals for the (6, 6) case as a representative example.
-Each eigenvalue of $A$ can be thought of as an excited state energy, given that $A$ is sparse and diagonally dominant.
-The highest excited state arises from electron excitations from the lowest occupied orbitals, which are $2p$ and $3d$ for KH and RbH cases respectively.
+Recalling that each eigenvalue of $A$ can be thought of as an excited state energy, the highest excited state arises from electron excitations from the lowest occupied orbitals, which are $2p$ and $3d$ for KH and RbH cases respectively.
 It is clear that $2p$ of K has much lower orbital energy than the $3d$ orbital of Rb.
@@
 Furthermore, in Figure 6(c), KH’s condition number reaches a staggering 4000 with all the virtuals considered, whereas in Figures 6(a) and (b), $\mathcal{\kappa}$ is substantially lower.
-This is not surprising, as for sub-figure 6(c), we consider 10 occupied orbitals, and KH has only 10 occupied orbitals and thus the active space contains the $1s$ orbital of K.
+This is not surprising, as for sub-figure 6(c), we consider 10 occupied orbitals.
+Since KH has only 10 occupied orbitals, the active space contains the $1s$ orbital of K.
 We can invoke the fact that the orbital energy gap is larger between the inner s orbitals, for example, and the density of orbitals becomes higher in the vicinity of highest occupied molecular orbital and the lowest unoccupied molecular orbital.
-Thus, if we include inner core orbitals in our active space, we expect that the excitation energy of the highest state, namely the largest eigenvalue of the $A$ matrix, to be larger.
+Thus, if we include inner core orbitals in our active space, we expect the excitation energy of the highest state, namely the largest eigenvalue of the $A$ matrix, to be larger.
 In contrast, for the 6, 8, and the 10 occupied orbitals’ cases, RbH and CsH have comparable condition numbers.
@@
 Since the orbital character of the lowest orbital is the same among the three active spaces, they yield very similar condition numbers.
-Figure 7: Figure showing a comparison of the performance of HHL and Psi-HHL approaches for two representative molecules, via PFDs of Li2 (sub-figure (a)) and KH (sub-figure (b)) molecules versus number of shots (denoted in the figure as ‘No.
-of shots’).
-The centre of each unfilled circle represents the average PFD over a range of 30 repetitions.
-Each Li2 calculation involves evaluating a 21-qubit circuit, with the $A$ matrix size being $64\times 64$ ($\mathcal{\kappa}$ of 64.07), while for KH, we evaluate a 26-qubit circuit involving an $A$ matrix size of $256\times 256$ ($\mathcal{\kappa}$ of 365.40).
-The zero PFD value is marked with a horizontal line for ease of visualization.
-Figure 8: Figure illustrating the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the classical value of the LCCSD total energy (sub-figure (a)) and PFDs (sub-figure (b)) vs bond lengths of the KH molecule obtained from HHL and Psi-HHL methods (26-qubit calculations involving $A$ matrix size of $256\times 256$; red and blue circles indicate HHL and Psi-HHL calculations respectively, averaged over 30 repetitions and with 50000 shots) and those computed with the classical LCC value.
-The associated value of $\mathcal{\kappa}$ is shown just above or below each data point.
-We add at this juncture that while considering only three molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may a reasonable direction to consider.
+We add at this juncture that while considering only three molecules does not allow any quantitative inference on the behaviour of condition number with system size, it gives a qualitative indication that quantum chemistry using HHL-LCCSD may be a reasonable direction to consider.
 However, further analysis would be required to assess what one would expect for $\mathcal{\kappa}(m)$ in general for an HHL-LCC computation.
@@
 We now discuss our results for the performance of the HHL and Psi-HHL algorithms for the molecules considered in this work.
-We begin by conducting a study similar to the one shown in Figure 5 for the toy matrices.
-It is not practical to carry out an analysis similar to that in Figure 4, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon.
-We next move to Figure 7 (with the accompanying data given in Table A12 of the Appendix), where we vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 30 repetitions.
+We begin by noting that it is not practical to carry out an analysis similar to that in Figure 4, since unlike the toy matrices case, the condition number grows much slower and the computations become too expensive too soon.
+We thus move to Figure 7 (with the accompanying data given in Table A15 of the Appendix), where we carry out a study on the same lines as Figure 5.
+We vary the number of shots and compare the performance of HHL and Psi-HHL in predicting the PFD, as well as the range across which the PFDs are spread, across 30 repetitions.
 For this analysis, we consider Li2 and KH.
@@
     KH: 26-qubit circuit with an $A$ matrix size of $256\times 256$ (built out of single and double excitations arising from 6 occupied and 6 virtual orbitals) in the Sapporo double zeta basis, with $\mathcal{\kappa}=365.40$.
-We pick these two systems, as in one case, $\mathcal{\kappa}$ is less than 100, and in the other much larger at over 300.
+We pick these two systems, since in one case, $\mathcal{\kappa}$ is less than 100, and in the other much larger at over 300.
 Repeating the calculations for RbH and CsH may not give any new insights, and hence we restrict the results only to Li2 and KH.
@@
 However, since we keep the number of shots fixed throughout the PEC, we observe a larger spread in PFDs.
-We verify this with a simple toy calculation, where we calculate the overlap $|\langle 0|(a|0\rangle+b|1\rangle)|$ using HHL, with 1000 shots and 20 repeats.
-We average our results over 20 repetitions.
-$a$ is varied from 0.26 to 0.99 (below 0.26, we get nan).
+We verify this with a simple toy calculation, where we calculate the overlap $|\langle 0|(a|0\rangle+b|1\rangle)|$ using the HOM circuit, with 1000 shots and average each of our results over 20 repetitions.
+$a$ is varied from 0.26 to 0.99 (below 0.26, we get ‘nan’ that is, situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result).
 We observe the same trend as in Figure 8(b) (see Figure A12 of Appendix).
+Figure 8: Figure illustrating the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the classical value of the LCCSD total energy (sub-figure (a)) and PFDs versus bond lengths (sub-figure (b)) of the KH molecule obtained from HHL and Psi-HHL methods (26-qubit calculations involving $A$ matrix size of $256\times 256$; red and blue circles indicate HHL and Psi-HHL calculations respectively, averaged over 30 repetitions and with 50000 shots).
+The associated value of $\mathcal{\kappa}$ is shown above or below each data point.
 IV Conclusion
@@
 Debashis Mukherjee for useful discussions on the quantum chemical sections of the work.
-KS acknowledges support from Quantum Leap Fragship Program (Grant No.
+VSP and PBT thank their group members for carefully reading the manuscript.
+KS acknowledges support from Quantum Leap Flagship Program (Grant No.
 JPMXS0120319794) from Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan; Center of Innovations for Sustainable Quantum AI (JPMJPF2221) from Japan Science and Technology Agency (JST), Japan; and Grants-in-Aid for Scientific Research C (21K03407) and for Transformative Research Area B (23H03819) from Japan Society for the Promotion of Sciences (JSPS), Japan.
@@
 Figure A1: Our choices for (a) $8\times 8$, (b) $16\times 16$, (c) $32\times 32$, and (d) $64\times 64$ toy $A$ matrices and un-normalized $\vec{b}$ for comparing the performance of HHL and the Psi-HHL algorithms.
-The cell with a star symbol denotes the matrix element with the variable $n_{r}$ (the number of clock register qubits from the HHL circuit), which sets $\mathcal{\kappa}$.
+The cell with a star symbol denotes the matrix element with variable $n_{r}$ (the number of clock register qubits from the HHL circuit), which sets $\mathcal{\kappa}$.
 A white cell denotes a zero matrix element.
+Refer to caption Refer to caption Refer to caption
+(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
+Refer to caption Refer to caption Refer to caption
+(d) $\alpha=60^{\circ}$ (e) $\alpha=70^{\circ}$ (f) $\alpha=80^{\circ}$
+Refer to caption Refer to caption Refer to caption
+(g) $\alpha=60^{\circ}$ (h) $\alpha=70^{\circ}$ (i) $\alpha=80^{\circ}$
+Figure A2: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{5}$ shots throughout.
+Sub-figures (a) through (c) consider the case where $A$ is diagonal and $\vec{b}$ is in equal superposition.
+Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal, and $\vec{b}$ is in unequal superposition.
+Each data point is generated by taking the average over 10 repetitions.
+Refer to caption Refer to caption Refer to caption
+(a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
+Refer to caption Refer to caption Refer to caption
+(d) $\alpha=60^{\circ}$ (e) $\alpha=70^{\circ}$ (f) $\alpha=80^{\circ}$
+Refer to caption Refer to caption Refer to caption
+(g) $\alpha=60^{\circ}$ (h) $\alpha=70^{\circ}$ (i) $\alpha=80^{\circ}$
+Figure A3: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $4\times 4$ matrices and using $10^{6}$ shots throughout.
+Sub-figures (a) through (c) consider the case where $A$ is diagonal and $\vec{b}$ is in equal superposition.
+Sub-figures (d) through (f) consider the case where $A$ is diagonal and $\vec{b}$ is in unequal superposition, and Sub-figures (g) through (i) consider the case where $A$ is not diagonal, and $\vec{b}$ is in unequal superposition.
+Each data point is generated by taking the average over 10 repetitions.
+Table A1: Table presenting the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$) for $4\times 4$ matrices, and in particular, for three cases: $A$ diagonal and $\vec{b}$ in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ equal’), $A$ diagonal and $\vec{b}$ not in equal superposition (denoted in the table as ‘$A$ diag, $\vec{b}$ unequal’), and $A$ not diagonal and $\vec{b}$ in an unequal superposition (denoted in the table as ‘$A$ not diag, $\vec{b}$ unequal’).
+The string ‘nan’ refers to situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result.
+Each data point is generated by taking the average over 10 repetitions.
+’# shots’ refers to the number of shots, while ‘# qubits’ denotes the number of qubits.
+$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
+# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1))$ PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$)
+$2^{2}$ 8 -0.44 (0.65, 0.35) -0.22 (0.21, 0.79) -1.01 (0.92, 0.08) -0.10 (0.06, 0.94) -0.87 (0.92, 0.08) -0.04 (0.08, 0.92)
+$2^{3}$ 9 -0.55 (0.72, 0.28) -0.07 (0.27, 0.73) -5.82 (0.97, 0.03) -0.32 (0.15, 0.85) -5.72 (0.97, 0.03) 0.08 (0.15, 0.85)
+$2^{4}$ 10 -0.81 (0.74, 0.26) 0.32 (0.32, 0.68) 14.39 (0.99, 0.01) -0.69 (0.20, 0.80) 17.20 (0.99, 0.01) 0.26 (0.20, 0.80)
+$2^{5}$ 11 -0.17 (0.75, 0.25) -0.02 (0.35, 0.65) 3.96 (0.99, 0.01) 2.85 (0.23, 0.77) -0.12 (0.99, 0.01) 0.32 (0.22, 0.78)
+$2^{6}$ 12 0.30 (0.75, 0.25) 0.05 (0.37, 0.63) -6.73 (0.99, 0.01) 1.66 (0.24, 0.76) -6.73 (0.99, 0.01) -1.56 (0.24, 0.76)
+$2^{7}$ 13 nan (0.75, 0.25) -0.45 (0.37, 0.63) nan (0.99, 0.01) 1.91 (0.25, 0.75) nan (0.99, 0.01) -2.62 (0.25, 0.75)
+$2^{8}$ 14 nan (0.75, 0.25) -0.01 (0.37, 0.63) nan (0.99, 0.01) -2.00 (0.25, 0.75) nan (0.99, 0.01) -3.87 (0.25, 0.75)
+$2^{9}$ 15 nan (0.75, 0.25) -0.22 (0.37, 0.63) nan (0.99, 0.01) -7.50 (0.25, 0.75) nan (0.99, 0.01) 8.02 (0.25, 0.75)
+$2^{10}$ 16 nan (0.75, 0.25) -0.26 (0.37, 0.63) nan (0.99, 0.01) -2.33 (0.25, 0.75) nan (0.99, 0.01) -7.77 (0.25, 0.75)
+$2^{11}$ 17 nan (0.75, 0.25) 0.08 (0.37, 0.63) nan (0.99, 0.01) 2.65 (0.25, 0.75) nan (0.99, 0.01) -0.55 (0.25, 0.75)
+$2^{12}$ 18 nan (0.75, 0.25) 0.03 (0.37, 0.63) nan (0.99, 0.01) 7.93 (0.25, 0.75) nan (0.99, 0.01) -2.41(0.25, 0.75)
+$2^{13}$ 19 nan (0.75, 0.25) 0.30 (0.38, 0.62) nan (0.99, 0.01) 4.67 (0.25, 0.75) nan (0.99, 0.01) 3.12 (0.25, 0.75)
+$2^{14}$ 20 nan (0.75, 0.25) 0.41 (0.38, 0.62) nan (0.99, 0.01) 9.11 (0.25, 0.75) nan (0.99, 0.01) -5.60 (0.25, 0.75)
+$2^{15}$ 21 nan (0.75, 0.25) 0.13 (0.38, 0.62) nan (0.99, 0.01) 9.42 (0.26, 0.74) nan (0.99, 0.01) -1.11 (0.25, 0.75)
+$2^{16}$ 22 nan (0.75, 0.25) 0.54 (0.38, 0.62) nan (0.99, 0.01) -0.25 (0.25, 0.75) nan (0.99, 0.01) 2.43 (0.25, 0.75)
+$2^{17}$ 23 nan (0.75, 0.25) -0.03 (0.38, 0.62) nan (0.99, 0.01) -0.40 (0.26, 0.74) nan (0.99, 0.01) -7.87 (0.25, 0.75)
+$2^{18}$ 24 nan (0.75, 0.25) 0.26 (0.37, 0.63) nan (0.99, 0.01) -8.56 (0.26, 0.74) nan (0.99, 0.01) 9.69 (0.26, 0.74)
+$2^{19}$ 25 nan (0.75, 0.25) 0.19 (0.37, 0.63) nan (0.99, 0.01) 6.91 (0.26, 0.74) nan (0.99, 0.01) 1.06 (0.25, 0.75)
+$10^{5}$ $2^{20}$ 26 nan (0.75, 0.25) -0.58 (0.37, 0.63) nan (0.99, 0.01) 0.30 (0.25, 0.75) nan (0.99, 0.01) 6.15 (0.25, 0.75)
+$2^{2}$ 8 -0.57 (0.65, 0.35) 0.00 (0.21, 0.79) -0.87 (0.92, 0.08) 0.03 (0.08, 0.92) -0.88 (0.92, 0.08) 0.04 (0.08, 0.92)
+$2^{3}$ 9 -0.61 (0.72, 0.28) -0.14 (0.27, 0.73) -5.82 (0.97, 0.03) -0.04 (0.15, 0.85) -5.82 (0.97, 0.03) 0.00 (0.15, 0.85)
+$2^{4}$ 10 -0.63 (0.74, 0.26) -0.04 (0.31, 0.69) 14.66 (0.99, 0.01) 0.11 (0.20, 0.8) 14.95 (0.99, 0.01) -0.15 (0.20, 0.80)
+$2^{5}$ 11 0.27 (0.75, 0.25) -0.06 (0.34, 0.66) 1.84 (0.99, 0.01) 0.16 (0.22, 0.78) 0.61 (0.99, 0.01) 0.44 (0.22, 0.78)
+$2^{6}$ 12 0.04 (0.75, 0.25) 0.00 (0.36, 0.64) -0.72 (0.99, 0.01) 0.01 (0.24, 0.76) 1.68 (0.99, 0.01) -0.70 (0.24, 0.76)
+$2^{7}$ 13 -0.07 (0.75, 0.25) 0.20 (0.37, 0.63) 1.35 (0.99, 0.01) -1.56 (0.25, 0.75) -4.87 (0.99, 0.01) -0.11 (0.25, 0.75)
+$2^{8}$ 14 -0.13 (0.75, 0.25) -0.02 (0.37, 0.63) 0.09 (0.99, 0.01) -0.04 (0.25, 0.75) -4.04 (0.99, 0.01) -1.15 (0.25, 0.75)
+$2^{9}$ 15 0.05 (0.75, 0.25) 0.06 (0.37, 0.63) -2.82 (0.99, 0.01) 2.82 (0.25, 0.75) nan (0.99, 0.01) 2.26 (0.25, 0.75)
+$2^{10}$ 16 0.04 (0.75, 0.25) -0.04 (0.37, 0.63) nan (0.99, 0.01) -0.12 (0.25, 0.75) nan (0.99, 0.01) -0.59 (0.25, 0.75)
+$2^{11}$ 17 0.19 (0.75, 0.25) -0.05 (0.37, 0.63) nan (0.99, 0.01) -1.49 (0.25, 0.75) nan (0.99, 0.01) 1.41 (0.25, 0.75)
+$2^{12}$ 18 0.08 (0.75, 0.25) -0.20 (0.37, 0.63) nan (0.99, 0.01) 0.11 (0.25, 0.75) nan (0.99, 0.01) 1.26 (0.25, 0.75)
+$2^{13}$ 19 0.12 (0.75, 0.25) 0.01 (0.37, 0.63) nan (0.99, 0.01) 0.13 (0.25, 0.75) nan (0.99, 0.01) -3.37 (0.25, 0.75)
+$2^{14}$ 20 -0.03 (0.75, 0.25) -0.00 (0.37, 0.63) nan (0.99, 0.01) -4.52 (0.25, 0.75) nan (0.99, 0.01) 1.47 (0.25, 0.75)
+$2^{15}$ 21 0.05 (0.75, 0.25) -0.05 (0.37, 0.63) nan (0.99, 0.01) -0.70 (0.25, 0.75) nan (0.99, 0.01) -0.79 (0.25, 0.75)
+$2^{16}$ 22 0.10 (0.75, 0.25) 0.12 (0.38, 0.63) nan (0.99, 0.01) 1.57 (0.25, 0.75) nan (0.99, 0.01) 0.84 (0.25, 0.75)
+$2^{17}$ 23 -0.03 (0.75, 0.25) -0.15 (0.38, 0.63) nan (0.99, 0.01) -0.40 (0.25, 0.75) nan (0.99, 0.01) 0.05 (0.25, 0.75)
+$2^{18}$ 24 0.11 (0.25, 0.75) 0.00 (0.37, 0.63) nan (0.99, 0.01) -1.22 (0.25, 0.75) nan (0.99, 0.01) -3.58 (0.25, 0.75)
+$2^{19}$ 25 0.08 (0.75, 0.25) -0.03 (0.37, 0.63) nan (0.99, 0.01) 0.78 (0.25, 0.75) nan (0.99, 0.01) 1.41 (0.25, 0.75)
+$10^{6}$ $2^{20}$ 26 0.16 (0.75, 0.25) -0.19 (0.37, 0.63) nan (0.99, 0.01) -1.52 (0.25, 0.75) nan (0.99, 0.01) 0.01 (0.25, 0.75)
+Table A2: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=70^{\circ}$) for $4\times 4$ matrices.
+The notations are the same as those used in Table A1.
+Each data point is generated by taking the average over 10 repetitions.
+$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
+# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$)
+$2^{2}$ 8 -0.04 (0.65, 0.35) -0.24 (0.23, 0.77) -1.01 (0.92, 0.08) -0.03 (0.02, 0.98) -0.87 (0.92, 0.08) 0.58 (0.02, 0.98)
+$2^{3}$ 9 -0.55 (0.72, 0.28) -0.28 (0.25, 0.75) -5.82 (0.97, 0.03) 0.14 (0.06, 0.94) -5.72 (0.97, 0.03) -0.41 (0.06, 0.94)
+$2^{4}$ 10 -0.81 (0.74, 0.26) 0.51 (0.27, 0.73) 14.39 (0.99, 0.01) -0.35 (0.12, 0.88) 17.20 (0.99, 0.01) 0.44 (0.09, 0.91)
+$2^{5}$ 11 -0.17 (0.75, 0.25) 0.05 (0.29, 0.71) 3.96 (0.99, 0.01) -0.99 (0.10, 0.90) -0.12 (0.99, 0.01) 1.17 (0.10, 0.90)
+$2^{6}$ 12 0.30 (0.75, 0.25) 0.54 (0.30, 0.70) nan (0.99, 0.01) 1.40 (0.12, 0.88) -6.73 (0.99, 0.01) 0.99 (0.11, 0.89)
+$2^{7}$ 13 nan (0.75, 0.25) 0.74 (0.30, 0.70) nan (0.99, 0.01) 2.34 (0.12, 0.88) nan (0.99, 0.01) 2.20 (0.02, 0.88)
+$2^{8}$ 14 nan (0.75, 0.25) 0.40 (0.30, 0.69) nan (0.99, 0.01) 3.69 (0.12, 0.88) nan (0.99, 0.01) 3.54 (0.02, 0.88)
+$2^{9}$ 15 nan (0.75, 0.25) -0.31 (0.31, 0.69) nan (0.99, 0.01) -8.88 (0.12, 0.88) nan (0.99, 0.01) 1.96 (0.02, 0.88)
+$2^{10}$ 16 nan (0.75, 0.25) 0.02 (0.31, 0.69) nan (0.99, 0.01) 7.45 (0.12, 0.88) nan (0.99, 0.01) 4.93 (0.02, 0.88)
+$2^{11}$ 17 nan (0.75, 0.25) -0.30 (0.31, 0.69) nan (0.99, 0.01) 4.81 (0.12, 0.88) nan (0.99, 0.01) -2.36 (0.02, 0.88)
+$2^{12}$ 18 nan (0.75, 0.25) 0.48 (0.31, 0.69) nan (0.99, 0.01) -2.98 (0.12, 0.88) nan (0.99, 0.01) 0.81 (0.02, 0.88)
+$2^{13}$ 19 nan (0.75, 0.25) 0.38 (0.31, 0.69) nan (0.99, 0.01) 5.64 (0.12, 0.88) nan (0.99, 0.01) -2.22 (0.02, 0.88)
+$2^{14}$ 20 nan (0.75, 0.25) 0.15 (0.31, 0.69) nan (0.99, 0.01) 10.48 (0.12, 0.88) nan (0.99, 0.01) 0.33 (0.02, 0.88)
+$2^{15}$ 21 nan (0.75, 0.25) 1.26 (0.31, 0.69) nan (0.99, 0.01) 9.59 (0.12, 0.88) nan (0.99, 0.01) 6.98 (0.02, 0.88)
+$2^{16}$ 22 nan (0.75, 0.25) -0.34 (0.31, 0.69) nan (0.99, 0.01) 5.62 (0.12, 0.88) nan (0.99, 0.01) -17.07 (0.02, 0.88)
+$2^{17}$ 23 nan (0.75, 0.25) 1.31 (0.31, 0.69) nan (0.99, 0.01) 5.56 (0.12, 0.88) nan (0.99, 0.01) 2.02 (0.02, 0.88)
+$2^{18}$ 24 nan (0.75, 0.25) 0.00 (0.69, 0.31) nan (0.99, 0.01) 0.00 (0.12, 0.88) nan (0.99, 0.01) 18.88 (0.13, 0.87)
+$2^{19}$ 25 nan (0.75, 0.25) 0.33 (0.31, 0.69) nan (0.99, 0.01) 0.33 (0.12, 0.88) nan (0.99, 0.01) 6.42 (0.12, 0.88)
+$10^{5}$ $2^{20}$ 26 nan (0.75, 0.25) -0.27 (0.31, 0.69) nan (0.99, 0.01) -6.81 (0.12, 0.88) nan (0.99, 0.01) 6.75 (0.12, 0.88)
+$2^{2}$ 8 -0.57 (0.64, 0.36) -0.01 (0.23, 0.77) -0.87 (0.92, 0.08) -0.02 (0.02, 0.98) -0.88 (0.92, 0.08) -0.01 (0.02, 0.98)
+$2^{3}$ 9 -0.61 (0.72, 0.28) -0.16 (0.22, 0.76) -5.82 (0.97, 0.03) 0.04 (0.06, 0.94) -5.82 (0.97, 0.03) 0.00 (0.06, 0.94)
+$2^{4}$ 10 -0.63 (0.74, 0.26) -0.05 (0.27, 0.73) 14.66 (0.99, 0.01) -0.26 (0.09, 0.91) 14.95 (0.99, 0.01) -0.27 (0.09, 0.91)
+$2^{5}$ 11 0.27 (0.75, 0.25) -0.08 (0.29, 0.71) 1.84 (0.99, 0.01) -0.26 (0.10, 0.90) 0.61 (0.99, 0.01) 0.50 (0.11, 0.89)
+$2^{6}$ 12 0.04 (0.75, 0.25) 0.02 (0.30, 0.70) -0.72 (0.99, 0.01) 1.32 (0.11, 0.89) 1.68 (0.99, 0.01) -0.27 (0.12, 0.88)
+$2^{7}$ 13 -0.07 (0.75, 0.25) -0.07 (0.31, 0.69) 1.35 (0.99, 0.01) -0.24 (0.12, 0.88) -4.87 (0.99, 0.01) 0.40 (0.12, 0.88)
+$2^{8}$ 14 -0.13 (0.75, 0.25) 0.12 (0.31, 0.69) 0.09 (0.99, 0.01) 0.43 (0.12, 0.88) -4.04 (0.99, 0.01) 1.91 (0.12, 0.88)
+$2^{9}$ 15 0.05 (0.75, 0.25) 0.19 (0.31, 0.69) -2.82 (0.99, 0.01) -1.81 (0.12, 0.88) nan (0.99, 0.01) 1.12 (0.12, 0.88)
+$2^{10}$ 16 -0.04 (0.75, 0.25) -0.26 (0.31, 0.69) nan (0.99, 0.01) 1.19 (0.12, 0.88) nan (0.99, 0.01) -2.59 (0.12, 0.88)
+$2^{11}$ 17 0.18 (0.75, 0.25) 0.04 (0.31, 0.69) nan (0.99, 0.01) -1.68 (0.12, 0.88) nan (0.99, 0.01) -0.85 (0.12, 0.88)
+$2^{12}$ 18 0.07 (0.75, 0.25) -0.44 (0.31, 0.69) nan (0.99, 0.01) 2.36 (0.12, 0.88) nan (0.99, 0.01) -5.96 (0.12, 0.88)
+$2^{13}$ 19 0.12 (0.75, 0.25) 0.15 (0.31, 0.69) nan (0.99, 0.01) 5.19 (0.12, 0.88) nan (0.99, 0.01) 1.52 (0.12, 0.88)
+$2^{14}$ 20 -0.04 (0.75, 0.25) 0.25 (0.31, 0.69) nan (0.99, 0.01) 2.50 (0.12, 0.88) nan (0.99, 0.01) -1.89 (0.12, 0.88)
+$2^{15}$ 21 0.06 (0.75, 0.25) -0.33 (0.31, 0.69) nan (0.99, 0.01) 0.41 (0.12, 0.88) nan (0.99, 0.01) -3.65 (0.12, 0.88)
+$2^{16}$ 22 0.10 (0.75, 0.25) -0.01 (0.31, 0.69) nan (0.99, 0.01) -3.09 (0.12, 0.88) nan (0.99, 0.01) -1.88 (0.12, 0.88)
+$2^{17}$ 23 -0.03 (0.75, 0.25) -0.07 (0.31, 0.69) nan (0.99, 0.01) 2.54 (0.12, 0.88) nan (0.99, 0.01) -0.92 (0.12, 0.88)
+$2^{18}$ 24 0.11 (0.75, 0.25) 0.07 (0.31, 0.69) nan (0.99, 0.01) -0.47 (0.88, 0.12) nan (0.99, 0.01) -0.46 (0.88, 0.12)
+$2^{19}$ 25 -0.08 (0.75, 0.25) 0.09 (0.31, 0.69) nan (0.99, 0.01) 1.81 (0.12, 0.88) nan (0.99, 0.01) -0.89 (0.12, 0.88)
+$10^{6}$ $2^{20}$ 26 0.16 (0.75, 0.25) -0.02 (0.31, 0.69) nan (0.99, 0.01) -2.00 (0.12, 0.88) nan (0.99, 0.01) 3.65 (0.12, 0.88)
+Table A3: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=80^{\circ}$) for $4\times 4$ matrices.
+The notations are the same as those used in Tables A1 and A2.
+Each data point is generated by taking the average over 10 repetitions.
+$A$ diag, $\vec{b}$ equal $A$ diag, $\vec{b}$ unequal $A$ not diag, $\vec{b}$ unequal
+# shots $\mathcal{\approx\mathcal{\kappa}}$ qubits PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$)
+$2^{2}$ 8 -0.44 (0.65, 0.35) 0.43 (0.21, 0.79) -1.01 (0.92, 0.08) 0.06 (0.01, 0.99) -0.87 (0.92, 0.08) 0.31 (0.02, 0.98)
+$2^{3}$ 9 -0.55 (0.72, 0.28) 0.29 (0.27, 0.73) -5.82 (0.97, 0.03) -0.59 (0.02, 0.98) -5.72 (0.97, 0.03) -0.92 (0.02, 0.98)
+$2^{4}$ 10 -0.81 (0.74, 0.26) 2.47 (0.32, 0.68) 14.39 (0.99, 0.01) 2.59 (0.02, 0.98) 17.20 (0.99, 0.01) 1.55 (0.02, 0.98)
+$2^{5}$ 11 -0.17 (0.75, 0.25) 0.67 (0.35, 0.65) 3.96 (0.99, 0.01) 1.63 (0.03, 0.97) -0.12 (0.99, 0.01) 1.89 (0.03, 0.97)
+$2^{6}$ 12 0.30 (0.75, 0.25) -0.17 (0.36, 0.64) -6.73 (0.99, 0.01) 5.34 (0.04, 0.96) -6.73 (0.99, 0.01) -4.59 (0.03, 0.97)
+$2^{7}$ 13 nan (0.75, 0.25) -0.30 (0.37, 0.63) nan (0.99, 0.01) 0.55 (0.04, 0.96) nan (0.99, 0.01) 0.37 (0.04, 0.96)
+$2^{8}$ 14 nan (0.75, 0.25) -2.17 (0.37, 0.63) nan (0.99, 0.01) 6.89 (0.04, 0.96) nan (0.99, 0.01) 0.46 (0.04, 0.96)
+$2^{9}$ 15 nan (0.75, 0.25) 0.04 (0.37, 0.63) nan (0.99, 0.01) -2.64 (0.04, 0.96) nan (0.99, 0.01) 4.65 (0.04, 0.96)
+$2^{10}$ 16 nan (0.75, 0.25) 1.01 (0.37, 0.63) nan (0.99, 0.01) 3.21 (0.04, 0.96) nan (0.99, 0.01) -4.06 (0.04, 0.96)
+$2^{11}$ 17 nan (0.75, 0.25) -0.20 (0.37, 0.63) nan (0.99, 0.01) -5.46 (0.04, 0.96) nan (0.99, 0.01) -0.93 (0.04, 0.96)
+$2^{12}$ 18 nan (0.75, 0.25) -0.55 (0.37, 0.63) nan (0.99, 0.01) -4.80 (0.04, 0.96) nan (0.99, 0.01) 9.42 (0.04, 0.96)
+$2^{13}$ 19 nan (0.75, 0.25) -0.17 (0.37, 0.63) nan (0.99, 0.01) 2.18 (0.04, 0.96) nan (0.99, 0.01) 7.83 (0.04, 0.96)
+$2^{14}$ 20 nan (0.75, 0.25) 0.56 (0.37, 0.63) nan (0.99, 0.01) 5.32 (0.04, 0.96) nan (0.99, 0.01) -0.86 (0.04, 0.96)
+$2^{15}$ 21 nan (0.75, 0.25) -0.11 (0.37, 0.63) nan (0.99, 0.01) 5.89 (0.04, 0.96) nan (0.99, 0.01) 3.92 (0.04, 0.96)
+$2^{16}$ 22 nan (0.75, 0.25) 0.24 (0.37, 0.63) nan (0.99, 0.01) -12.26 (0.04, 0.96) nan (0.99, 0.01) 9.60 (0.04, 0.96)
+$2^{17}$ 23 nan (0.75, 0.25) -0.94 (0.37, 0.63) nan (0.99, 0.01) -2.94 (0.04, 0.96) nan (0.99, 0.01) -5.61 (0.04, 0.96)
+$2^{18}$ 24 nan (0.75, 0.25) -1.64 (0.26, 0.74) nan (0.99, 0.01) -0.60 (0.04, 0.96) nan (0.99, 0.01) 5.92 (0.04, 0.96)
+$2^{19}$ 25 nan (0.75, 0.25) 1.47 (0.27, 0.73) nan (0.99, 0.01) -0.60 (0.04, 0.96) nan (0.99, 0.01) -4.92 (0.04, 0.96)
+$10^{5}$ $2^{20}$ 26 nan (0.75, 0.25) -1.70 (0.26, 0.74) nan (0.99, 0.01) -17.20 (0.04, 0.96) nan (0.99, 0.01) 5.03 (0.04, 0.96)
+$2^{2}$ 8 -0.57 (0.64, 0.36) -0.13 (0.28, 0.72) -0.87 (0.92, 0.08) 0.00 (0.02, 0.98) -0.88 (0.92, 0.08) 0.03 (0.02, 0.98)
+$2^{3}$ 9 -0.61 (0.72, 0.28) 0.14 (0.24, 0.76) -5.82 (0.97, 0.03) -0.19 (0.01, 0.99) -5.82 (0.97, 0.03) -0.06 (0.01, 0.99)
+$2^{4}$ 10 -0.63 (0.74, 0.26) -0.03 (0.25, 0.75) 14.66 (0.99, 0.01) -0.22 (0.02, 0.98) 14.95 (0.99, 0.01) -0.02 (0.02, 0.98)
+$2^{5}$ 11 0.27 (0.75, 0.25) 0.64 (0.26, 0.74) 1.84 (0.99, 0.01) 0.03 (0.03, 0.97) 0.61 (0.99, 0.01) 1.11 (0.03, 0.97)
+$2^{6}$ 12 0.04 (0.75, 0.25) -0.05 (0.26, 0.74) -0.72 (0.99, 0.01) 1.21 (0.04, 0.96) 1.68 (0.99, 0.01) 0.68 (0.03, 0.97)
+$2^{7}$ 13 -0.07 (0.75, 0.25) 0.17 (0.26, 0.74) 1.35 (0.99, 0.01) -1.22 (0.04, 0.96) -4.87 (0.99, 0.01) 0.24 (0.03, 0.97)
+$2^{8}$ 14 -0.13 (0.75, 0.25) -0.17 (0.26, 0.74) 0.09 (0.99, 0.01) -3.19 (0.04, 0.96) -4.04 (0.99, 0.01) 0.32 (0.04, 0.96)
+$2^{9}$ 15 0.05 (0.75, 0.25) -0.50 (0.26, 0.74) 2.82 (0.99, 0.01) -2.13 (0.04, 0.96) nan (0.99, 0.01) -2.48 (0.04, 0.96)
+$2^{10}$ 16 -0.03 (0.75, 0.25) -0.30 (0.26, 0.74) nan (0.99, 0.01) -0.59 (0.04, 0.96) nan (0.99, 0.01) -0.24 (0.04, 0.96)
+$2^{11}$ 17 0.18 (0.75, 0.25) 0.47 (0.26, 0.74) nan (0.99, 0.01) -2.93 (0.04, 0.96) nan (0.99, 0.01) 0.05 (0.04, 0.96)
+$2^{12}$ 18 0.07 (0.75, 0.25) 0.40 (0.26, 0.74) nan (0.99, 0.01) 0.62 (0.04, 0.96) nan (0.99, 0.01) -2.73 (0.04, 0.96)
+$2^{13}$ 19 0.11 (0.75, 0.25) 0.14 (0.26, 0.74) nan (0.99, 0.01) 0.12 (0.04, 0.96) nan (0.99, 0.01) 6.03 (0.04, 0.96)
+$2^{14}$ 20 -0.03 (0.75, 0.25) 0.62 (0.26, 0.74) nan (0.99, 0.01) 2.52 (0.04, 0.96) nan (0.99, 0.01) 0.28 (0.04, 0.96)
+$2^{15}$ 21 0.05 (0.75, 0.25) -0.34 (0.26, 0.74) nan (0.99, 0.01) 1.03 (0.04, 0.96) nan (0.99, 0.01) -6.30 (0.04, 0.96)
+$2^{16}$ 22 0.10 (0.75, 0.25) -0.59 (0.26, 0.74) nan (0.99, 0.01) -3.21 (0.04, 0.96) nan (0.99, 0.01) -2.40 (0.04, 0.96)
+$2^{17}$ 23 -0.03 (0.75, 0.25) 0.49 (0.26, 0.74) nan (0.99, 0.01) 0.45 (0.04, 0.96) nan (0.99, 0.01) 2.72 (0.04, 0.96)
+$2^{18}$ 24 0.11 (0.75, 0.25) 0.14 (0.27, 0.73) nan (0.99, 0.01) -2.46 (0.04, 0.96) nan (0.99, 0.01) -1.52 (0.04, 0.96)
+$2^{19}$ 25 0.08 (0.75, 0.25) 0.01 (0.27, 0.73) nan (0.99, 0.01) -1.95 (0.04, 0.96) nan (0.99, 0.01) -0.72 (0.04, 0.96)
+$10^{6}$ $2^{20}$ 26 0.16 (0.75, 0.25) -0.33 (0.27, 0.73) nan (0.99, 0.01) -2.07 (0.04, 0.96) nan (0.99, 0.01) -2.46 (0.04, 0.96)
 In our Results section in the main text, we discussed the case of $A$ not diagonal and the amplitude encoded state corresponding to $\vec{b}$ in an unequal superposition.
@@
 In particular, $A=\begin{pmatrix}2^{-(n_{r}-1)}&0.00&0.00&0.00\\ 0.00&0.75&0.00&0.00\\ 0.00&0.00&0.50&0.00\\ 0.00&0.00&0.00&1.00\end{pmatrix}$, and the un-normalized $\vec{b}=\begin{pmatrix}1.00\\ 1.00\\ 1.00\\ 1.00\end{pmatrix}$.
-The sparsity, $s$, of $A$ is set to 1.
+The sparsity, $s$, of $A$ is 1.
   2. 2.
@@
 In particular, $A$ is the same as in case 1 given above, and $\vec{b}=\begin{pmatrix}0.10\\ 0.01\\ 0.20\\ 1.00\end{pmatrix}$.
-We also present these results for $\alpha$ set to 60, 70, and 80 degrees (the main text only discussed the first case).
+We also present these results for $\alpha$ set to 60, 70, and 80 degrees (the main text only discusses the first case).
 Furthermore, while the main text considers 1 million shots, we also perform the same analyses for $10^{5}$ shots.
-All of these results are presented in Figures A2 and A3 for the cases considering $10^{5}$ and $10^{6}$ shots respectively.
-The accompanying data is provided in Tables A1, A2, and A3, for 60, 70, and 80 degrees examples respectively.
+All of these results are presented in Figures A4 and A5 for the cases considering $10^{5}$ and $10^{6}$ shots respectively.
+The accompanying data is provided in Tables A4, A5, and A6, for 60, 70, and 80 degrees examples respectively.
 Refer to caption Refer to caption Refer to caption
@@
 The string ‘nan’ refers to situations where we get the square root of a negative quantity for energy, which is unphysical, and hence is not a valid result.
+Each data point is generated by taking the average over 50 repetitions.
 $A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
@@
 $10^{6}$ $2^{17}$ 23 0.02 (0.75, 0.25) -0.01 (0.37, 0.63) nan (0.99, 0.01) -0.66 (0.25, 0.75) nan (0.99, 0.01) 0.45 (0.25, 0.75)
-Table A2: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=70^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Table A1.
+Table A5: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=70^{\circ}$) for $4\times 4$ matrices.
+The notations are the same as those used in Table A4.
+Each data point is generated by taking the average over 50 repetitions.
 $A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
-# shots $\approx\kappa$ qubits PFD(P(0), P(1)) PFD(P’(0), P’(1)) PFD(P(0), P(1)) PFD(P’(0), P’(1)) PFD(P(0), P(1)) PFD(P’(0), P’(1))
+# shots $\approx\kappa$ qubits PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$) PFD ($P(0),P(1)$) PFD ($P^{\prime}(0),P^{\prime}(1)$)
 $2^{2}$ 8 -0.08 (0.64, 0.36) -0.03 (0.23, 0.77) 0.03 (0.92, 0.08) -0.09 (0.02, 0.98) 0.07 (0.92, 0.08) -0.05 (0.02, 0.98)
@@
 $2^{7}$ 13 -0.01 (0.75, 0.25) 0.28 (0.30, 0.70) nan (0.99, 0.01) 0.74 (0.12, 0.88) nan (0.99, 0.01) 1.88 (0.11, 0.88)
-$2^{8}$ 14 0.15(0.75, 0.25) -0.09(0.3, 0.69) nan(0.99, 0.01) 2.95(0.12, 0.88) nan(0.99, 0.01) -1.04(0.12, 0.88)
+$2^{8}$ 14 0.15 (0.75, 0.25) -0.09 (0.30, 0.69) nan (0.99, 0.01) 2.95 (0.12, 0.88) nan (0.99, 0.01) -1.04 (0.12, 0.88)
 $2^{9}$ 15 0.12 (0.75, 0.25) 0.57 (0.31, 0.69) nan (0.99, 0.01) 0.33 (0.12, 0.88) nan (0.99, 0.01) 2.59 (0.12, 0.88)
@@
 $10^{6}$ $2^{17}$ 23 0.03 (0.75, 0.25) 0.06 (0.31, 0.69) nan (0.99, 0.01) 1.02 (0.12, 0.88) nan (0.99, 0.01) 0.45 (0.12, 0.88)
-Table A3: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=80^{\circ}$) for $4\times 4$ matrices. The notations are the same as those used in Tables A1 and A2.
+Table A6: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=80^{\circ}$) for $4\times 4$ matrices.
+The notations are the same as those used in Tables A4 and A5.
+Each data point is generated by taking the average over 50 repetitions.
 $A$ diag $\vec{b}$ equal $A$ diag $\vec{b}$ unequal $A$ not diag $\vec{b}$ unequal
@@
 $10^{6}$ $2^{17}$ 23 0.02 (0.75, 0.25) -0.24 (0.26, 0.74) nan (0.99, 0.01) -2.53 (0.04, 0.96) nan (0.99, 0.01) 0.52 (0.04, 0.96)
-Table A4: Summary of the fitted probabilities of HHL and Psi-HHL for the 4x4 case $A$ not diagonal and $\vec{b}$ in an unequal superposition of states, $\alpha=60^{\circ},70^{\circ},80^{\circ}$ and the number of shots is fixed to $10^{6}$.
+Table A7: Summary of the fitted probabilities obtained from HHL and Psi-HHL for the 4x4 case of $A$ not diagonal and $\vec{b}$ in an unequal superposition, for $\alpha=60^{\circ},70^{\circ},80^{\circ}$.
+The number of shots is fixed to $10^{6}$.
 $\alpha$ $\approx\mathcal{\kappa}$ $P_{\text{Fit}}(0)$ $P_{\text{Fit}}(1)$ $P_{\text{Fit}}(0)+P_{\text{Fit}}(1)$ $P^{\prime}_{\text{Fit}}(0)$ $P^{\prime}_{\text{Fit}}(1)$ $P^{\prime}_{\text{Fit}}(0)+P^{\prime}_{\text{Fit}}(1)$
@@
 (d) $\alpha=60^{\circ}$ (e) $\alpha=70^{\circ}$ (f) $\alpha=80^{\circ}$
-Figure A4: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $8\times 8$ matrices using $10^{5}$ shots ((a) through (c)) and using $10^{6}$ ((d) through (f))
+Figure A6: Sub-figures presenting the average probabilities, PFDs and range of PFDs of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $8\times 8$ matrices using $10^{5}$ shots ((a) through (c)) and using $10^{6}$ ((d) through (f)).
+In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition.
+Each data point is an average over 10 repetitions.
 Refer to caption Refer to caption Refer to caption
 (a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
-Figure A5: Sub-figures presenting the average probabilities of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $16\times 16$ matrices and using $10^{6}$ shots throughout.
+Figure A7: Sub-figures presenting the average probabilities, PFDs and range of PFDs of HHL and Psi-HHL with varying condition number $\mathcal{\kappa}$ of $A$, for $16\times 16$ matrices and using $10^{6}$ shots throughout.
+In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition.
+Each data point is an average over 10 repetitions.
 Refer to caption Refer to caption Refer to caption
 (a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
-Figure A6: Data for $32\times 32$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
+Figure A8: Data for $32\times 32$ matrices.
+Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
+In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition.
+Each data point is an average over 10 repetitions.
 Refer to caption Refer to caption Refer to caption
 (a) $\alpha=60^{\circ}$ (b) $\alpha=70^{\circ}$ (c) $\alpha=80^{\circ}$
-Figure A7: Data for $64\times 64$ matrices. Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
-Figure A8: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) of HHL and Psi-HHL, as the condition number grows from $2^{2}$ to $2^{17}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$.
+Figure A9: Data for $64\times 64$ matrices.
+Note that we have again reproduced here the sub-figure (a) with 1 million shots (which are already present in the main text).
+In all of the examples, $A$ is chosen to be a non-diagonal matrix and $\vec{b}$ is set in an unequal superposition.
+Each data point is an average over 10 repetitions.
+Figure A10: Figure showing average (over 10 repetitions) values of percentage fraction difference (PFD) obtained from HHL and Psi-HHL, as the condition number, $\mathcal{\kappa}$, grows from $2^{2}$ to $2^{20}$, for fixed $A$ matrix sizes going from $4\times 4$ to $64\times 64$.
 Only the case where $\alpha=60^{\circ}$, $A$ not diagonal, $\vec{b}$ in an unequal superposition, and number of shots = $10^{6}$ is considered.
 The blue cells indicate cases where HHL gives outputs, and red denotes cases where HHL fails.
-The white cells denote those cases for which we did not carry a calculation, due to computational cost.
+The white cells denote those cases for which we did not carry out a calculation, due to computational cost.
 Table A8: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $8\times 8$ matrices.
+$A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition.
+Each data point is an average over 10 repetitions.
 HHL Psi-HHL
@@
 Table A9: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $16\times 16$ matrices.
+$A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition.
+Each data point is an average over 10 repetitions.
 HHL Psi-HHL
@@
 Table A10: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $32\times 32$ matrices.
+$A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition.
+Each data point is an average over 10 repetitions.
 HHL Psi-HHL
@@
 Table A11: Table providing the data for HHL and Psi-HHL algorithms (with $\alpha=60^{\circ}$, $70^{\circ}$, $80^{\circ}$) for $64\times 64$ matrices.
+$A$ is a non-diagonal matrix and $\vec{b}$ is an unequal superposition.
+Each data point is an average over 10 repetitions.
 HHL Psi-HHL
@@
 $2^{2}$ 16 -3.13 (0.82, 0.17) 0.01 (0.12, 0.88) 0.21 (0.09, 0.91) 0.82 (0.10, 0.90)
-$2^{3}$ 17 -2.26(0.94, 0.05) -1.06(0.012, 0.88) -0.95(0.04, 0.96) -1.01(0.02, 0.98)
+$2^{3}$ 17 -2.26 (0.94, 0.05) -1.06 (0.01, 0.88) -0.95 (0.04, 0.96) -1.01 (0.02, 0.98)
 $2^{4}$ 18 -14.67 (0.98, 0.02) -1.82 (0.17, 0.83) 1.47 (0.03, 0.97) -1.56 (0.02, 0.98)
@@
 Appendix A2 Details on calculations involving quantum chemistry matrices
-A2.1 $A$ matrix and $\vec{b}$ vector generation
-We begin by explaining the procedure for generating the $A$ matrix of size $n$.
+A2.1 $A$ matrix and $\vec{b}$ generation
+We begin by explaining the procedure for generating an $A$ matrix of size $n\times n$.
 We use the configuration state function-based Graphical Unitary Group Approach-Configuration Interaction Singles and Doubles (GUGA-CISD) method available in the GAMESS program [27] to construct the Hamiltonian for a given molecule in a specified single particle basis.
@@
 Also, because it is a spin-adapted CI method, it avoids spin contamination of the wave function, thereby giving accurate results.
-We remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature).
-The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the HF energy and the correlation energy, with the latter being much lesser than the former).
-This is desirable when one executes HHL-LCC calculations on NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself.
+In order to generate the $A$ matrix, we remove the first row and the first column of the GUGA-CISD Hamiltonian matrix and subtract the diagonal entries of the resulting matrix with the energy obtained from Møller-Plesset theory to second order in perturbation (also called MP2 in literature).
+The latter step is carried out for the purpose of normal ordering, which in turn ensures that HHL would output the correlation part of the ground state energy directly instead of the total ground state energy (which can be written as a sum of the HF energy and the correlation energy, with the latter being much smaller than the former).
+This is desirable when one executes HHL-LCC calculations on NISQ/late-NISQ era quantum hardware, as the error in the total ground state energy can be much greater than the correlation energy itself.
 We also note that unlike the traditional approach where the HF energy is subtracted from the diagonal entries, we take away the MP2 energy, in order to make sure that $A$ is positive definite.
@@
 We now focus our attention on the elements of the $A$ matrix.
-The size of an $A$ matrix, $m$, generated from a quantum chemical calculation, which we can denote as $A_{mol}$, is not necessarily of the dimension $n\times n$.
+The size of an $A$ matrix, $m\times m$, generated from a quantum chemical calculation, which we can denote as $A_{mol}$, is not necessarily of the dimension $n\times n$.
 In such cases, we embed the $A_{mol}$ matrix of dimension ${m\times m}$ as a diagonal block in a larger matrix of dimension $n\times n$, with the other diagonal block being an identity matrix of size $(n-m)\times(n-m)$, denoted as $\mathcal{I}$.
 In such a case, we solve a system of linear equations given by
-$A=\begin{bmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{bmatrix}\begin{bmatrix}\vec{x}\\ 0\end{bmatrix}=\begin{bmatrix}\vec{b}\\ 0\end{bmatrix}.$ (A25)
-This approach ensures that the eigenvalues of $A_{mol}$ remain unaffected, as long as its maximum eigenvalue is $\approx$ 1.
+$A=\begin{pmatrix}A_{mol}&0\\ 0&\mathcal{I}\end{pmatrix}\begin{pmatrix}\vec{x}\\ \vec{0}\end{pmatrix}=\begin{pmatrix}\vec{b}\\ \vec{0}\end{pmatrix}.$ (A25)
+This approach ensures that the eigenvalues of $A_{mol}$ remain unaffected, as long as its maximum eigenvalue of the scaled matrix is $\approx$ 1.
 We choose our matrices carefully so that this condition is satisfied.
@@
 As an example, we picked a $3\times 3$ sparse matrix such that the largest diagonal entry is 0.70: $\begin{pmatrix}0.25&0.10&0.00\\ 0.10&0.45&0.20\\ 0.00&0.20&0.70\end{pmatrix}$, and considered two cases: padding with 1 and with 0.70.
-We found that kappa is larger in the latter (5.16) than in the former (4.21).
+We found that the condition number, $\mathcal{\kappa}$, is larger in the latter (5.16) than in the former (4.21).
 Since for the problems we consider, $\mathcal{\kappa}$ increases with system size monotonically, and further, going from an unpadded to a padded matrix always right shifts a data point, it is sufficient to carry out $\mathcal{\kappa}$ versus $m$ analysis for the unpadded case, as it upper bounds the complexity of the padded scenario of $\mathcal{\kappa}$ versus $n$.
@@
 (c) (d)
-Figure A9: Sub-figure (a) shows the increase in system size with number of virtuals, $n_{v}$, for KH, RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b), (c) and (d) show that the $A$ matrix of respectively KH, RbH and CsH increases as $n_{o}^{2}n_{v}^{2}$ for both cases, as one would expect.
+Figure A11: Sub-figure (a) shows the increase in system size ($m$) with the number of virtuals, $n_{v}$, for KH, RbH and CsH molecules in the LCCSD approximation, while the sub-figures (b), (c) and (d) show that $m$ increases as $n_{o}^{2}n_{v}^{2}$ for the three molecules, as one would expect.
 A2.2 Orbital character analysis
@@
 We begin with RbH, and then proceed to explain our procedure for CsH (with accompanying data provided in Table A12).
-We fix the number of core orbitals to 6, 8 and 10.
+We fix the number of occupied orbitals to 6, 8 and 10.
 We now describe the method adopted in choosing virtual molecular orbitals (MOs).
@@
 For the purposes of our analysis on condition number versus system size, we pick all the AOs that the basis allows in order to obtain more data points for fitting (details given in Table A12).
-We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 6 occupied MOs and 1 virtual one has a system size, $m$, of $209\times 209$, and subsequently upon padding, an $A$ matrix size, $n$, of $256\times 256$.
+We now move to some specifics for RbH: for the smallest system size considered (1 virtual MO), we pick the first virtual from our sorted MO list, which is the MO dominated by the Rb $5s-5p_{z}$ hybrid, and thus our active space that is built out of 6 occupied MOs and 1 virtual one has a system size, $m\times m$, of $209\times 209$, and subsequently upon padding, an $A$ matrix size, $n\times n$, of $256\times 256$.
 For the next system size, we add the next MO in the sorted list, which is dominated by Rb $5s$, leading to a $279\times 279$ matrix, which, after padding, becomes a matrix of size $512\times 512$.
-We continue to construct $A$ matrices in this fashion until we reach $m=2296$, for which $\mathcal{\kappa}$ is about 389.76.
+We continue to construct $A$ matrices in this fashion until we reach $m=2296$, for which $\mathcal{\kappa}$ is about 211.94.
 We do the same for KH and CsH molecules, we fix the number of occupied orbitals to 6, 8 and 10 occupied oritals while progressively increasing the number of valence orbitals.
 The order in which we add virtual MOs as well as each such virtual MO’s details are mentioned in Table A12.
-Table A9: Table providing the data for condition number ($\mathcal{\kappa}$) of the $m\times m$ $A$ matrix versus system size ($m$), for KH, RbH and CsH molecules, where we vary the virtual orbitals to increase $m$.
+Table A12: Table providing the data for condition number ($\mathcal{\kappa}$) of the $m\times m$ $A_{mol}$ matrix versus system size ($m$), for KH, RbH and CsH molecules, where we vary the virtual orbitals to increase $m$.
 In the table, $n_{o}$ is the number of active spin orbitals, and $n_{v}$ is the number of unoccupied spin orbitals.
 The column labelled ‘Reordered virtual MOs’ represents the atomic orbitals reordered according to the ones that contribute dominantly to the virtual molecular orbital.
-The column labelled ‘MO Energies’ gives the molecular virtual orbital energies in units of milli-Hartree.
+The column labelled ‘MO energies’ gives the virtual molecular orbital energies in units of milli-Hartree.
 Molecule Reordered virtual MOs MO energies $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
 m $\kappa$ $(n_{o},n_{v})$ m $\kappa$ $(n_{o},n_{v})$ m $\kappa$ $(n_{o},n_{v})$
-K $4s$ 1151.70 235 233.65(6, 6) 580 41.03(8, 8) 1484 2195.80(10, 10)
-K $4p_{y}$ 3609.70 313 234.44(6, 7) 745 41.57(8, 9) 1761 2206.59(10, 11)
-K $4p_{x}$ 3609.70 370 235.12(6, 8) 932 236.57(8, 10) 2046 2217.42(10, 12)
+K $4s$ 1151.70 202 233.65 (6, 6) 580 41.03 (8, 8) 1484 2195.80 (10, 10)
+K $4p_{y}$ 3609.70 274 234.44 (6, 7) 745 41.57 (8, 9) 1761 2206.59 (10, 11)
+K $4p_{x}$ 3609.70 337 235.12 (6, 8) 932 236.57 (8, 10) 2046 2217.42 (10, 12)
 K $4p_{z}$ 3663.70 432 236.80 (6, 9) 1110 237.74 (8, 11) 2451 2323.69 (10, 13)
 K $3d_{xy}$ 743.00 541 238.28 (6, 10) 1296 238.90 (8, 12) 2900 2849.98 (10, 14)
-K $3dy^{2}$ 743.00 630 244.66(6, 11) 1540 250.34(8, 13) 3393 2904.40(10, 15)
+K $3dy^{2}$ 743.00 655 244.66 (6, 11) 1540 250.34 (8, 13) 3393 2904.40 (10, 15)
 K $3d_{yz}$ 747.40 748 250.45 (6, 12) 1808 307.03 (8, 14) 3794 2933.74 (10, 16)
 K $3d_{xz}$ 747.40 889 265.31 (6, 13) 2100 312.83 (8, 15) 4203 2964.04 (10, 17)
-K $3dz^{2}$ 755.30 1044 340.59(6, 14) 2358 316.09(8, 16) 4788 3610.73(10, 18)
+K $3dz^{2}$ 755.30 1043 340.59 (6, 14) 2358 316.09 (8, 16) 4788 3610.73 (10, 18)
 K $6s$ 3.70 1213 348.68 (6, 15) 2624 319.44 (8, 17) 5265 3629.57 (10, 19)
-K $6p_{y}$ 72.70 1338 371.54(6, 16) 2972 389.11(8,18) 5750 3648.68(10,20)
+K $6p_{y}$ 72.70 1381 371.54 (6, 16) 2972 389.11 (8, 18) 5750 3648.68 (10, 20)
 K $6p_{x}$ 72.70 1510 374.07 (6, 17) 3278 391.14 (8, 19) 6427 3674.83 (10, 21)
 K $6p_{z}$ 82.80 1711 380.25 (6, 18) 3592 393.20 (8, 20)
-K $5s$ 147.20 1860 385.51(6, 19) 3996 395.77(8, 21)
+K $5s$ 147.20 1909 385.51 (6, 19) 3996 395.77 (8, 21)
 K $5p_{z}$ 692.70 2062 386.30 (6, 20)
-K $5p_{y}$ 702.90 2296 389.76(6, 21)
+K $5p_{y}$ 702.90 2295 389.76 (6, 21)
 K $5p_{x}$ 702.90
@@
 KH H $2p_{z}$ 1914.10
-Rb $5s-5p_{z}$ 940.00 209 102.73(6, 6) 632 104.88(8, 8) 1401 108.16(10,10)
+Rb $5s-5p_{z}$ hybrid 940.00 209 102.73 (6, 6) 632 104.88 (8, 8) 1401 108.16 (10, 10)
 Rb $5s$ 925.80 279 103.00 (6, 7) 744 105.81 (8, 9) 1673 111.85 (10, 11)
@@
 Rb $6s$ 130.80 2062 210.72 (6, 20)
-Rb $6p_{x}$ 477.70 2296 211.94(6, 21)
+Rb $6p_{x}$ 477.70 2295 211.94 (6, 21)
 Rb $6p_{y}$ 477.70
@@
 Cs $7p_{x}$ 25.50 759 97.83 (6, 12) 1808 112.65 (8, 14) 3537 184.86 (10, 16)
-Cs $7s$ 75.60 890 101.83(6, 13) 2100 156.01(8, 15) 3951 186.03(10,17)
+Cs $7s-7p_{y}$ hybrid 75.60 890 101.83 (6, 13) 2100 156.01 (8, 15) 3951 186.03 (10, 17)
 Cs $7s$ 85.50 1044 111.49 (6, 14) 2358 182.96 (8, 16) 4461 192.15 (10, 18)
@@
 CsH H $2p_{z}$ 1716.80
-Table A11: Table providing the minimum and maximum eigenvalues of the matrix $\mathrm{A}_{mol}$ of size $m$, denoted in the table as $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ and $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ respectively, for KH, RbH and CsH systems.
-$n_{o}$ and $n_{v}$ denote the number of occupied and virtual orbitals respectively.
-The energies are in units of milli-Hartree.
+Table A14: Table providing the minimum and maximum eigenvalues of the matrix $A_{mol}$ of size $m\times m$, denoted in the table as $\lambda_{\mathrm{min}}^{A_{mol}}$ and $\lambda_{\mathrm{max}}^{A_{mol}}$ respectively, for KH, RbH and CsH systems.
+$n_{o}$ and $n_{v}$ denote the number of occupied and virtual orbitals, respectively.
+$\lambda_{\mathrm{min}}^{A_{mol}}$ and $\lambda_{\mathrm{max}}^{A_{mol}}$ are in units of milli-Hartree.
 Molecule $n_{o}=6$ $n_{o}=8$ $n_{o}=10$
-m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{\mathrm{A_{mol}}}$ $\lambda_{\mathrm{max}}^{\mathrm{A_{mol}}}$
+m $n_{v}$ $\lambda_{\mathrm{min}}^{A_{mol}}$ $\lambda_{\mathrm{max}}^{A_{mol}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{A_{mol}}$ $\lambda_{\mathrm{max}}^{A_{mol}}$ m $n_{v}$ $\lambda_{\mathrm{min}}^{A_{mol}}$ $\lambda_{\mathrm{max}}^{A_{mol}}$
 202 6 733.73 25918.07 580 8 730.05 29960.23 1484 10 127.91 280908.57
@@
 RbH 2295 21 70.15 14868.69
-209 6 119.60 9216.72 632 8 114.40 9306.7 1484 10 99.94 9447.99
-279 7 119.41 9216.73 812 9 101.3 9338.37 1711 11 98.74 9448.39
+209 6 119.60 9216.72 632 8 114.40 9306.70 1484 10 99.94 9447.99
+279 7 119.41 9216.73 812 9 101.30 9338.37 1711 11 98.74 9448.39
 370 8 114.52 9224.27 1016 10 99.97 9338.71 2007 12 94.46 9448.85
@@
 CsH 1711 18 63.04 11884.73
-Table A12: Average PFDs and Range of PFDs of HHL and Psi-HHL for Li2 and KH molecules, both in Sapporo-DKH3-DZP-2012 basis set. The PFDs are calculated over a range of 30 repetitions.
+Table A15: Average PFDs and range of PFDs of HHL and Psi-HHL for Li2 and KH molecules, both in Sapporo-DKH3-DZP-2012 basis set. The average PFDs are calculated over a range of 30 repetitions.
 Li2 KH
@@
 $10^{6}$ (3.99, 5.09) (0.22, 4.05)
-Table A13: Table providing the details on the condition numbers ($\kappa$), the average total energy, the energy difference ($\Delta E$) between the total HHL/Psi-HHL energy and the LCCSD total energy and the PFDs over 30 repetitions with respect to the bond length of KH molecule.
+Table A16: Table providing the details on the condition numbers ($\kappa$), the average total energy, the energy difference ($\Delta E$) between the total HHL (Psi-HHL) energy and classical value of the LCCSD total energy and the average PFDs over 30 repetitions with respect to the bond length of KH molecule.
 All the energies are in units of milli-Hartree.
+‘Eq’ denotes the equilibrium geometry.
 Bond length (Bohr) $\kappa$ HHL Psi-HHL
@@
 3.76 367.32 -598967.447 -0.003 -1.96 -598967.447 -0.004 -0.26
-4.23 365.40 -598975.800 0.007 1.46 -598975.802 0.004 0.34
+4.23 (Eq) 365.40 -598975.800 0.007 1.46 -598975.802 0.004 0.34
 4.71 371.38 -598974.517 -0.010 -10.00 -598974.509 -0.007 1.39